### v3

# Corresponding Author

Abstract
Recent advancements in diffusion models have significantly impacted content creation, leading to the emergence of Personalized Content Synthesis (PCS). By utilizing a small set of user-provided examples featuring the same subject, PCS aims to tailor this subject to specific user-defined prompts. Over the past two years, more than 150 methods have been introduced in this area. However, existing surveys primarily focus on text-to-image generation, with few providing up-to-date summaries on PCS. This paper provides a comprehensive survey of PCS, introducing the general frameworks of PCS research, which can be categorized into test-time fine-tuning (TTF) and pre-trained adaptation (PTA) approaches. We analyze the strengths, limitations, and key techniques of these methodologies. Additionally, we explore specialized tasks within the field, such as object, face, and style personalization, while highlighting their unique challenges and innovations. Despite the promising progress, we also discuss ongoing challenges, including overfitting and the trade-off between subject fidelity and text alignment. Through this detailed overview and analysis, we propose future directions to further the development of PCS.
keywords:
Generative Models, Image Synthesis, Diffusion Models, Personalization
1 Introduction
Recently, generative models have shown remarkable progress in the field of natural language processing and computer vision. Notable examples, such as ChatGPT wu2023brief and text-to-image diffusion models croitoru2023diffusion , have showcased impressive capabilities in content creation. However, these advanced models often struggle to fulfill specific requirements, such as answering domain-specific queries or accurately depicting user’s portraits. This limitation highlights the importance of Personalized Content Synthesis (PCS), which enables users to customize models for their unique tasks and requirements. As a critical area of research, PCS is increasingly recognized as essential for achieving Artificial General Intelligence (AGI), evidenced by the growing number of companies releasing products that support personalized content creation, like utilizing Reinforcement Learning to fine-tune language models uc2023survey . In this paper, we focus on PCS within the context of diffusion models in computer vision. The objective of PCS is to learn the subject of interest (SoI) from a small set of user-uploaded samples and generate images that align with user-defined contexts. For instance, as illustrated in Fig. 1, PCS can customize a unique cat into various scenarios, such as “wearing pink sunglasses” and “in the snow”.
Figure 1: Given a few reference images of a subject (e.g., a cat 2ruiz2023dreambooth or face 54xiao2023fastcomposer ), PCS aims to generate new renditions of the subject that align with user-defined textual prompts. The task requires preserving the subject’s identity (fidelity) while adapting to diverse contexts. The examples are generated by this survey using DreamBooth 2ruiz2023dreambooth and InstantID 37wang2024instantid .
The emergence of diffusion models has significantly facilitated text-guided content generation, leading to a rapid expansion of PCS. As depicted in Fig. 2, the number of research papers on PCS has surged over time. Following the release of key innovations such as DreamBooth 2ruiz2023dreambooth and Textual Inversion 1gal2022image in August 2022, over 150 methods have been proposed in a remarkably short period. These methods can be classified using several criteria. First, in terms of training strategy, we differentiate between test-time fine-tuning (TTF) and pre-trained adaptation (PTA) approaches. TTF methods fine-tune generative models for each personalization request during inference phase, while PTA methods strive to pre-train a unified model capable of generating a wide range of subjects of interest (SoI). Second, we categorize the techniques employed in PCS into four primary areas: attention-based operations, mask-guided generation, data augmentation, and regularization. As shown in Fig. 2, we summarize various influential PCS methods based on these criteria. Third, the scope of personalization has expanded considerably. The SoI now includes not only general objects but also extends to human faces, artistic styles, actions, and other intricate semantic elements. This growing interest encompasses the creation of complex compositions and the combination of several conditions. Additionally, the research landscape has broadened from static images to other modalities such as video and 3D representations.
Figure 2: A chronological overview of classical PCS methods as surveyed, illustrating the evolution of techniques through months. The number of related works has rapidly increased over the past two years. We divide PCS methods with 3 different criteria: training strategy, personalization scope, and technique.
While current methods in PCS have shown impressive performance, several challenges remain unresolved. A primary concern is the issue of overfitting, which often arises from the limited number of reference images available. This limitation can lead to a neglect of the textual context in the generated outputs, as shown in Fig. 3. Another related challenge is the trade-off between image alignment and text fidelity, as illustrated in Fig. 3. When a model successfully reconstructs the fine-grained details of the subject of interest (SoI), it often sacrifices controllability. Conversely, enhancing the model’s editability can result in underfitting with a loss of fidelity to the original SoI. Additionally, there are other challenges including the absence of robust evaluation metrics, a lack of standardized test datasets, and the need for faster processing times. This paper discusses these challenges in detail and establishes a benchmark for evaluating classical methods. By addressing these obstacles, we aim to advance the field of personalized content synthesis and enhance its practical applications.
Figure 3: The trade-off between text alignment and visual fidelity in personalized image synthesis, illustrated through DreamBooth-generated 2ruiz2023dreambooth examples of a customized cat wearing sunglasses. Overfitting occurs when the model focuses solely on reconstructing the cat, disregarding the sunglasses context. Underfitting, on the other hand, reflects the model’s attempt to satisfy the text prompt but fails to accurately represent the personalized cat. Collapse signifies a failure to meet both criteria.
This survey aims to provide a comprehensive overview of the existing methods, frameworks, and challenges of PCS to help readers understand current trends in this evolving landscape and promote further improvements in this area. The structure of this survey is organized into several key sections: In Sec. 2, We begin with a brief introduction to diffusion models for better understanding of PCS. In Sec. 3, we introduce two primary frameworks for PCS: test-time fine-tuning and pre-trained adaptation. In Sec. 4, we categorize common techniques into four main parts, attention-based methods, mask-based generation, data augmentation, and regularization. In Sec. 5, we summarize various PCS methods based on different personalization tasks, including personalization on object, face, style, and etc. In Sec. 6, we demonstrate PCS’s growing impact in video, 3D, and other areas beyond traditional image generation tasks. In Sec. 7, we review the current evaluation dataset and metrics and construct a benchmark on the existing methods for further research. In Sec. 8, we discuss current challenges faced in PCS and propose potential future directions to overcome these problems.
2 Fundamentals
Modern diffusion modeling has evolved into a sophisticated framework that unifies discrete and continuous generative paradigms through stochastic differential equations (SDEs) song2020score and ordinary differential equations (ODEs) song2020score ; karras2022elucidating . This section introduces the core mathematical foundations, emphasizing innovations in the diffusion process and conditional mechanisms critical for personalized generation.
2.1 Denoising Diffusion Probabilistic Models (DDPMs)
The architecture of diffusion models comprises two complementary processes: a forward diffusion that systematically perturbs data distributions, and a learned reverse process that reconstructs signals through iterative refinement. Originating from discrete-time Markov chains in Denoising Diffusion Probabilistic Model (DDPM) ho2020denoising , the forward process applies Gaussian noise corruption over $T$ steps:
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},\quad\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ (1)
where $\mathbf{x}_{t}$ denotes the noised data at diffusion step $t$, $\mathbf{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ is the standard Gaussian noise vector that corrupts the original data $\mathbf{x}_{0}$, and $\alpha_{s}$ governs the noise scheduling policy.
The generative capability resides in learning to invert this degradation through parameterized reverse transitions. For DDPMs, this inversion is achieved via Bayesian reconstruction:
$p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t}\right)=\mathcal{N}\left(\mathbf{x}_{t-1};\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t)\right),\beta_{t}I\right)$ (2)
Here, $p_{\theta}$ represents the parameterized reverse transition probability, $\epsilon_{\theta}$ is the neural network predicting noise components, and $\beta_{t}$ is a hyperparameter before model training. And the neural network $\epsilon_{\theta}$ learns to predict the applied noise component $\epsilon_{\theta}$ through:
$\mathcal{L}=\mathbb{E}_{\mathbf{x}_{0},t,\mathbf{\epsilon}}\left[\|\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}(\mathbf{x}_{t},t)\|_{2}^{2}\right]$ (3)
This objective enables end-to-end training of the denoising trajectory.
2.2 Stochastic Differential Equations (SDEs)
Although DDPMs establish basic denoising dynamics, this discrete formulation imposes constraints on flexible noise scheduling and computational efficiency.
To overcome these limitations, the framework was generalized through continuous-time SDE song2020score , which provide a unified perspective for modeling diffusion dynamics:
$d\mathbf{x}=\mathbf{f}(\mathbf{x},t)dt+g(t)d\mathbf{w}$ (4)
where $\mathbf{f}(\mathbf{x},t)$ encodes deterministic drift components, $g(t)$ modulates stochastic diffusion via Wiener process $\mathbf{w}$. This continuous perspective subsumes DDPMs as special cases while enabling adaptive noise scheduling strategies such as variance-preserving (VP) and variance-exploding (VE) schedules through strategic choices of $\mathbf{f}$ and $g$ karras2022elucidating .
For the reverse process, a crucial result anderson1982reverse shows that the reverse diffusion process also follows an SDE:
$d\mathbf{x}=\left[\mathbf{f}(\mathbf{x},t)-g(t)^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right]dt+g(t)d\mathbf{\bar{w}}$ (5)
where $\bar{w}$ denotes reverse-time Wiener increments, and $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ is a score function. This formulation enables generation by integrating backward from $\mathbf{x}_{T}\sim\mathcal{N}(0,I)$ to $\mathbf{x}_{0}$, conditioned on learned score estimates.
2.3 Ordinary Differential Equations (ODEs)
While SDEs provide a comprehensive framework for modeling diffusion dynamics, their inherent randomness introduces critical challenges in practical deployment. The Wiener process $\mathbf{w}$ in SDE-based sampling generates pathwise variability, requiring extensive Monte Carlo averaging to converge to stable solutions. Furthermore, stochastic trajectories exhibit erratic curvature profiles, forcing fixed-step solvers to adopt conservative discretization schemes that often necessitate 1,000+ steps for high-fidelity generation.
These limitations motivate the derivation of deterministic sampling trajectories through probability flow ordinary differential equations (ODEs) song2020score ; karras2022elucidating , obtained by eliminating the stochastic term from the SDE formulation:
$d\mathbf{x}=\left[\mathbf{f}(\mathbf{x},t)-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right]dt$ (6)
The ODE’s deterministic nature arises from two synergistic components: the original drift term $\mathbf{f}(\mathbf{x},t)$ and a correction term involving the score function $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$. This score-guided adjustment preserves the marginal data distribution $p_{t}(\mathbf{x})$ while eliminating pathwise stochasticity. Crucially, the resulting trajectories exhibit intrinsic geometric regularity, allowing the complex denoising process to operate in a simplified mathematical space. This enabling 5-10x faster sampling than SDEs through adaptive ODE solvers like DPM-Solver lu2022dpm and DPM-Solver++ lu2022dpmplus . Recent work chen2024trajectory further identifies strong geometric regularity in these ODE-based sampling trajectories, demonstrating that they inherently follow a linear-nonlinear-linear structure regardless of generated content. This trajectory regularity enables dynamic programming-based time scheduling optimization with negligible computational overhead.
2.4 Conditional Generation Mechanisms
Recently, conditional synthesis has emerged as the critical capability bridging theoretical diffusion frameworks with real-world applications. It enables precise alignment of outputs with multimodal guidance signals (text prompts, subject embeddings, anatomical masks), fulfilling domain-specific requirements for reliability and reproducibility.
Building upon the unconditional framework, conditional synthesis is formalized through extended score matching:
$\mathcal{L}=\mathbb{E}_{\mathbf{x}_{0},t,\mathbf{\epsilon}}\left[\|\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}(\mathbf{x}_{t},t,c)\|_{2}^{2}\right]$ (7)
where the conditioning signal $c$ can be integrated via multiple synergistic mechanisms, such as cross-modal attention rombach2022high and spatial modulation zhang2023adding .
This conditional paradigm directly enables the applications of PCS. Leading text-to-image systems, such as Stable Diffusion (SD) rombach2022high and DALLE ramesh2022hierarchical , are widely adopted to empower users in controlling customized content through text instructions.
3 Generic Framework
In this survey, we broadly categorize PCS frameworks into two paradigms: TTF and PTA approaches. These two frameworks fundamentally differ in their adaptation mechanisms. TTF methods dynamically adjust model parameters for each new subject during inference, prioritizing visual fidelity at the cost of computational overhead. Conversely, PTA frameworks employ reference-aware architectures trained on large datasets to enable single-pass personalization without parameter updates during inference. We introduce two generic frameworks in the following sections.
3.1 Test-time Fine-tuning (TTF) Framework
Figure 4: Illustration of the TTF framework for the test-time fine-tuning process and generation phase. During the inference phase, the model fine-tunes its parameters by reconstructing the reference images for each SoI group. The unique modifier V* is employed to represent the SoI and used to formulate new inference prompts for generating personalized images.
TTF methods represent the foundational approach for personalized synthesis, where models adapt to new subjects through instance-specific optimization during inference. As illustrated in Fig. 4, this framework operates through two core principles: 1) Test-time adaptation fine-tunes model parameters to learn the key visual element of the SoI. 2) Semantic-aware modifier system represents the SoI at the token level to bridge the gap between visual adaptation and textual control. In the following Sec. 3.1.1 and Sec. 3.1.2, we will detail these principles respectively.
3.1.1 Test-time Adaption
Test-time fine-tuning. For each reference set $\mathbf{X}_{SoI}$ of SoI, the optimization process adjusts a subset of parameters $\mathbf{\theta^{\prime}}$ to reconstruct the SoI conditioned on the reference prompt. The fine-tuning objective is defined by a reconstruction loss:
$\mathcal{L}=\mathbb{E}_{\mathbf{x}_{0}\in\mathbf{X}_{SoI},t,\mathbf{\epsilon}}\left[\|\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta^{\prime}}(\mathbf{x}_{t},t,c)\|_{2}^{2}\right]$ (8)
where $c$ represents the condition signal, typically the reference image caption. Compared to the large-scale pre-training described in Eq. 7, key differences lie in the training data and learnable parameters. The training samples are typically restricted to the SoI references, sometimes supplemented with a regularization dataset to mitigate overfitting 2ruiz2023dreambooth . For the selection of learnable parameters $\mathbf{\theta^{\prime}}$, the commonly adopted options include token embeddings 1gal2022image ; 23voynov2023p+ , the entire diffusion model 2ruiz2023dreambooth ; 151zhang2024generative , specific subsets of parameters 6kumari2023multi ; 7tewel2023key ; 8han2023svdiff , or introducing new parameters such as adapters 72xiang2023closer ; 5sohn2024styledrop , and LoRA 46gu2024mix ; 58ruiz2023hyperdreambooth ; 115kong2024omg , which will be discussed in Sec. 3.1.3.
Test-time generation. The test-time generation is proceeded once the model has been fine-tuned with the optimized parameters $\mathbf{\theta^{\prime}}$. By composing novel input prompts that incorporate the SoI’s unique identifier (introduced in Sec. 3.1.2), the adapted model can synthesize diverse images while preserving the subject’s distinctive characteristics. This approach enables flexible control over the generated content through natural language descriptions.
3.1.2 Unique Modifier
The unique modifier is a critical component in PCS, enabling text-based representation of the SoI for flexible prompt composition. As illustrated in Fig. 4, this modifier serves as a textual description for the SoI, allowing combination with other descriptions (e.g., “V* on the beach”) during inference. Normally, the construction of the unique modifier can be divided into three categories:
Learnable token embedding 1gal2022image ; 23voynov2023p+ ; 61alaluf2023neural ; 30zhang2023prospect ; 48wang2023hifi ; 66pang2023cross . This method adds a new token and its corresponding embedding vector to the word dictionary of the tokenizer. An intuitive example is that this method creates a new word that does not exist in the dictionary. This word acts as the modifier, with adjustable weights during fine-tuning, while the embeddings of other tokens in the pre-defined dictionary will not be influenced. This approach requires only a few kilobytes of additional parameters and maintains the base model’s capabilities for non-customized generation.
Plain text 2ruiz2023dreambooth ; 6kumari2023multi ; 7tewel2023key ; 5sohn2024styledrop ; 151zhang2024generative . This approach utilizes an explicit text description to represent the SoI. For example, words such as cat could directly represent the user’s cat in the references. This setting typically requires fine-tuning the parameters of the diffusion model components such as UNet ronneberger2015unet or Transformer blocks vaswani2017attention . The plain text offers user-friendly prompt construction and injects subject prior information to ease tuning difficulty. However, this technique may over-specialize common terms, limiting their broader applicability as the model learns to associate general words with specific SoI characteristics.
Rare token 2ruiz2023dreambooth . This technique employs infrequently used tokens to minimize their impact on commonly used vocabulary. Similar to the plain text approach, the embeddings of rare tokens remain unchanged during fine-tuning. However, these rare tokens fail to provide useful subject prior information and still exhibit weak interference with unrelated vocabulary, potentially leading to ambiguity between the original meaning and the intended SoI reference.
3.1.3 Training Parameter Selection
The selection of trainable parameters represents a critical design consideration in PCS, directly impacting many key performance metrics: subject fidelity, training efficiency, and model storage requirements. Current tunable parameters can be broadly categorized into token embedding, existing model components, parameter-efficient techniques, and hybrid approaches.
Token embedding. As introduced in Sec. 3.1.2, token embedding optimization 1gal2022image introduces learnable tokens as the unique modifier to represent SoI through noise-to-image reconstruction. While achieving remarkable parameter efficiency, this approach faces challenges in detail preservation and prolonged training time due to the inherent compression of complex features into low-dimensional embeddings. Subsequent works 23voynov2023p+ ; 61alaluf2023neural ; 30zhang2023prospect ; 48wang2023hifi ; 66pang2023cross aim to address these limitations through different strategies, which are summarized in Sec. 5.1.
Existing model parameters. This paradigm directly optimizes pre-trained components, such as the text encoder, U-Net blocks, and transformer layers 2ruiz2023dreambooth ; 151zhang2024generative ; 6kumari2023multi ; 7tewel2023key ; 8han2023svdiff . Benefitting from the advanced representation capacity of these modules, the fine-tuning phase can achieve faster convergence (5-10 minutes) and superior visual fidelity compared to token-only methods, though at the cost of significant storage overhead. In addition, these modules inherently support attention mechanisms to facilitate feature enhancement operations.
Parameter-efficient extensions. Recent advanced methods have introduced parameter-efficient techniques into PCS, such as LoRA 46gu2024mix ; 58ruiz2023hyperdreambooth ; 115kong2024omg ; 51lu2024object ; 95achlioptas2023stellar ; 76yeh2023navigating and adapter modules 72xiang2023closer ; 5sohn2024styledrop , which inject small, trainable components into the base model. These methods achieve comparable performance to full parameter fine-tuning while dramatically reducing storage requirements.
Combined strategy. Since the aforementioned strategies are not conflicting, some methods allocate different learning rates and training phases to each component type to achieve optimal balance between fidelity and efficiency. For instance, the fine-tuned token embedding can be regarded as an effective initialization for the subsequent model weight fine-tuning 90hyung2023magicapture . Also, these two parts can be simultaneously optimized with different learning rates 74avrahami2023break ; 86hua2023dreamtuner .
3.1.4 Prompt Engineering
The construction of training prompts for samples typically starts with adding prefix words before the modifier token. A simplest example is “Photo of V*”. However, DreamBooth 2ruiz2023dreambooth noted that such a simple description causes a long training time and unsatisfactory performance. To address this, they incorporate the unique modifier with a class noun to describe the SoI in the references (e.g., “Photo of V* cat”). Also, the training caption for each training reference can be more precious for better disentanglement of SoI and irrelevant concepts 57he2023data , such as “Photo of V* cat on the chair”. This follows the tending that high-quality captions in the training set could assist in further improvement of accurate text control BetkerImprovingIG .
3.2 Pre-trained Adaptation (PTA) Framework
Figure 5: Illustration of the PTA method for personalized image synthesis. This framework utilizes a large-scale dataset to train a unified model that can process diverse personalization requests. The diffusion model is adapted to process hybrid inputs derived from both visual and textual features. Additionally, the concatenation of image and text features can be implemented in various ways, such as placeholder-based and reference-conditioned.
The PTA framework has emerged as a breakthrough approach for PCS, aiming to eliminate the computational burden of per-request fine-tuning while maintaining high-quality, subject-specific generation capabilities. To achieve this goal, this approach combines large-scale pretraining with reference-aware architectures to enable single-pass personalization, as shown in Fig. 5. Based on this architecture, three critical design factors are considered to ensure practical viability: 1) Preservation of semantic-critical features to guarantee visual consistency with reference inputs, 2) Effective fusion that combines reference features with text guidance to achieve desired generation, 3) Optimization of training dataset scale to achieve robust generalization without overfitting. In the following sub-section, we first present overall architecture introduction in Sec. 3.2.1 and then delve into three detailed factors.
3.2.1 Pre-trianed Adaption
Pre-training. During the pre-training phase, the PTA framework aims to establish direct mappings between reference characteristics (e.g., facial features, object textures) and synthesized outputs. To achieve this, the reference inputs are processed through dedicated feature extractors and fused with text prompts to serve as conditions to guide the generation, as shown in Fig. 5. A reconstruction loss is optimized that enforces the alignment between generated images and the large-scale dataset $X_{data}$:
$\mathcal{L}=\mathbb{E}_{\mathbf{x}_{0}\in\mathbf{X}_{data},t,\mathbf{\epsilon}}\left[\|\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta^{\prime}}(\mathbf{x}_{t},t,c^{\prime})\|_{2}^{2}\right],\quad c^{\prime}=\mathcal{F}(c,\mathbf{x}_{0})$ (9)
where $\mathcal{F}$ represents the fusion operation combining text condition $c$ and reference image $\mathbf{x}_{0}$. The tunable parameters $\theta^{\prime}$ include visual encoder weights, text encoder components, diffusion modules, and injected adapter modules.
Inference. During inference, the PTA framework processes a reference image through its visual encoder to extract discriminative features, which are then fused with text embeddings using the pretrained conditioning module. This fused representation guides the diffusion model to generate personalized outputs. This approach effectively eliminates test-time optimization to ensure fast generation.
3.2.2 Subject Feature Extraction
Extracting representative features of the SoI is crucial in the creation of personalized content. A common approach is to employ an encoder, leveraging pre-trained models such as CLIP radford2021learning and BLIP li2022blip . While these models excel at capturing global features, they often include irrelevant information that can detract from the fidelity, potentially compromising the quality of the personalized output, such as including the same background in the generation. To mitigate this issue, some studies incorporate additional prior knowledge to guide the learning process so as to focus on the targeted SoI. For instance, the SoI-specific mask 21zhang2023personalize ; 48wang2023hifi ; 74avrahami2023break ; 81safaee2023clic ; 51lu2024object ; 95achlioptas2023stellar contributes to the effective exclusion of the influence of the background. Moreover, using facial landmarks 37wang2024instantid in the context of human face customization helps improve identity preservation. We will discuss a more detailed technical summary of mask-assistant generation in Sec. 4.2.
Handling multiple input references presents another challenge but is essential for real-world deployment. This necessitates an ensemble of features from the multiple reference images to augment the framework’s adaptability. Yet, the majority of current PTA systems are limited to supporting one reference input. Some research works 37wang2024instantid ; 28li2023photomaker propose to average or stack features extracted from multiple references to form a composite SoI representation.
3.2.3 Subject Feature Fusion
Personalized content synthesis systems typically process two input modalities: reference images and textual descriptions, as illustrated in Fig. 5. Effective fusion of these heterogeneous features represents a critical technical challenge in PTA frameworks. Current methodologies can be categorized into four primary approaches:
Concatenation-based Fusion 19wei2023elite ; 28li2023photomaker ; 54xiao2023fastcomposer ; 89chen2023dreamidentity ; 85cai2023decoupled ; 96peng2023portraitbooth . This method makes the unique modifier a placeholder token to encapsulate visual subject characteristics. The placeholder token embedding, initialized with image features from visual encoders, is concatenated with text embeddings from the language model. This combined representation subsequently guides generation through standard cross-attention layers in the diffusion process, enabling basic subject-text alignment while maintaining architectural simplicity.
Cross-attention Fusion 40ye2023ip ; 11chen2024subject ; 37wang2024instantid ; 91li2023stylegan ; 212liu2024stylecrafter ; 221zhang2024ssr ; 220personalizationms . This paradigm extends the U-Net architecture with specialized attention mechanisms that jointly process visual and textual conditions. For example, IP-Adapter 40ye2023ip introduces decoupled cross-attention layers that maintain separate query projections for image and text features. In this case, the unique modifier directly displays the plain text.
Multimodal Encoder Fusion 10li2024blip ; 84zhou2023customization ; song2024moma ; 77pan2023kosmos ; 71ma2023unified . This approach leverages powerful multimodal encoder architectures (e.g., BLIP-2 li2023blip ) to jointly embed visual and textual subject descriptors. BLIP-Diffusion 10li2024blip exemplifies this strategy by learning a compact subject prompt embedding that fuses image patches with textual names through Q-Former modules.
Hybrid Fusion 16ma2023subject ; 156zhu2024multibooth . Moreover, some systems integrate multiple fusion strategies. For example, Subject-Diffusion 16ma2023subject integrate both concatenation and cross-attention fusion, taking advantage of the strengths of each approach to enhance the overall personalization capability.
3.2.4 Training Data
Training a PTA model for PCS necessitates a large-scale dataset. There are primarily two types of training samples utilized:
Triplet Data (Reference Image, Target Image, Target Caption). This dataset format is directly aligned with the PCS objectives, establishing a clear relation between the reference and the personalized content. However, such large-scale triplet samples are not widely available. Several strategies have been proposed to mitigate this issue: 1) Data Augmentation. Techniques such as foreground segmentation followed by placement in a different background are used to construct triplet data 10li2024blip . 2) Synthetic Sample Generation. Methods like SuTI 11chen2024subject utilize multiple TTF models to generate synthetic samples, which are then paired with original references. 3) Utilizing Recognizable SoIs. Collecting images of easily recognizable subjects, such as celebrities, significantly facilitates face personalization 60yuan2023inserting .
Dual Data (Reference Image, Reference Caption). This dataset is essentially a simplified version of the triplet format, where the personalized content is the original image itself. Such datasets are more accessible, including collections like LAION schuhmann2022laion and LAION-FACE zheng2022general . However, a notable drawback is that training tends to focus more on reconstructing the reference image rather than incorporating the text prompts. Consequently, models trained on this type of data might struggle with complex prompts that require substantial modifications or interactions with objects.
3.3 Hybrid Framework
Recently, some works have started to explore the combination of TTF and PTA methods. HyperDreamBooth 58ruiz2023hyperdreambooth states that PTA methods provide a general framework capable of handling a wide range of common objects, while TTF techniques enable fine-tuning to specific instances, which can improve the preservation of fine-grained details. They first develop a PTA network, followed by a subject-driven fine-tuning. Similarly, DreamTuner 86hua2023dreamtuner pre-trains a subject encoder that outputs diffusion conditions for accurate reconstruction. Then an additional fine-tuning stage is conducted for fine identity preservation. In contrast, SuTI 11chen2024subject first applies TTF methods to generate synthetic pairwise samples, which can be used for training the PTA network.
4 Techniques in Personalized Content Synthesis
Building upon the architectural framework discussed in Sec. 3, this section analyzes learning optimization techniques applicable to both frameworks. We focus on four categories: attention mechanisms, mask-guided generation, data augmentation, and regularization strategies. These methods aim to address key challenges in PCS, like enhancing subject fidelity, minimizing the interference of redundant semantics, enhancing generalization, and avoiding overfitting.
4.1 Attention-based Operation
Attention-based operations have become a crucial technique in model learning, particularly for processing features effectively hassanin2024visual . In diffusion models, these operations generally involve manipulating the way a model focuses on different parts of data, often through a method known as the Query-Key-Value (QKV) scheme. While large-scale pre-training has equipped this module with strong feature extraction capabilities, there remains significant ongoing work to enhance its performance for customization tasks.
Explicit attention weight manipulation. A cluster of studies focuses on restricting the influence of the SoI token within the attention layers. For example, Mix-of-Show 46gu2024mix designs region-aware cross-attention where the feature map is initially generated by the global prompt and replaced with distinct regional features corresponding to each entity. This avoids the misalignment between the words and visual regions. DreamTuner 86hua2023dreamtuner designs a self-subject-attention layer to further refine the subject identity. This attention module takes the features of the generated image as query, the concatenation of generated features as key, and reference features as value. Layout-Control 152chen2024training adjusts attention weights specifically around the layout without additional training. Cones 2 62liu2023cones also defines some negative attention areas to penalize the illegal occupation to allow multiple object generation. VICO 20hao2023vico inserts a new attention layer where a binary mask is deployed to selectively obscure the attention map between the noisy latent and the reference image features.
Implicit attention guidance. In addition to these explicit attention weights modification methods, many researchers 54xiao2023fastcomposer ; 8han2023svdiff ; 74avrahami2023break ; 16ma2023subject ; 81safaee2023clic ; 44rahman2024visual ; 90hyung2023magicapture employ localization supervision in the cross-attention module. Specifically, they train cross-attention modules using coordinate-aware loss functions, forcing attention maps to align with annotated subject positions. DreamTuner 86hua2023dreamtuner further refines this approach by designing an attention layer that effectively integrates features from different parts of the image.
4.2 Mask-guided Generation
Since reference images contain both the SoI and irrelevant visual elements, masks as a crucial prior indicating the position and contour of the specified object can effectively minimize the influence of redundant information.
Pixel-level mask. Benefitting from advanced segmentation methods like SAM kirillov2023segment , the SoI can be precisely isolated from the background. Based on this strategy, plenty of studies 21zhang2023personalize ; 48wang2023hifi ; 74avrahami2023break ; 81safaee2023clic ; 51lu2024object ; 95achlioptas2023stellar ; 90hyung2023magicapture ; 92tang2023retrieving ; 34li2023generate choose to discard the pixels of the background area so that the reconstruction loss can focus on the targeted object and exclude irrelevant disturbances. Another technique 85cai2023decoupled further adds the masked background reconstruction for better disentanglement. In addition, the layout indicated by the pixel mask can be incorporated into the attention modules as a supervision signal 54xiao2023fastcomposer ; 8han2023svdiff ; 74avrahami2023break ; 16ma2023subject ; 81safaee2023clic ; 44rahman2024visual ; 90hyung2023magicapture to adjust the attention’s concentration adaptively. Moreover, the mask can stitch specific feature maps to construct more informative semantic patterns 19wei2023elite ; 24jia2023taming ; 16ma2023subject .
Feature-level mask. In addition to the pixel-level manipulation, masks can be extended to feature-level operations. DisenBooth 18chen2023disenbooth defines an identity-irrelevant embedding with a learnable mask. By maximizing the cosine similarity between the identity-preservation embedding and identity-irrelevant embedding, the mask will adaptively exclude the redundant information, and thus the subject appearance can be better preserved. AnyDoor 14chen2023anydoor defines a high-frequency mask that stores detailed SoI features as a condition for the image generation process. Face-Diffuser 52wang2023high determines the mask through augmentation from the noise predicted by both a pre-trained text-to-image diffusion model and a PTA personalized model. Each model makes its own noise prediction, and the final noise output is a composite created through mask-guided concatenation.
4.3 Data Augmentation
Since the optimization of neural networks requires extensive data, existing PCS methods often struggle to capture complete semantic information of the SoI from limited references, resulting in poor images. To address this, various data augmentation strategies are employed to enrich the diversity of SoI references.
Compositional augmentation. Some methods enhance data diversity through classical image augmentation like blending and spatial rearrangement. SVDiff 8han2023svdiff manually constructs mixed images of multiple SoI as new training data, thereby enhancing the model’s exposure to complex scenarios. Such concept composition is also used in other works 16ma2023subject ; 138chen2023videodreamer ; 145wang2024customvideo . BLIP-Diffusion 10li2024blip segments the foreground subject and composes it in a random background so that the original text-image pairs are expanded to a larger dataset. Besides, StyleAdapter 26wang2023styleadapter chooses to shuffle the image patches to break the irrelevant subject and preserve the desired style. PACGen 34li2023generate shows that the spatial position also entangles with the identity information. Rescaling, center crop, and relocation are effective solutions for this issue.
Synthetic data. Generated synthetic data can provide a large amount of training resources when coupled with quality assurance mechanisms. SuTI 11chen2024subject establishes a cascaded pipeline where TTF models first generate diverse variations of each SoI. These synthetic samples then train the target PTA model. Similarly, DreamIdentity 89chen2023dreamidentity leverages the existing knowledge of celebrities embedded in large-scale pre-trained diffusion model to generate both the source image and the edited face image. StyleDrop 5sohn2024styledrop and GAL 151zhang2024generative implement iterative refinement pipelines where high-quality synthetic outputs from early training phases are incorporated into subsequent rounds.
External sources. It is intuitive to leveraging web resources to expand training datasets. COTI 3yang2023controllable adopts a scorer network to progressively expand the training set by selecting semantic-relevant samples with high aesthetic quality from a large web-crawled data pool.
4.4 Regularization
Regularization is an effective method that is used to regularize the weight update to avoid overfitting and enhance generalization.
Auxiliary data regularization. To mitigate the overfitting problem where the PCS system consistently produces identical outputs as the references, a commonly adopted method is to use an additional dataset composed of images with the same category of the SoI 2ruiz2023dreambooth . By reconstructing these images, the personalized model is required to generate diverse instances of the class while adapting to the target subject. Building on this strategy, StyleBoost 53park2023styleboost introduces an auxiliary style-specific auxiliary data to separate content and aesthetic adaptation. Later, a dataset 57he2023data is curated with detailed textual prompts (specifying attributes/contexts) to improve disentanglement between subject characteristics and background features.
Text embedding constraints. The semantic richness of pre-trained text (e.g., subject class name) provides a powerful regularization signal for personalized generation. By strategically constraining how subject-specific representations interact with established linguistic concepts in the embedding space, these approaches can achieve better generalization ability. For example, Perfusion 7tewel2023key constrains key projections toward class noun embeddings while learning value projections from subject images. Inspired by coached active learning wei2012coaching ; wei2011coached , which uses anchor concepts for optimization guidance, Compositional Inversion 9zhang2023compositional employs a set of semantically related tokens as anchors to constrain the token embedding search. In addition, some works 12gal2023encoder ; 51lu2024object regularize learnable token offsets relative to pre-trained CLIP embeddings. By minimizing the offset, the final word embedding is able to achieve better text alignment. Similarly, Cones 2 62liu2023cones minimizes the offset by reconstructing the features of 1,000 sentences containing the class noun. And 66pang2023cross optimizes the learnable token towards the mean textual embedding of 691 well-known names. Domain-Agnostic 17arar2023domain proposes to use a contrastive loss to guide the SoI text embedding close to its nearest CLIP tokens pre-trained on large-scale samples. On the other hand, VICO 20hao2023vico empirically finds that the end-of-text token <|EOT|> keeps the semantic consistency of SoI. To leverage this discovery, an L2 loss is leveraged to reduce the difference of attention similarity logits between the SoI token and <|EOT|>.
5 Categorization of Personalization Tasks
As shown in Fig. 6, personalization covers a range of areas, including objects, styles, faces, etc. The following subsections analyze these tasks through the frameworks established in Sec. 3, examining both TTF and PTA approaches for each domain.
Figure 6: Visualization of several personalized content synthesis tasks. Given text prompts and a few images of the subject of interest, personalization approaches are required to produce expected images.
5.1 Personalized Object Generation
As the foundational task, personalized object generation requires learning discriminative features from general instances (e.g., toys, vehicles, or architecture) and rendering them in novel contexts specified by textual prompts.
TTF framework. TTF methods perform instance-specific optimization during inference by fine-tuning model parameters on reference images. This achieves superior subject fidelity and handles rare attributes effectively.
An important branch in TTF methods is to optimize the learnable token embeddings. The first work started from Textual Inversion 1gal2022image , which applies a simple yet effective method that introduces the unique modifier as a new token to represent the SoI. One of the significant benefits of this method is its minimal storage requirement as the new tokens consumes just a few kilobytes. However, the method compresses complex visual features into a small set of parameters, which can lead to long convergence times and a potential loss in visual fidelity. Recent work aims to address these limitations through several directions. DVAR 69voronov2024loss improves training efficiency by proposing a clear stopping criterion by removing all randomness to indicate the convergence. For enhanced representation, P+ 23voynov2023p+ introduces distinct learnable tokens across different layers of the U-Net architecture, thereby offering better attribute control through additional learnable parameters. NeTI 61alaluf2023neural advances this concept by proposing a neural mapper that adaptively outputs token embeddings based on the denoising timestep and specific U-Net layers. ProSpect 30zhang2023prospect recommends optimizing multiple token embeddings tailored to different denoising timesteps based on the observation that different types of prompts, like layout, color, structure, and texture, are activated at different stages of the denoising process. Similarly, a study by 79agarwal2023image shows layered activation insight to learn distinct attributes by selectively activating the tokens within their respective scopes. Later, HiFiTuner 48wang2023hifi integrates multiple techniques into the learnable token, including mask-guided loss function, parameter regularization, time-dependent embedding, and generation refinement assisted by the nearest reference. Alternative approaches, like DreamArtist 68dong2022dreamartist , choose to optimize both negative and positive prompt embeddings to refine the detail preservation. In addition to these token-level refinement approaches, the field continues to evolve with novel techniques such as InstructBooth’s reinforcement learning framework 56chae2023instructbooth and gradient-free evolutionary optimization 73fei2023gradient . In summary, recent developments following the foundational work of Textual Inversion focus on reducing training times and enhancing the visual quality of generated images.
In the realm of TTF methods for PCS, there is a clear shift towards fine-tuning model weights rather than just the token embeddings. This approach often addresses the limitations where token embeddings alone struggle to capture complex semantics uncovered in the pre-training data 46gu2024mix ; 80zhao2023catversion . DreamBooth 2ruiz2023dreambooth proposes to use a unique modifier by a rare token to represent the SoI and fine-tune the whole parameters of the diffusion model. Besides, a regularization dataset containing 20-30 images with the same category as SoI is adopted to overcome the overfitting problem. These two combined approaches achieve impressive performance that largely promotes the progress of the research on image personalization. However, fine-tuning the entire model for each new object causes considerable storage costs, potentially rendering widespread application. To address this, Custom Diffusion 6kumari2023multi focuses on identifying and fine-tuning critical parameters, particularly the key-value projections in cross-attention layers, to achieve a balance of visual fidelity and storage efficiency. Further approach, Perfusion 7tewel2023key , also adopts the cross-attention fine-tuning and proposes to regularize the update direction of the K (key) projection towards the super-category token embedding and the V (value) projection towards the learnable token embedding. COMCAT 75xiao2023comcat introduces a low-rank approximation of attention matrices, which drastically reduces storage requirements to 6 MB while maintaining high fidelity in the outputs. Additionally, methods like adapters 72xiang2023closer ; 5sohn2024styledrop and LoRA variants 46gu2024mix ; 58ruiz2023hyperdreambooth ; 115kong2024omg ; 51lu2024object ; 95achlioptas2023stellar ; 76yeh2023navigating are increasingly utilized in personalized generation for parameter-efficient fine-tuning. It is worth noting that the token embedding fine-tuning is compatible with diffusion weight fine-tuning. Multiple methods 90hyung2023magicapture ; 74avrahami2023break ; 86hua2023dreamtuner have started using combined weight fine-tuning.
PTA framework. For practical deployment of PCS systems, fast response time is a crucial factor. PTA methods enable real-time generation (less than 10 sec/subject) by leveraging large-scale pre-training to avoid per-subject optimization during the inference phase. Re-Imagen 67chen2022re introduces a retrieval-augmented generative approach, which leverages features from text-image pairs retrieved via a specific prompt. While it is not specifically tailored for object personalization, it demonstrates the feasibility of training reference-conditioned frameworks. Later, ELITE 19wei2023elite specifically targets image personalization by combining the global reference features with text embedding while incorporating local features that exclude irrelevant backgrounds. Both fused features and local features serve as conditions for the denoising process. Similarly, InstantBooth 22shi2023instantbooth retrains CLIP models to extract image features and patch features, which are injected into the diffusion model via the attention mechanism and learnable adapter, respectively. Additionally, UMM-Diffusion 71ma2023unified designs a multi-modal encoder that produces fused features based on the reference image and text prompt. The text features and multi-modal hidden state are seen as guidance signals to predict a mixed noise. Another work, SuTI 11chen2024subject , adopts the same architecture as Re-Imagen. The difference lies in the training samples which are produced by a massive number of TTF models, each tuned on a particular subject set. This strategy promotes a more precise alignment with personalization at an instance level rather than the class level of Re-Imagen. Moreover, Domain-Agnostic 17arar2023domain combines a contrastive-based regularization technique to push the pseudo embedding produced by the image encoder towards the existing nearest pre-trained token. Besides, They introduce a dual-path attention module separately conditioned on the nearest token and pseudo embedding. Compared to the methods that use separate encoders to process a single modality, some works have explored the usage of pre-trained multi-modal large language models (MLLM) that can process text and image modality within a unified framework. For example, BLIP-Diffusion 10li2024blip utilizes the pre-trained BLIP2 li2023blip that encodes multimodal inputs including the SoI reference and a class noun. The output embedding is then concatenated with context description and serves as a condition to generate images. Further, Customization Assistant 84zhou2023customization and KOSMOS-G 77pan2023kosmos replace the text encoder of Stable Diffusion with a pre-trained MLLM to output a fused feature based on the reference and context description. Meanwhile, to meet the standard format of Stable Diffusion, a network is trained to align the dimension of the output embedding.
5.2 Personalized Style Generation
Personalized style generation seeks to tailor the aesthetic elements of reference images. The concept of “style” now includes a wide range of artistic elements, such as brush strokes, material textures, color schemes, structural forms, lighting techniques, and cultural influences.
TTF framework. In this field, StyleDrop 5sohn2024styledrop leverages adapter tuning to efficiently capture the style from a single reference image. This method demonstrates the effectiveness through iterative training, utilizing synthesized images refined by feedback mechanisms, like human evaluations and CLIP scores. This approach not only enhances style learning but also ensures that the generated styles align closely with human aesthetic judgments. Later, GAL 151zhang2024generative proposes an uncertainty-based evaluation strategy to filter high-quality synthetic-style data and uses a weighted schema to balance the contribution of the additional samples and the original reference. Furthermore, StyleAligned 98hertz2023style focuses on maintaining stylistic consistency across a batch of images. This is achieved by using the first image as a reference, which acts as an additional key and value in the self-attention layers, ensuring that all subsequent images in the batch adhere to the same stylistic guidelines. Style-friendly 211choi2024style introduces a novel diffusion model fine-tuning approach that enhances personalized artistic style generation by adaptively biasing noise sampling toward higher noise levels, where stylistic features emerge.
PTA framework. For the PTA framework, StyleAdapter 26wang2023styleadapter employs a dual-path cross-attention mechanism within the PTA framework. This model introduces a specialized embedding module designed to extract and integrate global features from multiple style references. Diptych Prompting 209shin2024large utilizes an inpainting mechanism to draw another image with the same style of the reference part.
5.3 Personalized Face Generation
Personalized face generation aims to generate diverse ID images that adhere to text prompt specifications, utilizing only a few initial face images. Compared to general object personalization, the scope is narrowed to a specific class, humans. It is easy to acquire large-scale human-centric datasets zheng2022general ; yang2020facescape ; cao2018vggface2 and utilize pre-trained models in well-developed areas, like face landmark detection wu2019facial ; khabarlak2021fast and face recognition adjabi2020past .
TTF framework. As for the TTF methods, PromptNet 55zhou2023enhancing trains a diffusion-based network that encodes the input image and noisy latent to a word embedding. To alleviate the overfitting problem, the noises predicted by the word embedding and context description are balanced through fusion sampling through classifier-free guidance. HyperDreamBooth 58ruiz2023hyperdreambooth proposes a second-time fine-tuning strategy to further enhance the fidelity after training a PTA model on a large-scale dataset. Additionally, Celeb Basis 60yuan2023inserting provides a novel idea that the personalized ID can be viewed as the composition of celebrity’s face which has been learned by the pre-trained diffusion model. Based on this hypothesis, a simple MLP is optimized to transform face features into the celeb embedding space.
PTA framework. Thanks to the abundance of available datasets featuring the same individual in various contexts, which provide valuable data for pre-training PTA methods, the number of works in PTA frameworks is rapidly increasing. Face0 88valevski2023face0 crops the face region to extract refined embeddings and concatenates them with text features. During the sampling phase, the output of classifier-free guidance is replaced by a weighted combination of the noise patterns predicted by face-only embedding, text-only embedding, and concatenated face-text embedding. The $\mathcal{W}+$ Adapter 91li2023stylegan constructs a mapping network and residual cross-attention modules to transform the facial features from the StyleGAN karras2019style $\mathcal{W}+$ space into the text embedding space of Stable Diffusion. FaceStudio 31yan2023facestudio adapts the cross-attention layer to support hybrid guidance including stylized images, facial images, and textual prompts. Moreover, PhotoMaker 28li2023photomaker constructs a high-quality dataset through a meticulous data collection and filtering pipeline. They use a two-layer MLP to fuse ID features and class embeddings for an overall representation of human portrait. PortraitBooth 96peng2023portraitbooth also employs a simple MLP, which fuses the text condition and shallow features of a pre-trained face recognition model. To ensure expression manipulation and facial fidelity, they add another expression token and incorporate the identity preservation loss and mask-based cross-attention loss. InstantID 37wang2024instantid additionally introduces a variant of ControlNet that takes facial landmarks as input, providing stronger guiding signals compared to the methods that solely rely on attention fusion.
5.4 Multiple Subject Composition
Multiple subject composition refers to the scenario where users intend to compose multiple SoI displayed in one or more references.
TTF framework. This task presents a challenge for the TTF methods, particularly in how to integrate the parameters within the same module which are separately fine-tuned for individual SoI. Some works focus on the one-for-one generation following a fusion mechanism. For instance, Custom Diffusion 6kumari2023multi proposes a constrained optimization method to merge the cross-attention key-value projection weights with the goal of maximizing reconstruction performance for each subject. Mix-of-Show 46gu2024mix fuses the LoRA hu2021lora weights with the same optimization objective. StyleDrop 5sohn2024styledrop dynamically summarizes noise predictions from each personalized diffusion model. In OMG 115kong2024omg , the latent predicted by each LoRA-tuned model is spatially composited using the subject mask. Joint training is another strategy to cover all expected subjects. SVDiff 8han2023svdiff employs a data augmentation method called Cut-Mix to compose several subjects together and applies a location loss to regularize attention maps, ensuring alignment between each subject and its corresponding token. Similar strategies are found in other works 6kumari2023multi ; 60yuan2023inserting which train a single model by reconstructing the appearance of every SoI. There are advanced control mechanisms designed to manage multiple subjects. Cones 4liu2023cones proposes to find a small cluster of neurons that preserve the most information about SoI. The neurons belonging to different SoI will be simultaneously activated to generate the combination. Compositional Inversion 9zhang2023compositional introduces spatial region assignment to different subjects to improve the composition success rate.
PTA framework. For PTA frameworks, multi-subject generation is achieved through specialized architectural designs. Fastcomposer 54xiao2023fastcomposer , Subject-Diffusion 16ma2023subject , and $\lambda$-eclipse 222patel2024lambda place each subject feature in its corresponding placeholder in the text embedding, ensuring a seamless and efficient combination. CustomNet 29yuan2023customnet and MIGC 174zhou2024migc train a PTA network that supports location control for each subject. SSR-Encoder 221zhang2024ssr implements an encoder to selectively preserve the desired subject feature and a cross-attention module to support multi-subject feature fusion.
5.5 High-level Semantic Personalization
Recently, the field of image personalization has started to include complex semantic relationships and high-level concepts. Different approaches have been developed to enhance the capability of models to understand and manipulate these abstract elements.
TTF framework. For now, all researches in this field are based on the TTF framework. ReVersion 63huang2023reversion intends to invert object relations from references. Specifically, they use a contrastive loss to guide the optimization of the token embedding towards specific clusters of Part-of-Speech tags, such as prepositions, nouns, and verbs. Meanwhile, they also increase the likelihood of adding noise at larger timesteps during the training process to emphasize the extraction of high-level semantic features. Lego 45motamed2023lego focuses on more general concepts, such as adjectives, which are frequently intertwined with the subject appearance. The concept can be learned from a contrastive loss applied to the dataset comprising clean subject images and images that embody the desired adjectives. Moreover, ADI 43huang2023learning aims to learn the action-specific identifier from the references. To ensure the inversion only focuses on the desired action, ADI extracts gradient invariance from a constructed triplet sample and applies a threshold to mask out the irrelevant feature channels.
6 Extensions of Personalized Content Synthesis
While the core PCS system focuses on image generation from reference subjects, recent advances have expanded its capabilities across multiple dimensions. This section examines several frontier extensions that push the boundaries of personalization technology.
6.1 Personalization on Extra Conditions
Recent personalization tasks tend to include additional conditions for diverse content customization. One common application is to condition on a fixed source image. For example, PhotoSwap 15gu2024photoswap introduces a new task that replaces the subject in the source image with the SoI from the reference image. To address this requirement, they first fine-tunes a diffusion model on the references to obtain a personalized model. To preserve the background of the source image, they initialize the noise with DDIM inversion song2020denoising and replace the intermediate feature maps with those derived from source image generation during inference time. Later, MagiCapture 91li2023stylegan broadens the scope to face customization. Another similar application can be found in Virtual Try-on, which aims to fit selected clothing onto a target person. The complexities of this task have been thoroughly analyzed in another survey islam2024deep .
Additional conditions in personalization tasks may include adjusting the layout 152chen2024training , transforming sketches 94chen2023democaricature , controlling viewpoint 29yuan2023customnet ; 163kumari2024customizing , or modifying poses 37wang2024instantid . Each of these conditions presents unique challenges and requires specialized approaches to integrate these elements seamlessly into the personalized content.
6.2 Personalized Video Generation
With the rising popularity of video generation xing2024survey , video personalization has also begun to attract attention. In video personalization, the SoI can be classified into three distinct categories: appearance, motion, and the combination of both appearance and motion.
Appearance-based video personalization. This task focuses on transferring subject appearance from static images to video sequences. The standard TTF pipeline utilizes reference images as appearance anchors and fine-tunes video diffusion models (VDM) for temporal synthesis. The process involves leveraging sophisticated methods from 2D personalization, such as parameter-efficient fine-tuning 145wang2024customvideo , data augmentation 138chen2023videodreamer ; 145wang2024customvideo , and attention manipulation 83jiang2023videobooth ; 212liu2024stylecrafter ; 145wang2024customvideo . Additionally, several studies 82zhao2023videoassembler ; 83jiang2023videobooth ; 212liu2024stylecrafter ; 105feng2023dreamoving have explored the PTA framework. These diffusion models are specifically tailored to synthesize videos based on the image references.
Motion-based video personalization. In this task, the reference input switches to a video clip containing a consistent action. A common approach is to fine-tune the video diffusion model by reconstructing the action clip 150wu2023tune ; 103song2023save ; 140jeong2023vmc ; 148esser2023structure ; 136he2023animate ; 102wu2023lamp ; 104materzynska2023customizing . However, distinguishing between appearance and motion within the reference video can be challenging. To solve this problem, SAVE 103song2023save applies appearance learning to ensure that appearance is excluded from the motion learning phase. Additionally, VMC 140jeong2023vmc removes the background information during training prompt construction.
Appearance and motion personalization. When integrating both subject appearance and motion, innovative methods are employed to address the complexities of learning both aspects simultaneously. MotionDirector 101zhao2023motiondirector utilizes spatial and temporal losses to facilitate learning across these dimensions. Another approach, DreamVideo 142wei2023dreamvideo , incorporates residual features from a randomly selected frame to emphasize subject information. This technique enables the fine-tuned module to primarily focus on learning motion dynamics.
In summary, video personalization strategies vary significantly based on the specific aspects. Moreover, due to the current limitations in robust video feature representation, PTA video personalization that is directly conditioned on video input remains an area under exploration.
6.3 Personalized 3D generation
Personalized 3D Generation refers to the process of creating customized 3D models or scenes based on 2D SoI images. Basically, the pipeline begins by fine-tuning a 2D diffusion model using TTF methods. This tuned model then guides the optimization of a 3D Neural Radiance Field (NeRF) model mildenhall2021nerf for each specific prompt 120lin2023magic3d ; 126shi2023mvdream . In the second phase, DreamFusion poole2022dreamfusion is the key technique, which introduces Score Distillation Sampling (SDS) to train a 3D model capable of rendering images aligned with 2D diffusion model. Building on this foundation, several methods have been developed to improve the workflow. DreamBooth3D 121raj2023dreambooth3d structures the process into three phases: initializing and optimizing a NeRF from a DreamBooth model, rendering multi-view images, and fine-tuning a secondary DreamBooth for the final 3D NeRF refinement. Consist3D 127ouyang2023chasing enhances text embeddings by training two distinct tokens, a semantic and a geometric token, during 3D model optimization. TextureDreamer 133yeh2024texturedreamer focuses on extracting texture maps from optimized spatially-varying bidirectional reflectance distribution (BRDF) fields for rendering texture on wide-range 3D subjects.
Additionally, advancements extend to 3D avatar rendering and dynamic scenes. Animate124 128zhao2023animate124 and Dream-in-4D 130zheng2023unified integrate video diffusion for 4D dynamic scene support within the 3D optimization process. In avatar rendering, PAS 122azadi2023text generates 3D body poses configurable by avatar settings, StyleAvatar3D 123zhang2023styleavatar3d facilitates 3D avatar generation based on images, and AvatarBooth 124zeng2023avatarbooth employs dual fine-tuned diffusion models for separate face and body generation.
6.4 Attack and Defense
This rapid advancement raises concerns about the ethical implications of PCS, particularly in areas such as misinformation, privacy violations, and the creation of deepfakes. There is an increased risk that individuals or organizations may exploit them to produce misleading content or manipulate public perception. To mitigate this, Anti-DreamBooth 13van2023anti aims to add a subtle noise perturbation to the references so that any personalized model trained on these samples only produces terrible results. The basic idea is to maximize the reconstruction loss of the surrogate model. Additionally, 50zhang2023backdooring suggests to predefine a collection of trigger words and meaningless images. These data are paired and incorporated during the training phase. Once the trigger words are encountered, the synthesized image will be intentionally altered for safeguarding.
6.5 Other Emerging Directions
Several works are exploring different personalization extensions. For example, SVG personalization is introduced by 36zhang2023text , in which a parameter-efficient fine-tuning method is applied to create SVGs. After first-step generation, the SVGs are refined through a process that includes semantic alignment and a dual optimization approach, which utilizes both image-level and vector-level losses to enhance the final output. Another application, 360-degree panorama customization 25wang2024customizing , is also emerging as a potential tool for personalization in the digital imaging realm.
7 Evaluation
7.1 Evaluation Dataset
To assess the performance of personalized models, various datasets have been developed:
DreamBench 2ruiz2023dreambooth serves as the primary evaluation benchmark for DreamBooth 2ruiz2023dreambooth , containing 30 diverse subjects (e.g., backpacks, animals, vehicles, and toys) with 25 unique prompts per subject.
DreamBench-v2 11chen2024subject expands the evaluation scope of DreamBench by adding 220 test prompts for each subject.
Custom-10 6kumari2023multi used in Custom Diffusion 6kumari2023multi evaluates 10 subjects, each with 20 specific test prompts, and includes tests for multi-subject composition with 5 pairs of subjects and 8 prompts for each pair.
Custom-101 6kumari2023multi is the latest released dataset by the authors of Custom Diffusion 6kumari2023multi , which comprises 101 subjects to provide a broader scope of evaluation.
Stellar 95achlioptas2023stellar specifically targets human-centric evaluation, featuring 20,000 prompts on 400 human identities.
Despite these contributions, they remain fragmented across different research groups and the research community still lacks a benchmark tested on a large number of personalized generation tasks. To address this gap, this survey introduces a comprehensive evaluation dataset [^1] This dataset is available on: https://github.com/zhangxulu1996/awesome-personalization designed for the most common personalized object and face personalization. Section 7.3 details our benchmark design and evaluation results.
7.2 Evaluation Metrics
As PCS aims to maintain fidelity to the SoI while ensuring alignment with textual conditions, the metrics are designed from two aspects, text alignment and visual fidelity.
The text alignment metrics quantify how precisely generated outputs reflect prompt semantics:
CLIP-T measures semantic alignment using the cosine similarity between CLIP radford2021learning embeddings of generated images and their text prompts.
ImageReward xu2023imagereward , HPS Score (v1/v2) wu2023human ; wu2023humanv2 , and PickScore kirstain2023pick employ learned models trained on human judgments to better correlate with perceptual quality to reflect human preference better.
To determine how closely the generated subject resembles the SoI, visual fidelity can be assessed via the following metrics:
CLIP-I evaluates subject preservation through CLIP image embedding similarity between generations and references. Optimal values balance fidelity (high scores) against overfitting (excessive scores that ignore text guidance).
DINO-I caron2021emerging provides complementary assessment using DINO’s instance-aware features, particularly effective for object-level similarity.
Fréchet Inception Distance (FID) FIDheusel2017gans quantifies the statistical similarity between generated and real image distributions through Inception-V3.
Inception Score (IS) ISsalimans2016improved evaluates both image quality and diversity through classifier confidence.
In addition to these commonly adopted metrics, there are some discussions of specialized metrics for PCS system evaluation: LyCORIS 76yeh2023navigating introduces a 5-dimensional assessment covering fidelity, controllability, diversity, base model preservation, and image quality. Stellar 95achlioptas2023stellar develops six human-centric metrics including soft-penalized CLIP text score, Identity Preservation Score, Attribute Preservation Score, Stability of Identity Score, Grounding Objects Accuracy, and Relation Fidelity Score. These metrics ensure a structured and detailed evaluation of personalized models. This evolving metric landscape reflects the growing sophistication of PCS systems, with optimal evaluation typically requiring combinations of general and task-specific measures.
7.3 New Benchmark on SoTA Methods
Although multiple evaluation datasets have been proposed in this area, there is still an urgent need for a standardized benchmark that systematically assesses performance across diverse PCS methodologies.To address this limitation, we present a comprehensive evaluation built upon existing works.
For Object. We collect a dataset of 47 subjects from available resources 1gal2022image ; 2ruiz2023dreambooth ; 6kumari2023multi . Following the methodology of DreamBooth 2ruiz2023dreambooth , we categorize the subjects into two classes: objects and live pets, based on whether the subject is a living entity. Specifically, 10 out of 47 subjects are pets and the remaining 37 are various objects. To evaluate the performance, we utilize the text prompts from DreamBooth 2ruiz2023dreambooth . This includes 20 recontextualization prompts and 5 property modification prompts for objects, along with 10 recontextualization, 10 accessorization, and 5 property modification prompts for pets, totally 25 prompts per category.
For Face. We collect 15 subjects from Celeb-A liu2018large dataset. We use 40 prompts for evaluation, including 10 accessory prompts, 10 style prompts, 10 action prompts, and 10 context prompts.
Settings. To assess existing representative PCS methods using our constructed test dataset, we compute CLIP-T to assess the text alignment. We calculate CLIP-I for subject fidelity in object generation. For subject fidelity in face generation, we detect faces in both the generated and target images using MTCNN zhang2016mtcnn and calculate the pairwise identity similarity using FaceNet schroff2015facenet . To uniformly evaluate and compare existing personalized generation methods, we select 22 representative PCS methods for evaluation. We generate 4 images for each test prompt and set the same random seed for all methods.
Results. The evaluation results are shown in Tab.1. It is evident that no method excels simultaneously in both visual fidelity and text alignment metrics. This highlights a significant challenge currently faced by PCS methods: achieving the optimal trade-off between subject preservation and editability. Striking this balance is difficult, as high subject fidelity often comes at the cost of prompt fidelity and vice versa. Furthermore, we note that higher visual fidelity does not always equate to better performance. The generated images sometimes exhibit patterns that closely mirror the reference images and ignore the prompt guidance. This phenomenon primarily arises from model overfitting on the reference input, which hinders the model’s ability to generalize. Consequently, the visual similarity metric yields higher results based on the mirrored output and the reference, rather than providing an accurate representation of the performance.
Table 1: Evaluation results of representative methods on our evaluation dataset.
Type Methods Framework Backbone CLIP-T CLIP-I
Object Textual Inversion 1gal2022image TTF SD 1.5 0.199 0.749
Dreambooth 2ruiz2023dreambooth TTF SD 1.5 0.286 0.772
P+ 23voynov2023p+ TTF SD 1.4 0.244 0.643
Custom Diffusion 6kumari2023multi TTF SD 1.4 0.307 0.722
NeTI 61alaluf2023neural TTF SD 1.4 0.283 0.801
SVDiff 8han2023svdiff TTF SD 1.5 0.282 0.776
Perfusion 7tewel2023key TTF SD 1.5 0.273 0.691
ELITE 19wei2023elite PTA SD 1.4 0.292 0.765
BLIP-Diffusion 10li2024blip PTA SD 1.5 0.292 0.772
IP-Adapter 40ye2023ip PTA SD 1.5 0.2722 0.825
SSR Encoder 221zhang2024ssr PTA SD 1.5 0.288 0.792
MoMA song2024moma PTA SD 1.5 0.322 0.748
Diptych Prompting 209shin2024large PTA FLUX 1.0 dev 0.327 0.722
$\lambda$-eclipse 222patel2024lambda PTA Kandinsky 2.2 0.272 0.824
MS-Diffusion 220personalizationms PTA SDXL 0.298 0.777
Face CrossInitialization 66pang2023cross TTF SD 2.1 0.261 0.469
Face2Diffusion 170shiohara2024face2diffusion PTA SD 1.4 0.265 0.588
SSR Encoder 221zhang2024ssr PTA SD 1.5 0.233 0.490
FastComposer 54xiao2023fastcomposer PTA SD 1.5 0.230 0.516
IP-Adapter 40ye2023ip PTA SD 1.5 0.292 0.462
IP-Adapter 40ye2023ip PTA SDXL 0.292 0.642
PhotoMaker 28li2023photomaker PTA SDXL 0.311 0.547
InstantID 37wang2024instantid PTA SDXL 0.278 0.707
8 Challenge and Outlook
8.1 Overfitting Problem
As discussed in Sec. 7.3, current PCS systems face a critical challenge of overfitting because of a limited set of reference images. This overfitting problem manifests in two ways: 1) Loss of SoI editability. The personalized model tends to produce images that rigidly mirror the SoI in the reference, such as consistently depicting a cat in an identical pose. 2) Irrelevant semantic inclusion. The irrelevant elements in the references are generated in the output, such as backgrounds or objects not pertinent to the current context.
To investigate the rationale behind, Compositional Inversion 9zhang2023compositional observes that the learned token embedding is located in an out-of-distribution area compared to the center distribution formed by pre-trained words. This is also found in another work 66pang2023cross that the learnable token embeddings deviate significantly from the distribution of the initial embedding. In addition, there is evidence 9zhang2023compositional ; 24jia2023taming ; 20hao2023vico suggesting that the unique modifier dominates in the cross-attention layers compared to the other context tokens, leading to the absence of other semantic appearances.
To address this issue, many solutions have been proposed. Most methods discussed in Sec. 4 contribute to the alleviation of the overfitting problem, such as the exclusion of redundant background, attention manipulation, regularization of the learnable parameters, and data augmentation. However, it has not been solved yet, especially in the cases where the SoI has a non-rigid appearance 9zhang2023compositional or the context prompt has a similar semantic correlation with the irrelevant elements in the reference 151zhang2024generative . It is clear that addressing overfitting in PCS is not merely a technical challenge but a necessity for ensuring the practical deployment and scalability of these systems in varied and dynamic real-world environments. Therefore, an effective strategy and robust evaluation metrics are urgently needed to achieve broader adoption and greater satisfaction in practical applications.
8.2 Trade-off on Subject Fidelity and Text Alignment
The ultimate goal of personalized content synthesis is to create systems that not only render the SoI with high fidelity but also effectively respond to textual prompts. However, achieving excellence in both areas simultaneously presents a notable conflict. Specifically, high subject fidelity typically involves capturing and reproducing detailed and specific features of the SoI. This often requires the model to minimize the reconstruction loss to replicate delicate characteristics accurately. Conversely, text alignment necessitates that the system flexibly adapts the SoI according to varying textual descriptions. These descriptions may suggest changes in pose, expression, environment, or stylistic alterations that do not aim to reconstruct the exact visualization in the reference. As a result, it becomes challenging to achieve flexible adaptation in different contexts while simultaneously pushing the model to capture fine-grained details. To address this inherent conflict, Perfusion 7tewel2023key proposes to regularize the attention projections by these two items. 35cao2023concept decouples the conditional guidance into two separate processes, which allows for the distinct handling of subject fidelity and textual alignment. Despite these efforts, there still remains room for further exploration and refinement of this issue. Enhanced model architectures, innovative training methodologies, and more dynamic data handling strategies could potentially provide new pathways to better balance the demands of subject and text fidelity in PCS systems.
8.3 Standardization and Evaluation
Despite the popularity of personalization, there is a noticeable lack of standardized test datasets and robust evaluation metrics that accurately capture the performance of different strategies. Currently, a widely used metric for assessing visual fidelity relies on CLIP image similarity. However, this approach may wrongly exaggerate the value when the model is overfitted to the references. Therefore, future efforts should focus on creating comprehensive and widely accepted benchmarks that can evaluate various aspects of PCS models, including but not limited to visual fidelity and subject editability.
8.4 Multimodal Autoregressive Frameworks
Recent advancements in multimodal autoregressive models present novel solutions for PCS by unifying cross-modal understanding and generation. Models like Emu3 wang2024emu3 demonstrate that autoregressive architectures can natively handle image-text-video sequences through discrete tokenization and joint transformer training. This paradigm enables seamless integration of user-provided multimodal references (e.g., text descriptions paired with SoI images) while maintaining contextual coherence across generation steps. Besides, this framework natively support for subject editing via multi-round chat, effectively addressing the overfitting limitations observed in diffusion-based PCS.
8.5 Interactive Personalization Workflow
The evolution of interactive generation systems presents valuable opportunities for PCS, particularly in addressing core challenges like the overfitting problem while enhancing user control. Recent advancements in conversational PCS systems, such as Gemini-2.0-flash’s interactive image editing team2023gemini , demonstrate how iterative refinement through natural language can optimize both subject fidelity and prompt alignment. This aligns with PCS’s fundamental goal of balancing precise subject representation with flexible editing capability.
9 Conclusion
This survey has provided a thorough review of personalized content synthesis with diffusion models, particularly focusing on 2D image customization. We explore two main frameworks, TTF and PTA methods, and delve into their mechanics. We also cover the recent progress in specific customization areas, including object, face, style, video, 3D synthesis. In addition to the impressive techniques, we propose several challenges that still need to be addressed. These include preventing overfitting, finding the right balance between reconstruction quality and editability, and standardizing evaluation methods. To support ongoing research, we collect a test dataset from existing literature and evaluate the classical method to provide a clear comparison. By providing detailed analysis and outlining targeted recommendations, we hope to promote further innovation and collaboration within the PCS community.
Table 2: Paper summary on personalization.
Paper Scope Framework Backbone Technique
Textual Inversion 1gal2022image Object TTF LDM Token Embedding Fine-tuning
DreamBooth 2ruiz2023dreambooth Object TTF Imagen Diffusion Model Fine-tuning; Regularization Dataset
DreamArtist 68dong2022dreamartist Object TTF LDM Negative Prompt Fine-tuning
DVAR 69voronov2024loss Object TTF SD 1.5 Randomness Erasing
HiPer 33han2023highly Object TTF SD Token Embedding Enhancement
P+ 23voynov2023p+ Object TTF SD 1.4 Token Embedding Enhancement
Unet-finetune 72xiang2023closer Object TTF SD Parameter-efficient Fine-tuning
Jia et al. 24jia2023taming Object TTF Imagen Mask-assisted Generation; Regularization
COTI 3yang2023controllable Object TTF SD 2.0 Data Augmentation
Gradient-Free TI 73fei2023gradient Object TTF SD Evolutionary Strategy
PerSAM 21zhang2023personalize Object TTF SD Mask-assisted Generation
DisenBooth 18chen2023disenbooth Object TTF SD 2.1 Regularization
NeTI 61alaluf2023neural Object TTF SD 1.4 Token Embedding Enhancement
Prospect 30zhang2023prospect Object TTF SD 1.4 Token Embedding Enhancement
Break-a-Scene 74avrahami2023break Object TTF SD 2.1 Attention-based Operation; Attention-based Operation; Mask-assisted Generation
COMCAT 75xiao2023comcat Object TTF SD 1.4 Attention-based Operation
ViCo 20hao2023vico Object TTF SD Attention-based Operation; Regularization
OFT 59qiu2024controlling Object TTF SD 1.5 Finetuning Strategy
LyCORIS 76yeh2023navigating Object TTF SD 1.5 Attention-based Operation
He et al. 57he2023data Object TTF SD, SDXL Data Augmentation
DIFFNAT 78roy2023diffnat Object TTF SD Regularization
MATTE 79agarwal2023image Object; Style; High-level Semantic TTF SD Regularization
LEGO 45motamed2023lego Object TTF SD Regularization
Catversion 80zhao2023catversion Object TTF SD 1.5 Embedding Concatenation
CLiC 81safaee2023clic Object TTF SD 1.4 Attention-based Operation; Mask-assisted Generation
HiFi Tuner 48wang2023hifi Object TTF SD 1.4 Attention-based Operation; Mask-assisted Generation; Regularization
InstructBooth 56chae2023instructbooth Object TTF SD 1.5 Reinforcement Learning
DETEX 85cai2023decoupled Object TTF SD 1.4 Mask-assisted Generation
DreamDistribution 114zhao2023dreamdistribution Object TTF SD 2.1 Token Embedding Enhancement
DreamTuner 86hua2023dreamtuner Object TTF SD Attention-based Operation
Xie et al. 51lu2024object Object TTF SD Regularization; Mask-assisted Generation
Infusion 157zeng2024infusion Object TTF SD 1.5 Attention-based Operation
Pair Customization 160jones2024customizing Object TTF SDXL Disentanglement Approach
DisenDreamer 171chen2024disendreamer Object TTF SD 2.1 Disentanglement Approach
DreamBooth+++++ + 219fan2024dreambooth Object TTF SD Attention-based Operation
CoRe 218wu2024core Object TTF SD 2.1 Regularization
DCI 216jin2025customized Object TTF SD Token Embedding Optimization
Re-Imagen 67chen2022re Object PTA Imagen Retrieval-augmented Paradigm
Versatile Diffusion 109xu2023versatile Object PTA SD 1.4 Architecture Design
Tuning-Encoder 12gal2023encoder Object PTA SD Regularization
ELITE 19wei2023elite Object PTA SD 1.4 Attention-based Operation; Mask-assisted Generation
UMM-Diffusion 71ma2023unified Object PTA SD 1.5 Reference Feature Injection
SuTI 11chen2024subject Object PTA Imagen Data Augmentation
InstantBooth 22shi2023instantbooth Object PTA SD 1.4 Patch Feature Extraction
BLIP-Diffusion 10li2024blip Object PTA SD 1.5 Data Augmentation; Mask-assisted Generation
Domain-Agnostic 17arar2023domain Object PTA SD Regularization; Attention-based Operation
IP-Adapter 40ye2023ip Object PTA SD 1.5 Reference Feature Injection
Kosmos-G 77pan2023kosmos Object; Face PTA SD 1.5 Multimodal Language Modeling
Kim et al. 47kim2023personalized Object PTA SD 1.4 Reference Feature Injection
CAFÉ 84zhou2023customization Object; Face PTA SD 2.0 Multimodal Language Modeling
SAG 99pan2023towards Object; Face; Style PTA SD 1.5 Gradient-based Guidance
BootPIG 41purushwalkam2024bootpig Object PTA SD Data Augmentation
CustomContrast 217chen2024customcontrast Object PTA SD 1.5 Contrastive Learning
MoMA song2024moma Object PTA SD 1.5 Multimodal LLM Adapter
Diptych Prompting 209shin2024large Object; Style PTA FLUX Attention-based Operation; Mask-assisted Generation
Table 3: Paper summary on personalization.
Paper Scope Framework Technique Backbone
StyleDrop 5sohn2024styledrop Style; Multiple Subjects TTF Muse Data Augmentation
StyleBoost 53park2023styleboost Style TTF SD 1.5 Regularization Dataset
StyleAligned 98hertz2023style Style TTF SDXL Regularization
GAL 151zhang2024generative Style; Object TTF SD 1.5 Data Augmentation
StyleForge 153park2024text Style TTF SD 1.5 Parameter-efficient Fine-tuning
FineStyle 213zhangfinestyle Style TTF Muse Mask-assisted Generation
Style-Friendly 211choi2024style Style TTF FLUX-dev; SD 3.5 SNR Sampler
StyleAdapter 26wang2023styleadapter Style PTA SD Data Augmentation; Reference Feature Fxtraction
ArtAdapter 42chen2023artadapter Style PTA SD 1.5 Data Augmentation
ProFusion 55zhou2023enhancing Face TTF SD 2.0 Fusion Sampling
Celeb Basis 60yuan2023inserting Face TTF SD Face Feature Representation
HyperDreamBooth 58ruiz2023hyperdreambooth Face TTF; PTA SD 1.5 Pretraining and Fast Fine-tuning
Banerjee et al. 39banerjee2023identity Face TTF SD 1.4 Regularization
Magicapture 90hyung2023magicapture Face; Multiple Subjects TTF SD 1.5 Attention-based Operation; Mask-assisted Generation; Identity Loss
Concept-centric 35cao2023concept Face TTF SD 1.5 Modified Classifier-free Guidance
Cross Initialization 66pang2023cross Face TTF SD 2.1 Regularization
OMG 115kong2024omg Face; Multiple Subjects TTF SDXL Attention-based Operation; Mask-assisted Generation; Identity Loss
InstantFamily 161kim2024instantfamily Face; Multiple Subjects; Extra Conditions TTF SD 1.5 Mask-assisted Generation
PersonaMagic 203li2024personamagic Face TTF SD 1.4 Attention-based Operation
Su et al. 49su2023identity Face PTA LDM; StyleGAN Identity Loss; Multi-task Learning
FastComposer 54xiao2023fastcomposer Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
Face0 88valevski2023face0 Face PTA SD 1.4 Fusion Sampling
DreamIdentity 89chen2023dreamidentity Face PTA SD 2.1 Attention-based Operation
Face-Diffuser 52wang2023high Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
W-Plus-Adapter 91li2023stylegan Face PTA SD 1.5; StyleGAN Face Feature Representation
Portrait Diffusion 32liu2023portrait Face PTA SD 1.5 Mask-assisted Generation
RetriNet 92tang2023retrieving Face PTA SD Mask-assisted Generation
FaceStudio 31yan2023facestudio Face PTA SD Attention-based Operation
PVA 27xu2024personalized Face PTA LDM Mask-assisted Generation
DemoCaricature 94chen2023democaricature Face; Extra Conditions PTA SD 1.5 Regularization
PhotoMaker 28li2023photomaker Face PTA SDXL Data Collection; Reference Feature Extraction
Stellar 95achlioptas2023stellar Face PTA SDXL Evaluation Prompts and Metrics
PortraitBooth 96peng2023portraitbooth Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
InstantID 37wang2024instantid Face; Extra Conditions PTA SDXL Multi-feature Injection
ID-Aligner 162chen2024id Face PTA SD 1.5; SDXL Feedback Learning
MoA 164ostashev2024moa Face; Multiple Subjects PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
IDAdapter 166cui2024idadapter Face PTA SD 2.1 Mixed Facial Features
Infinite-ID 169wu2024infinite Face PTA SDXL Reference Feature Injection
Face2Diffusion 170shiohara2024face2diffusion Face PTA SD Multi-feature Injection
FreeCure 210cai2024foundation Face PTA SD 1.5; SDXL Mask-assisted Generation
Omni-ID 205qian2024omni Face PTA FLUX Face representation enhancement
Table 4: Paper summary on personalization.
Paper Scope Framework Technique Backbone
Custom Diffusion 6kumari2023multi Multiple Subjects TTF SD 1.4 Parameter-efficient Fine-tuning; Constrained Optimization
Cones 4liu2023cones Multiple Subjects TTF SD 1.4 Concept Neurons Activation
SVDiff 8han2023svdiff Multiple Subjects TTF SD Data Augmentation; Attention-based Operation
Perfusion 7tewel2023key Multiple Subjects TTF SD 1.5 Regularization
Mix-of-Show 46gu2024mix Multiple Subjects TTF SD 1.5 Attention-based Operation; Mask-assisted Generation
Cones-2 62liu2023cones Multiple Subjects TTF SD 2.1 Attention-based Operation; Regularization; Mask-assisted Generation
PACGen 34li2023generate Multiple Subjects TTF SD 1.4 Mask-assisted Generation
Compositional Inversion 9zhang2023compositional Multiple Subjects TTF SD Attention-based Operation; Regularization
EM-style 44rahman2024visual Multiple Subjects TTF SD Attention-based Operation; Mask-assisted Generation
MC2 154jiang2024mc Multiple Subjects TTF SD 1.5 Attention-based Operation
MultiBooth 156zhu2024multibooth Multiple Subjects TTF SD 1.5 Mask-assisted Generation
Matsuda et al. 165matsuda2024multi Multiple Subjects TTF LDM Mask-assisted Generation
MagicTailor zhou2024magictailor Multiple Subjects TTF SD 2.1 Mask-assisted Generation; Attention-based Operation
AnyDoor 14chen2023anydoor Multiple Subjects PTA SD 2.1 Mask-assisted Generation
Subject-Diffusion 16ma2023subject Multiple Subjects PTA SD 2.0 Attention-based Operation; Data Augmentation; Mask-assisted Generation
CustomNet 29yuan2023customnet Multiple Subjects; Extra Conditions PTA SD Multi-condition Integration
MIGC 174zhou2024migc Multiple Subjects PTA SD Attention-based Operation; Mask-assisted Generation
Emu2 223sun2024generative Multiple Subjects PTA SDXL In-Context Learning
SSR-Encoder 221zhang2024ssr Multiple Subjects PTA SD 1.5 Attention-based Operation
$\lambda$-eclipse 222patel2024lambda Multiple Subjects PTA Kandinsky v2.2 Contrasstive Learning
MS-Diffusion 220personalizationms Multiple Subjects PTA SD Attention-based Operation
ReVersion 63huang2023reversion High-level Semantic TTF SD Regularization
Inv-ReVersion 159zhang2024inv High-level Semantic TTF SD 1.5 Regularization
CusConcept 215xu2024cusconcept High-level Semantic TTF SD 2.1 Regularization
ADI 43huang2023learning High-level Semantic TTF SD 2.1 Mask-assisted Generation
PhotoSwap 15gu2024photoswap Extra Conditions TTF SD 2.1 Attention-based Operation
PE-VITON 38zhang2023two Extra Conditions TTF SD Shape and Texture Control
Layout-Control 152chen2024training Extra Conditions TTF SD Attention-based Operation
SwapAnything 155gu2024swapanything Extra Conditions TTF SD 2.1 Mask-assisted Generation
PE-VITON 38zhang2023two Extra Conditions TTF SD Shape and Texture Control
Viewpoint Control 163kumari2024customizing Extra Conditions TTF SDXL 3D Feature Incorporation
Prompt-Free Diffusion 110xu2023prompt Extra Conditions PTA SD 2.0 Reference Feature Injection
Uni-ControlNet 111zhao2024uni Extra Conditions PTA SD Multi-feature Injection
ViscoNet 93cheong2023visconet Extra Conditions PTA SD Multi-feature Injection
Context Diffusion 112najdenkoska2023context Extra Conditions PTA LDM Multi-feature Injection
FreeControl 113mo2023freecontrol Extra Conditions PTA SD 1.5; SD 2.1; SDXL Fusion Guidance
Li et al. 168li2024tuning Extra Conditions PTA SD Attention-based Operation; Mask-assisted Generation
Table 5: Paper summary on personalization.
Paper Scope Framework Technique Backbone
Tune-A-Video 150wu2023tune Video TTF SD 1.4 Parameter-efficient Fine-tuning
Gen-1 148esser2023structure Video TTF LDM Multi-feature Injection
Make-A-Protagonist 135zhao2023make Video TTF LDM Attention-based Operation
Animate-A-Story 136he2023animate Video TTF SD Retrieval-augmented Paradigm
MotionDirector 101zhao2023motiondirector Video TTF ZeroScope; ModelScope Disentanglement Approach
LAMP 102wu2023lamp Video TTF SDXL; SD 1.4 First-frame Conditioned Pipeline
VMC 140jeong2023vmc Video TTF Show-1 Parameter-efficient Fine-tuning
SAVE 103song2023save Video TTF VDM Attention-based Operation
Customizing Motion 104materzynska2023customizing Video TTF ZeroScope Regularization
DreamVideo 142wei2023dreamvideo Video TTF ModelScope Disentanglement Approach
MotionCrafter 106zhang2023motioncrafter Video TTF VDM Disentanglement Approach
Customize-A-Video 172ren2024customize Video TTF ModelScope Multi-stage Refinement
Magic-Me 173ma2024magic Video TTF SD 1.5; AnimateDiff Multi-stage Refinement
CustomTTT 202bi2024customttt Video TTF VDM Fine-tuning and Alignment
PersonalVideo 208li2024personalvideo Video; Face TTF AnimateDiff Identity Injection
CustomVideo 145wang2024customvideo Video; Multiple Subjects TTF ZeroScope Data Augmentation; Attention-based Operation
VideoDreamer 138chen2023videodreamer Video; Multiple Subjects TTF SD 2.1 Data Augmentation
VideoAssembler 82zhao2023videoassembler Video PTA VidRD Reference Feature Injection
VideoBooth 83jiang2023videobooth Video PTA VDM Reference Feature Injection
VideoMaker 201wu2024videomaker Video PTA VDM Reference Feature Injection
SUGAR 204zhou2024sugar Video PTA VDM Attention-based Operation
StyleCrafter 212liu2024stylecrafter Video; Style PTA SD 1.5 Reference Feature Injection
ID-Animator 158he2024id Video; Face PTA AnimateDiff Reference Feature Injection
ConsisID 207yuan2024identity Video; Face PTA CogVideoX frequency decomposition
MotionCharacter 206fang2024motioncharacter Video; Face PTA VDM ID-Preserving Optimization; Motion Control Enhancement
DreaMoving 105feng2023dreamoving Video; Face; Extra Conditions PTA SD Multi-feature Injection
Magic3D 120lin2023magic3d 3D TTF Imagen; LDM Score Distillation Sampling
DreamBooth3D 121raj2023dreambooth3d 3D TTF Imagen Multi-stage Refinement
PAS 122azadi2023text 3D PTA SD Text-to-3D-Pose
StyleAvatar3D 123zhang2023styleavatar3d 3D PTA LDM Multi-view Alignment
AvatarBooth 124zeng2023avatarbooth 3D TTF SD Dual Model Fine-tuning
MVDream 126shi2023mvdream 3D TTF SD 2.1 Score Distillation Sampling
Consist3D 127ouyang2023chasing 3D TTF SD Token Embedding Enhancement
Animate124 128zhao2023animate124 3D TTF SD 1.5 Multi-stage Refinement
Dream-in-4D 130zheng2023unified 3D TTF SD 2.1 Multi-stage Refinement
TextureDreamer 133yeh2024texturedreamer 3D TTF LDM Texture Extraction
TIP-Editor 134zhuang2024tip 3D TTF SD Multi-stage Refinement; Mask-assisted Generation
Anti-DreamBooth 13van2023anti Attack and Defense TTF SD 2.1 Perturbation Learning
Concept Censorship 50zhang2023backdooring Attack and Defense TTF SD 1.4 Trigger Injection
Huang et al. 167huang2024personalization Attack and Defense TTF SD backdoor attack
Continual Diffusion 64smith2023continual Others TTF SD Continual Learning
SVGCustomization 36zhang2023text Others TTF SD 1.5 Fine-tuning and Alignment
StitchDiffusion 25wang2024customizing Others TTF LDM Parameter-efficient Fine-tuning
MC-TI 214wang2024multi Others TTF SD 1.5 Regularization
Acknowledgments
This work was supported in part by Chinese National Natural Science Foundation Projects U23B2054, 62276254, 62372314, Beijing Natural Science Foundation L221013, InnoHK program, and Hong Kong Research Grants Council through Research Impact Fund (project no. R1015-23).
Declarations of Conflict of Interest
The authors declared that they have no conflicts of interest to this work.

### v4

@@
-Corresponding Author
+
 Abstract
@@
 keywords:
-Generative Models, Image Synthesis, Diffusion Models, Personalization
+Generative Models, Image Synthesis, Diffusion Models, Personalized Content Synthesis, Subject Customization
+00footnotetext: Accepted by Machine Intelligence Research, DOI: 10.1007/s11633-025-1563-3
 1 Introduction
 Recently, generative models have shown remarkable progress in the field of natural language processing and computer vision.
-Notable examples, such as ChatGPT wu2023brief and text-to-image diffusion models croitoru2023diffusion , have showcased impressive capabilities in content creation.
+Notable examples, such as ChatGPT [1] and text-to-image diffusion models [2], have showcased impressive capabilities in content creation.
 However, these advanced models often struggle to fulfill specific requirements, such as answering domain-specific queries or accurately depicting user’s portraits.
 This limitation highlights the importance of Personalized Content Synthesis (PCS), which enables users to customize models for their unique tasks and requirements.
-As a critical area of research, PCS is increasingly recognized as essential for achieving Artificial General Intelligence (AGI), evidenced by the growing number of companies releasing products that support personalized content creation, like utilizing Reinforcement Learning to fine-tune language models uc2023survey .
+As a critical area of research, PCS is increasingly recognized as essential for achieving Artificial General Intelligence (AGI), evidenced by the growing number of companies releasing products that support personalized content creation, like utilizing Reinforcement Learning to fine-tune language models [3].
 In this paper, we focus on PCS within the context of diffusion models in computer vision.
@@
 1, PCS can customize a unique cat into various scenarios, such as “wearing pink sunglasses” and “in the snow”.
-Figure 1: Given a few reference images of a subject (e.g., a cat 2ruiz2023dreambooth or face 54xiao2023fastcomposer ), PCS aims to generate new renditions of the subject that align with user-defined textual prompts.
-The task requires preserving the subject’s identity (fidelity) while adapting to diverse contexts.
-The examples are generated by this survey using DreamBooth 2ruiz2023dreambooth and InstantID 37wang2024instantid .
+Figure 1: Given a few reference images of a subject (e.g., a cat [4] or face [5]), PCS aims to generate new renditions of the subject that align with user-defined textual prompts.
+The task requires preserving the subject’s identity while adapting to diverse contexts.
+The examples are generated by this survey using DreamBooth [4] and InstantID [6].
 The emergence of diffusion models has significantly facilitated text-guided content generation, leading to a rapid expansion of PCS.
@@
 2, the number of research papers on PCS has surged over time.
-Following the release of key innovations such as DreamBooth 2ruiz2023dreambooth and Textual Inversion 1gal2022image in August 2022, over 150 methods have been proposed in a remarkably short period.
+Following the release of key innovations such as DreamBooth [4] and Textual Inversion [7] in August 2022, over 150 methods have been proposed in a remarkably short period.
 These methods can be classified using several criteria.
@@
 We divide PCS methods with 3 different criteria: training strategy, personalization scope, and technique.
+Figure 3: The trade-off between text alignment and visual fidelity in personalized image synthesis, illustrated through DreamBooth-generated [4] examples of a customized cat wearing sunglasses.
+Overfitting occurs when the model focuses solely on reconstructing the cat, disregarding the sunglasses context.
+Underfitting, on the other hand, reflects the model’s attempt to satisfy the text prompt but fails to accurately represent the personalized cat.
+Collapse signifies a failure to meet both criteria.
 While current methods in PCS have shown impressive performance, several challenges remain unresolved.
@@
 By addressing these obstacles, we aim to advance the field of personalized content synthesis and enhance its practical applications.
-Figure 3: The trade-off between text alignment and visual fidelity in personalized image synthesis, illustrated through DreamBooth-generated 2ruiz2023dreambooth examples of a customized cat wearing sunglasses.
-Overfitting occurs when the model focuses solely on reconstructing the cat, disregarding the sunglasses context.
-Underfitting, on the other hand, reflects the model’s attempt to satisfy the text prompt but fails to accurately represent the personalized cat.
-Collapse signifies a failure to meet both criteria.
 This survey aims to provide a comprehensive overview of the existing methods, frameworks, and challenges of PCS to help readers understand current trends in this evolving landscape and promote further improvements in this area.
-The structure of this survey is organized into several key sections: In Sec.
-2, We begin with a brief introduction to diffusion models for better understanding of PCS.
-In Sec.
-3, we introduce two primary frameworks for PCS: test-time fine-tuning and pre-trained adaptation.
-In Sec.
-4, we categorize common techniques into four main parts, attention-based methods, mask-based generation, data augmentation, and regularization.
-In Sec.
-5, we summarize various PCS methods based on different personalization tasks, including personalization on object, face, style, and etc.
-In Sec.
-6, we demonstrate PCS’s growing impact in video, 3D, and other areas beyond traditional image generation tasks.
-In Sec.
-7, we review the current evaluation dataset and metrics and construct a benchmark on the existing methods for further research.
-In Sec.
-8, we discuss current challenges faced in PCS and propose potential future directions to overcome these problems.
+The structure of this survey is organized into several key sections: In Section 2, We begin with a brief introduction to diffusion models for better understanding of PCS.
+In Section 3, we introduce two primary frameworks for PCS: test-time fine-tuning and pre-trained adaptation.
+In Section 4, we categorize common techniques into four main parts, attention-based methods, mask-based generation, data augmentation, and regularization.
+In Section 5, we summarize various PCS methods based on different image personalization tasks, including personalization on object, face, style, and etc.
+In Section 6, we demonstrate PCS’s growing impact in video, 3D, and other areas beyond traditional image generation tasks.
+In Section 7, we review the current evaluation dataset and metrics, and introduce a benchmark on the existing methods to promote further research.
+In Section 8, we discuss current challenges faced in PCS and propose potential future directions in this area.
+Additionally, we present a comprehensive summary of all relerated papers in Table 2, Table 3, and Table 4.
 2 Fundamentals
-Modern diffusion modeling has evolved into a sophisticated framework that unifies discrete and continuous generative paradigms through stochastic differential equations (SDEs) song2020score and ordinary differential equations (ODEs) song2020score ; karras2022elucidating .
-This section introduces the core mathematical foundations, emphasizing innovations in the diffusion process and conditional mechanisms critical for personalized generation.
+Modern diffusion modeling has evolved into a sophisticated framework that unifies discrete and continuous generative paradigms through stochastic differential equations (SDEs) [8], ordinary differential equations (ODEs) [8, 9].
+This section introduces the core mathematical foundations, emphasizing innovations in the diffusion process and conditional mechanisms.
+By covering these theoretical developments, this section lays the groundwork for the personalization tasks explored in subsequent sections.
 2.1 Denoising Diffusion Probabilistic Models (DDPMs)
 The architecture of diffusion models comprises two complementary processes: a forward diffusion that systematically perturbs data distributions, and a learned reverse process that reconstructs signals through iterative refinement.
-Originating from discrete-time Markov chains in Denoising Diffusion Probabilistic Model (DDPM) ho2020denoising , the forward process applies Gaussian noise corruption over $T$ steps:
+Originating from discrete-time Markov chains in Denoising Diffusion Probabilistic Model (DDPM) [10], the forward process applies Gaussian noise corruption over $T$ steps:
 $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},\quad\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ (1)
@@
 The generative capability resides in learning to invert this degradation through parameterized reverse transitions. For DDPMs, this inversion is achieved via Bayesian reconstruction:
-$p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t}\right)=\mathcal{N}\left(\mathbf{x}_{t-1};\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t)\right),\beta_{t}I\right)$ (2)
+$p_{\theta}\left(\mathbf{x}_{t-1}\mid\mathbf{x}_{t}\right)=\\ \mathcal{N}\left(\mathbf{x}_{t-1};\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{x}_{t},t)\right),\beta_{t}I\right)$ (2)
 Here, $p_{\theta}$ represents the parameterized reverse transition probability, $\epsilon_{\theta}$ is the neural network predicting noise components, and $\beta_{t}$ is a hyperparameter before model training.
@@
 This objective enables end-to-end training of the denoising trajectory.
+Once training is finished, the model can apply the learned reverse diffusion transitions to any arbitrary noise input, progressively denoising it into a coherent data sample.
+This capability allows DDPMs to function as general-purpose generative engines, producing high-quality outputs from random noise vectors.
 2.2 Stochastic Differential Equations (SDEs)
 Although DDPMs establish basic denoising dynamics, this discrete formulation imposes constraints on flexible noise scheduling and computational efficiency.
-To overcome these limitations, the framework was generalized through continuous-time SDE song2020score , which provide a unified perspective for modeling diffusion dynamics:
+To overcome these limitations, the framework was generalized through continuous-time SDE [8], which provide a unified perspective for modeling diffusion dynamics:
 $d\mathbf{x}=\mathbf{f}(\mathbf{x},t)dt+g(t)d\mathbf{w}$ (4)
 where $\mathbf{f}(\mathbf{x},t)$ encodes deterministic drift components, $g(t)$ modulates stochastic diffusion via Wiener process $\mathbf{w}$.
-This continuous perspective subsumes DDPMs as special cases while enabling adaptive noise scheduling strategies such as variance-preserving (VP) and variance-exploding (VE) schedules through strategic choices of $\mathbf{f}$ and $g$ karras2022elucidating .
-For the reverse process, a crucial result anderson1982reverse shows that the reverse diffusion process also follows an SDE:
+This continuous perspective subsumes DDPMs as special cases while enabling adaptive noise scheduling strategies such as variance-preserving (VP) and variance-exploding (VE) schedules through strategic choices of $\mathbf{f}$ and $g$ [9].
+For the reverse process, a crucial result [11] shows that the reverse diffusion process also follows an SDE:
 $d\mathbf{x}=\left[\mathbf{f}(\mathbf{x},t)-g(t)^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right]dt+g(t)d\mathbf{\bar{w}}$ (5)
@@
 Furthermore, stochastic trajectories exhibit erratic curvature profiles, forcing fixed-step solvers to adopt conservative discretization schemes that often necessitate 1,000+ steps for high-fidelity generation.
-These limitations motivate the derivation of deterministic sampling trajectories through probability flow ordinary differential equations (ODEs) song2020score ; karras2022elucidating , obtained by eliminating the stochastic term from the SDE formulation:
+These limitations motivate the derivation of deterministic sampling trajectories through probability flow ordinary differential equations (ODEs) [8, 9], obtained by eliminating the stochastic term from the SDE formulation:
 $d\mathbf{x}=\left[\mathbf{f}(\mathbf{x},t)-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right]dt$ (6)
@@
 Crucially, the resulting trajectories exhibit intrinsic geometric regularity, allowing the complex denoising process to operate in a simplified mathematical space.
-This enabling 5-10x faster sampling than SDEs through adaptive ODE solvers like DPM-Solver lu2022dpm and DPM-Solver++ lu2022dpmplus .
-Recent work chen2024trajectory further identifies strong geometric regularity in these ODE-based sampling trajectories, demonstrating that they inherently follow a linear-nonlinear-linear structure regardless of generated content.
+This enabling 5-10x faster sampling than SDEs through adaptive ODE solvers like DPM-Solver [12] and DPM-Solver++ [13].
+Recent work [14] further identifies strong geometric regularity in these ODE-based sampling trajectories, demonstrating that they inherently follow a linear-nonlinear-linear structure regardless of generated content.
 This trajectory regularity enables dynamic programming-based time scheduling optimization with negligible computational overhead.
@@
 $\mathcal{L}=\mathbb{E}_{\mathbf{x}_{0},t,\mathbf{\epsilon}}\left[\|\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}(\mathbf{x}_{t},t,c)\|_{2}^{2}\right]$ (7)
-where the conditioning signal $c$ can be integrated via multiple synergistic mechanisms, such as cross-modal attention rombach2022high and spatial modulation zhang2023adding .
+where the conditioning signal $c$ can be integrated via multiple synergistic mechanisms, such as cross-modal attention [15] and spatial modulation [16].
 This conditional paradigm directly enables the applications of PCS.
-Leading text-to-image systems, such as Stable Diffusion (SD) rombach2022high and DALLE ramesh2022hierarchical , are widely adopted to empower users in controlling customized content through text instructions.
+Leading text-to-image systems, such as Stable Diffusion (SD) [15] and DALLE [17], are widely adopted to empower users in controlling customized content through text instructions.
 3 Generic Framework
@@
 2) Semantic-aware modifier system represents the SoI at the token level to bridge the gap between visual adaptation and textual control.
-In the following Sec.
-3.1.1 and Sec.
-3.1.2, we will detail these principles respectively.
+In the following Section 3.1.1 and Section 3.1.2, we will detail these principles respectively.
 3.1.1 Test-time Adaption
@@
 where $c$ represents the condition signal, typically the reference image caption.
-Compared to the large-scale pre-training described in Eq.
-7, key differences lie in the training data and learnable parameters.
-The training samples are typically restricted to the SoI references, sometimes supplemented with a regularization dataset to mitigate overfitting 2ruiz2023dreambooth .
-For the selection of learnable parameters $\mathbf{\theta^{\prime}}$, the commonly adopted options include token embeddings 1gal2022image ; 23voynov2023p+ , the entire diffusion model 2ruiz2023dreambooth ; 151zhang2024generative , specific subsets of parameters 6kumari2023multi ; 7tewel2023key ; 8han2023svdiff , or introducing new parameters such as adapters 72xiang2023closer ; 5sohn2024styledrop , and LoRA 46gu2024mix ; 58ruiz2023hyperdreambooth ; 115kong2024omg , which will be discussed in Sec.
-3.1.3.
+Compared to the large-scale pre-training described in Equation 7, key differences lie in the training data and learnable parameters.
+The training samples are typically restricted to the SoI references, sometimes supplemented with a regularization dataset to mitigate overfitting [4].
+For the selection of learnable parameters $\mathbf{\theta^{\prime}}$, the commonly adopted options include token embeddings [7, 18], the entire diffusion model [4, 19], specific subsets of parameters [20, 21, 22], or introducing new parameters such as adapters [23, 24], and LoRA [25, 26, 27], which will be discussed in Section 3.1.3.
 Test-time generation.
 The test-time generation is proceeded once the model has been fine-tuned with the optimized parameters $\mathbf{\theta^{\prime}}$.
-By composing novel input prompts that incorporate the SoI’s unique identifier (introduced in Sec.
-3.1.2), the adapted model can synthesize diverse images while preserving the subject’s distinctive characteristics.
+By composing novel input prompts that incorporate the SoI’s unique identifier (introduced in Section 3.1.2), the adapted model can synthesize diverse images while preserving the subject’s distinctive characteristics.
 This approach enables flexible control over the generated content through natural language descriptions.
 3.1.2 Unique Modifier
-The unique modifier is a critical component in PCS, enabling text-based representation of the SoI for flexible prompt composition.
+A unique modifier is a textual token or short phrase that uniquely represents a SoI, enabling text-based representation of the SoI for flexible prompt composition.
 As illustrated in Fig.
@@
 Normally, the construction of the unique modifier can be divided into three categories:
-Learnable token embedding 1gal2022image ; 23voynov2023p+ ; 61alaluf2023neural ; 30zhang2023prospect ; 48wang2023hifi ; 66pang2023cross .
-This method adds a new token and its corresponding embedding vector to the word dictionary of the tokenizer.
-An intuitive example is that this method creates a new word that does not exist in the dictionary.
-This word acts as the modifier, with adjustable weights during fine-tuning, while the embeddings of other tokens in the pre-defined dictionary will not be influenced.
-This approach requires only a few kilobytes of additional parameters and maintains the base model’s capabilities for non-customized generation.
-Plain text 2ruiz2023dreambooth ; 6kumari2023multi ; 7tewel2023key ; 5sohn2024styledrop ; 151zhang2024generative .
+Plain text [4, 20, 21, 24, 19].
 This approach utilizes an explicit text description to represent the SoI.
 For example, words such as cat could directly represent the user’s cat in the references.
-This setting typically requires fine-tuning the parameters of the diffusion model components such as UNet ronneberger2015unet or Transformer blocks vaswani2017attention .
+This setting typically requires fine-tuning the parameters of the diffusion model components, such as UNet [28] or Transformer blocks [29], to enable the model to associate the SoI’s visual features with the plain‐text token.
 The plain text offers user-friendly prompt construction and injects subject prior information to ease tuning difficulty.
 However, this technique may over-specialize common terms, limiting their broader applicability as the model learns to associate general words with specific SoI characteristics.
-Rare token 2ruiz2023dreambooth .
+Rare token [4].
 This technique employs infrequently used tokens to minimize their impact on commonly used vocabulary.
 Similar to the plain text approach, the embeddings of rare tokens remain unchanged during fine-tuning.
-However, these rare tokens fail to provide useful subject prior information and still exhibit weak interference with unrelated vocabulary, potentially leading to ambiguity between the original meaning and the intended SoI reference.
+However, these rare tokens often fail to provide useful subject prior information and still exhibit weak interference with unrelated vocabulary, potentially leading to ambiguity between the original meaning and the intended SoI reference.
+Learnable token embedding [7, 18, 30, 31, 32, 33].
+This method adds a new token and its corresponding embedding vector to the dictionary of the tokenizer.
+An intuitive example is that this method creates a new word that does not exist in the dictionary.
+This inserted token has adjustable weights during fine-tuning, while the embeddings of other tokens in the pre-defined dictionary will remain unchanged.
+This approach requires only a few kilobytes of additional parameters and maintains the base model’s capabilities for non-customized generation.
+Similar to the rare token approach, it is not very user-friendly in practice, since users must learn and remember an unfamiliar token to reference their subject of interest.
 3.1.3 Training Parameter Selection
 The selection of trainable parameters represents a critical design consideration in PCS, directly impacting many key performance metrics: subject fidelity, training efficiency, and model storage requirements.
-Current tunable parameters can be broadly categorized into token embedding, existing model components, parameter-efficient techniques, and hybrid approaches.
+Current tunable parameters can be broadly categorized into the following four types.
 Token embedding.
-As introduced in Sec.
-3.1.2, token embedding optimization 1gal2022image introduces learnable tokens as the unique modifier to represent SoI through noise-to-image reconstruction.
-While achieving remarkable parameter efficiency, this approach faces challenges in detail preservation and prolonged training time due to the inherent compression of complex features into low-dimensional embeddings.
-Subsequent works 23voynov2023p+ ; 61alaluf2023neural ; 30zhang2023prospect ; 48wang2023hifi ; 66pang2023cross aim to address these limitations through different strategies, which are summarized in Sec.
-5.1.
+As introduced in Section 3.1.2, token embedding optimization [7] introduces learnable tokens as the unique modifier to represent SoI through noise-to-image reconstruction.
+While achieving remarkable parameter efficiency, this approach faces challenges in detail preservation and prolonged training time (typically larger than 20 minutes) due to the inherent compression of complex features into low-dimensional embeddings.
+Subsequent works [18, 30, 31, 32, 33] aim to address these limitations through different strategies, which are summarized in Section 5.1.
 Existing model parameters.
-This paradigm directly optimizes pre-trained components, such as the text encoder, U-Net blocks, and transformer layers 2ruiz2023dreambooth ; 151zhang2024generative ; 6kumari2023multi ; 7tewel2023key ; 8han2023svdiff .
+This paradigm directly optimizes pre-trained model components, such as the text encoder, U-Net blocks, and transformer layers [4, 19, 20, 21, 22].
 Benefitting from the advanced representation capacity of these modules, the fine-tuning phase can achieve faster convergence (5-10 minutes) and superior visual fidelity compared to token-only methods, though at the cost of significant storage overhead.
@@
 Parameter-efficient extensions.
-Recent advanced methods have introduced parameter-efficient techniques into PCS, such as LoRA 46gu2024mix ; 58ruiz2023hyperdreambooth ; 115kong2024omg ; 51lu2024object ; 95achlioptas2023stellar ; 76yeh2023navigating and adapter modules 72xiang2023closer ; 5sohn2024styledrop , which inject small, trainable components into the base model.
+Recent advanced methods have introduced parameter-efficient techniques into PCS, such as LoRA [25, 26, 27, 34, 35, 36] and adapter modules [23, 24], which inject small, trainable components into the base model.
 These methods achieve comparable performance to full parameter fine-tuning while dramatically reducing storage requirements.
@@
 Since the aforementioned strategies are not conflicting, some methods allocate different learning rates and training phases to each component type to achieve optimal balance between fidelity and efficiency.
-For instance, the fine-tuned token embedding can be regarded as an effective initialization for the subsequent model weight fine-tuning 90hyung2023magicapture .
-Also, these two parts can be simultaneously optimized with different learning rates 74avrahami2023break ; 86hua2023dreamtuner .
+For instance, the fine-tuned token embedding can be regarded as an effective initialization for the subsequent model weight fine-tuning [37].
+Also, these two parts can be simultaneously optimized with different learning rates [38, 39].
 3.1.4 Prompt Engineering
@@
 A simplest example is “Photo of V*”.
-However, DreamBooth 2ruiz2023dreambooth noted that such a simple description causes a long training time and unsatisfactory performance.
+However, DreamBooth [4] noted that such a simple description causes a long training time and unsatisfactory performance.
 To address this, they incorporate the unique modifier with a class noun to describe the SoI in the references (e.g., “Photo of V* cat”).
-Also, the training caption for each training reference can be more precious for better disentanglement of SoI and irrelevant concepts 57he2023data , such as “Photo of V* cat on the chair”.
-This follows the tending that high-quality captions in the training set could assist in further improvement of accurate text control BetkerImprovingIG .
+Also, the training caption for each training reference can be more precious for better disentanglement of SoI and irrelevant concepts [40], such as “Photo of V* cat on the chair”.
+This follows the tending that high-quality captions in the training set could assist in further improvement of accurate text control [41].
+The PTA framework has emerged as a breakthrough approach for PCS, aiming to eliminate the computational burden of per-request fine-tuning while maintaining high-quality, subject-specific generation capabilities.
+To achieve this goal, this approach combines large-scale pretraining with reference-aware architectures to enable single-pass personalization, as shown in Fig.
+5.
+Based on this architecture, three critical design factors are considered to ensure practical viability: 1) Preservation of semantic-critical features to guarantee visual consistency with reference inputs, 2) Effective fusion that combines reference features with text guidance to achieve desired generation, 3) Optimization of training dataset scale to achieve robust generalization without overfitting.
+In the following sub-section, we first present overall architecture introduction in Section 3.1.5 and then delve into three detailed factors.
+3.1.5 Pre-trianed Adaption
 3.2 Pre-trained Adaptation (PTA) Framework
@@
 Additionally, the concatenation of image and text features can be implemented in various ways, such as placeholder-based and reference-conditioned.
-The PTA framework has emerged as a breakthrough approach for PCS, aiming to eliminate the computational burden of per-request fine-tuning while maintaining high-quality, subject-specific generation capabilities.
-To achieve this goal, this approach combines large-scale pretraining with reference-aware architectures to enable single-pass personalization, as shown in Fig.
-5.
-Based on this architecture, three critical design factors are considered to ensure practical viability: 1) Preservation of semantic-critical features to guarantee visual consistency with reference inputs, 2) Effective fusion that combines reference features with text guidance to achieve desired generation, 3) Optimization of training dataset scale to achieve robust generalization without overfitting.
-In the following sub-section, we first present overall architecture introduction in Sec.
-3.2.1 and then delve into three detailed factors.
-3.2.1 Pre-trianed Adaption
 Pre-training.
@@
 Extracting representative features of the SoI is crucial in the creation of personalized content.
-A common approach is to employ an encoder, leveraging pre-trained models such as CLIP radford2021learning and BLIP li2022blip .
+A common approach is to employ an encoder, leveraging pre-trained models such as CLIP [42] and BLIP [43].
 While these models excel at capturing global features, they often include irrelevant information that can detract from the fidelity, potentially compromising the quality of the personalized output, such as including the same background in the generation.
 To mitigate this issue, some studies incorporate additional prior knowledge to guide the learning process so as to focus on the targeted SoI.
-For instance, the SoI-specific mask 21zhang2023personalize ; 48wang2023hifi ; 74avrahami2023break ; 81safaee2023clic ; 51lu2024object ; 95achlioptas2023stellar contributes to the effective exclusion of the influence of the background.
-Moreover, using facial landmarks 37wang2024instantid in the context of human face customization helps improve identity preservation.
-We will discuss a more detailed technical summary of mask-assistant generation in Sec.
-4.2.
+For instance, the SoI-specific mask [44, 32, 38, 45, 34, 35] contributes to the effective exclusion of the influence of the background.
+Moreover, using facial landmarks [6] in the context of human face customization helps improve identity preservation.
+We will discuss a more detailed technical summary of mask-assistant generation in Section 4.2.
 Handling multiple input references presents another challenge but is essential for real-world deployment.
@@
 Yet, the majority of current PTA systems are limited to supporting one reference input.
-Some research works 37wang2024instantid ; 28li2023photomaker propose to average or stack features extracted from multiple references to form a composite SoI representation.
+Some research works [6, 46] propose to average or stack features extracted from multiple references to form a composite SoI representation.
 3.2.2 Subject Feature Fusion
@@
 Current methodologies can be categorized into four primary approaches:
-Concatenation-based Fusion 19wei2023elite ; 28li2023photomaker ; 54xiao2023fastcomposer ; 89chen2023dreamidentity ; 85cai2023decoupled ; 96peng2023portraitbooth .
+Concatenation-based Fusion [47, 46, 5, 48, 49, 50].
 This method makes the unique modifier a placeholder token to encapsulate visual subject characteristics.
@@
 This combined representation subsequently guides generation through standard cross-attention layers in the diffusion process, enabling basic subject-text alignment while maintaining architectural simplicity.
-Cross-attention Fusion 40ye2023ip ; 11chen2024subject ; 37wang2024instantid ; 91li2023stylegan ; 212liu2024stylecrafter ; 221zhang2024ssr ; 220personalizationms .
+Cross-attention Fusion [51, 52, 6, 53, 54, 55, 56].
 This paradigm extends the U-Net architecture with specialized attention mechanisms that jointly process visual and textual conditions.
-For example, IP-Adapter 40ye2023ip introduces decoupled cross-attention layers that maintain separate query projections for image and text features.
+For example, IP-Adapter [51] introduces decoupled cross-attention layers that maintain separate query projections for image and text features.
 In this case, the unique modifier directly displays the plain text.
-Multimodal Encoder Fusion 10li2024blip ; 84zhou2023customization ; song2024moma ; 77pan2023kosmos ; 71ma2023unified .
-This approach leverages powerful multimodal encoder architectures (e.g., BLIP-2 li2023blip ) to jointly embed visual and textual subject descriptors.
-BLIP-Diffusion 10li2024blip exemplifies this strategy by learning a compact subject prompt embedding that fuses image patches with textual names through Q-Former modules.
-Hybrid Fusion 16ma2023subject ; 156zhu2024multibooth .
+Multimodal Encoder Fusion [57, 58, 59, 60, 61].
+This approach leverages powerful multimodal encoder architectures (e.g., BLIP-2 [62]) to jointly embed visual and textual subject descriptors.
+BLIP-Diffusion [57] exemplifies this strategy by learning a compact subject prompt embedding that fuses image patches with textual names through Q-Former modules.
+Hybrid Fusion [63, 64].
 Moreover, some systems integrate multiple fusion strategies.
-For example, Subject-Diffusion 16ma2023subject integrate both concatenation and cross-attention fusion, taking advantage of the strengths of each approach to enhance the overall personalization capability.
+For example, Subject-Diffusion [63] integrate both concatenation and cross-attention fusion, taking advantage of the strengths of each approach to enhance the overall personalization capability.
 3.2.3 Training Data
@@
 Several strategies have been proposed to mitigate this issue: 1) Data Augmentation.
-Techniques such as foreground segmentation followed by placement in a different background are used to construct triplet data 10li2024blip .
+Techniques such as foreground segmentation followed by placement in a different background are used to construct triplet data [57].
 2) Synthetic Sample Generation.
-Methods like SuTI 11chen2024subject utilize multiple TTF models to generate synthetic samples, which are then paired with original references.
+Methods like SuTI [52] utilize multiple TTF models to generate synthetic samples, which are then paired with original references.
 3) Utilizing Recognizable SoIs.
-Collecting images of easily recognizable subjects, such as celebrities, significantly facilitates face personalization 60yuan2023inserting .
+Collecting images of easily recognizable subjects, such as celebrities, significantly facilitates face personalization [65].
 Dual Data (Reference Image, Reference Caption).
 This dataset is essentially a simplified version of the triplet format, where the personalized content is the original image itself.
-Such datasets are more accessible, including collections like LAION schuhmann2022laion and LAION-FACE zheng2022general .
+Such datasets are more accessible, including collections like LAION [66] and LAION-FACE [67].
 However, a notable drawback is that training tends to focus more on reconstructing the reference image rather than incorporating the text prompts.
@@
 Recently, some works have started to explore the combination of TTF and PTA methods.
-HyperDreamBooth 58ruiz2023hyperdreambooth states that PTA methods provide a general framework capable of handling a wide range of common objects, while TTF techniques enable fine-tuning to specific instances, which can improve the preservation of fine-grained details.
+HyperDreamBooth [26] states that PTA methods provide a general framework capable of handling a wide range of common objects, while TTF techniques enable fine-tuning to specific instances, which can improve the preservation of fine-grained details.
 They first develop a PTA network, followed by a subject-driven fine-tuning.
-Similarly, DreamTuner 86hua2023dreamtuner pre-trains a subject encoder that outputs diffusion conditions for accurate reconstruction.
+Similarly, DreamTuner [39] pre-trains a subject encoder that outputs diffusion conditions for accurate reconstruction.
 Then an additional fine-tuning stage is conducted for fine identity preservation.
-In contrast, SuTI 11chen2024subject first applies TTF methods to generate synthetic pairwise samples, which can be used for training the PTA network.
+In contrast, SuTI [52] first applies TTF methods to generate synthetic pairwise samples, which can be used for training the PTA network.
 4 Techniques in Personalized Content Synthesis
-Building upon the architectural framework discussed in Sec.
-3, this section analyzes learning optimization techniques applicable to both frameworks.
+Building upon the architectural framework discussed in Section 3, this section analyzes learning optimization techniques applicable to both frameworks.
 We focus on four categories: attention mechanisms, mask-guided generation, data augmentation, and regularization strategies.
@@
 4.1 Attention-based Operation
-Attention-based operations have become a crucial technique in model learning, particularly for processing features effectively hassanin2024visual .
+Attention-based operations have become a crucial technique in model learning, particularly for processing features effectively [68].
 In diffusion models, these operations generally involve manipulating the way a model focuses on different parts of data, often through a method known as the Query-Key-Value (QKV) scheme.
@@
 A cluster of studies focuses on restricting the influence of the SoI token within the attention layers.
-For example, Mix-of-Show 46gu2024mix designs region-aware cross-attention where the feature map is initially generated by the global prompt and replaced with distinct regional features corresponding to each entity.
+For example, Mix-of-Show [25] designs region-aware cross-attention where the feature map is initially generated by the global prompt and replaced with distinct regional features corresponding to each entity.
 This avoids the misalignment between the words and visual regions.
-DreamTuner 86hua2023dreamtuner designs a self-subject-attention layer to further refine the subject identity.
+DreamTuner [39] designs a self-subject-attention layer to further refine the subject identity.
 This attention module takes the features of the generated image as query, the concatenation of generated features as key, and reference features as value.
-Layout-Control 152chen2024training adjusts attention weights specifically around the layout without additional training.
-Cones 2 62liu2023cones also defines some negative attention areas to penalize the illegal occupation to allow multiple object generation.
-VICO 20hao2023vico inserts a new attention layer where a binary mask is deployed to selectively obscure the attention map between the noisy latent and the reference image features.
+Layout-Control [69] adjusts attention weights specifically around the layout without additional training.
+Cones 2 [70] also defines some negative attention areas to penalize the illegal occupation to allow multiple object generation.
+VICO [71] inserts a new attention layer where a binary mask is deployed to selectively obscure the attention map between the noisy latent and the reference image features.
 Implicit attention guidance.
-In addition to these explicit attention weights modification methods, many researchers 54xiao2023fastcomposer ; 8han2023svdiff ; 74avrahami2023break ; 16ma2023subject ; 81safaee2023clic ; 44rahman2024visual ; 90hyung2023magicapture employ localization supervision in the cross-attention module.
+In addition to these explicit attention weights modification methods, many researchers [5, 22, 38, 63, 45, 72, 37] employ localization supervision in the cross-attention module.
 Specifically, they train cross-attention modules using coordinate-aware loss functions, forcing attention maps to align with annotated subject positions.
-DreamTuner 86hua2023dreamtuner further refines this approach by designing an attention layer that effectively integrates features from different parts of the image.
+DreamTuner [39] further refines this approach by designing an attention layer that effectively integrates features from different parts of the image.
 4.2 Mask-guided Generation
@@
 Pixel-level mask.
-Benefitting from advanced segmentation methods like SAM kirillov2023segment , the SoI can be precisely isolated from the background.
-Based on this strategy, plenty of studies 21zhang2023personalize ; 48wang2023hifi ; 74avrahami2023break ; 81safaee2023clic ; 51lu2024object ; 95achlioptas2023stellar ; 90hyung2023magicapture ; 92tang2023retrieving ; 34li2023generate choose to discard the pixels of the background area so that the reconstruction loss can focus on the targeted object and exclude irrelevant disturbances.
-Another technique 85cai2023decoupled further adds the masked background reconstruction for better disentanglement.
-In addition, the layout indicated by the pixel mask can be incorporated into the attention modules as a supervision signal 54xiao2023fastcomposer ; 8han2023svdiff ; 74avrahami2023break ; 16ma2023subject ; 81safaee2023clic ; 44rahman2024visual ; 90hyung2023magicapture to adjust the attention’s concentration adaptively.
-Moreover, the mask can stitch specific feature maps to construct more informative semantic patterns 19wei2023elite ; 24jia2023taming ; 16ma2023subject .
+Benefitting from advanced segmentation methods like SAM [73], the SoI can be precisely isolated from the background.
+Based on this strategy, plenty of studies [44, 32, 38, 45, 34, 35, 37, 74, 75] choose to discard the pixels of the background area so that the reconstruction loss can focus on the targeted object and exclude irrelevant disturbances.
+Another technique [49] further adds the masked background reconstruction for better disentanglement.
+In addition, the layout indicated by the pixel mask can be incorporated into the attention modules as a supervision signal [5, 22, 38, 63, 45, 72, 37] to adjust the attention’s concentration adaptively.
+Moreover, the mask can stitch specific feature maps to construct more informative semantic patterns [47, 76, 63].
 Feature-level mask.
 In addition to the pixel-level manipulation, masks can be extended to feature-level operations.
-DisenBooth 18chen2023disenbooth defines an identity-irrelevant embedding with a learnable mask.
+DisenBooth [77] defines an identity-irrelevant embedding with a learnable mask.
 By maximizing the cosine similarity between the identity-preservation embedding and identity-irrelevant embedding, the mask will adaptively exclude the redundant information, and thus the subject appearance can be better preserved.
-AnyDoor 14chen2023anydoor defines a high-frequency mask that stores detailed SoI features as a condition for the image generation process.
-Face-Diffuser 52wang2023high determines the mask through augmentation from the noise predicted by both a pre-trained text-to-image diffusion model and a PTA personalized model.
+AnyDoor [78] defines a high-frequency mask that stores detailed SoI features as a condition for the image generation process.
+Face-Diffuser [79] determines the mask through augmentation from the noise predicted by both a pre-trained text-to-image diffusion model and a PTA personalized model.
 Each model makes its own noise prediction, and the final noise output is a composite created through mask-guided concatenation.
@@
 Some methods enhance data diversity through classical image augmentation like blending and spatial rearrangement.
-SVDiff 8han2023svdiff manually constructs mixed images of multiple SoI as new training data, thereby enhancing the model’s exposure to complex scenarios.
-Such concept composition is also used in other works 16ma2023subject ; 138chen2023videodreamer ; 145wang2024customvideo .
-BLIP-Diffusion 10li2024blip segments the foreground subject and composes it in a random background so that the original text-image pairs are expanded to a larger dataset.
-Besides, StyleAdapter 26wang2023styleadapter chooses to shuffle the image patches to break the irrelevant subject and preserve the desired style.
-PACGen 34li2023generate shows that the spatial position also entangles with the identity information.
-Rescaling, center crop, and relocation are effective solutions for this issue.
+SVDiff [22] manually constructs mixed images of multiple SoI as new training data, thereby enhancing the model’s exposure to complex scenarios.
+Such concept composition is also used in other works [63, 80, 81].
+BLIP-Diffusion [57] segments the foreground subject and composes it in a random background so that the original text-image pairs are expanded to a larger dataset.
+StyleAdapter [82] chooses to shuffle the image patches to break the irrelevant subject and preserve the desired style.
+PACGen [75] shows that the spatial position entangles with the identity information.
+Thus rescaling, center crop, and relocation are effective augmentation solutions.
 Synthetic data.
 Generated synthetic data can provide a large amount of training resources when coupled with quality assurance mechanisms.
-SuTI 11chen2024subject establishes a cascaded pipeline where TTF models first generate diverse variations of each SoI.
+SuTI [52] establishes a cascaded pipeline where TTF models first generate diverse variations of each SoI.
 These synthetic samples then train the target PTA model.
-Similarly, DreamIdentity 89chen2023dreamidentity leverages the existing knowledge of celebrities embedded in large-scale pre-trained diffusion model to generate both the source image and the edited face image.
-StyleDrop 5sohn2024styledrop and GAL 151zhang2024generative implement iterative refinement pipelines where high-quality synthetic outputs from early training phases are incorporated into subsequent rounds.
+Similarly, DreamIdentity [48] leverages the existing knowledge of celebrities embedded in large-scale pre-trained diffusion model to generate both the source image and the edited face image.
+StyleDrop [24] and GAL [19] implement iterative refinement pipelines where high-quality synthetic outputs from early training phases are incorporated into subsequent rounds.
 External sources.
 It is intuitive to leveraging web resources to expand training datasets.
-COTI 3yang2023controllable adopts a scorer network to progressively expand the training set by selecting semantic-relevant samples with high aesthetic quality from a large web-crawled data pool.
+COTI [83] adopts a scorer network to progressively expand the training set by selecting semantic-relevant samples with high aesthetic quality from a large web-crawled data pool.
 4.4 Regularization
@@
 Auxiliary data regularization.
-To mitigate the overfitting problem where the PCS system consistently produces identical outputs as the references, a commonly adopted method is to use an additional dataset composed of images with the same category of the SoI 2ruiz2023dreambooth .
+To mitigate the overfitting problem where the PCS system consistently produces identical outputs as the references, studies start to use an additional dataset composed of images with the same category of the SoI [4].
 By reconstructing these images, the personalized model is required to generate diverse instances of the class while adapting to the target subject.
-Building on this strategy, StyleBoost 53park2023styleboost introduces an auxiliary style-specific auxiliary data to separate content and aesthetic adaptation.
-Later, a dataset 57he2023data is curated with detailed textual prompts (specifying attributes/contexts) to improve disentanglement between subject characteristics and background features.
+Building on this strategy, StyleBoost [84] introduces an auxiliary style-specific data to separate content and aesthetic adaptation.
+Later, a dataset [40] is curated with detailed textual prompts (specifying attributes/contexts) to improve disentanglement between subject characteristics and background features.
 Text embedding constraints.
@@
 By strategically constraining how subject-specific representations interact with established linguistic concepts in the embedding space, these approaches can achieve better generalization ability.
-For example, Perfusion 7tewel2023key constrains key projections toward class noun embeddings while learning value projections from subject images.
-Inspired by coached active learning wei2012coaching ; wei2011coached , which uses anchor concepts for optimization guidance, Compositional Inversion 9zhang2023compositional employs a set of semantically related tokens as anchors to constrain the token embedding search.
-In addition, some works 12gal2023encoder ; 51lu2024object regularize learnable token offsets relative to pre-trained CLIP embeddings.
+For example, Perfusion [21] constrains key projections toward class noun embeddings while learning value projections from subject images.
+Inspired by coached active learning [85, 86], which uses anchor concepts for optimization guidance, Compositional Inversion [87] employs a set of semantically related tokens as anchors to constrain the token embedding search.
+In addition, some works [88, 34] regularize learnable token offsets relative to pre-trained CLIP embeddings.
 By minimizing the offset, the final word embedding is able to achieve better text alignment.
-Similarly, Cones 2 62liu2023cones minimizes the offset by reconstructing the features of 1,000 sentences containing the class noun.
-And 66pang2023cross optimizes the learnable token towards the mean textual embedding of 691 well-known names.
-Domain-Agnostic 17arar2023domain proposes to use a contrastive loss to guide the SoI text embedding close to its nearest CLIP tokens pre-trained on large-scale samples.
-On the other hand, VICO 20hao2023vico empirically finds that the end-of-text token <|EOT|> keeps the semantic consistency of SoI.
+Similarly, Cones 2 [70] minimizes the offset by reconstructing the features of 1,000 sentences containing the class noun.
+And [33] optimizes the learnable token towards the mean textual embedding of 691 well-known names.
+Domain-Agnostic [89] proposes to use a contrastive loss to guide the SoI text embedding close to its nearest CLIP tokens pre-trained on large-scale samples.
+On the other hand, VICO [71] empirically finds that the end-of-text token <|EOT|> keeps the semantic consistency of SoI.
 To leverage this discovery, an L2 loss is leveraged to reduce the difference of attention similarity logits between the SoI token and <|EOT|>.
-5 Categorization of Personalization Tasks
+5 Categorization of Image Personalization Tasks
 As shown in Fig.
 6, personalization covers a range of areas, including objects, styles, faces, etc.
-The following subsections analyze these tasks through the frameworks established in Sec.
-3, examining both TTF and PTA approaches for each domain.
-Figure 6: Visualization of several personalized content synthesis tasks.
+The following subsections analyze these tasks through the frameworks established in Section 3, examining both TTF and PTA approaches for each domain.
+We also summarize these studies in Table 2, Table 3, and Table 4 to provide a clear overview and facilitate quick comparison.
+Figure 6: Visualization of several personalized image synthesis tasks.
 Given text prompts and a few images of the subject of interest, personalization approaches are required to produce expected images.
@@
 An important branch in TTF methods is to optimize the learnable token embeddings.
-The first work started from Textual Inversion 1gal2022image , which applies a simple yet effective method that introduces the unique modifier as a new token to represent the SoI.
+The first work started from Textual Inversion [7], which applies a simple yet effective method that introduces the unique modifier as a new token to represent the SoI.
 One of the significant benefits of this method is its minimal storage requirement as the new tokens consumes just a few kilobytes.
@@
 Recent work aims to address these limitations through several directions.
-DVAR 69voronov2024loss improves training efficiency by proposing a clear stopping criterion by removing all randomness to indicate the convergence.
-For enhanced representation, P+ 23voynov2023p+ introduces distinct learnable tokens across different layers of the U-Net architecture, thereby offering better attribute control through additional learnable parameters.
-NeTI 61alaluf2023neural advances this concept by proposing a neural mapper that adaptively outputs token embeddings based on the denoising timestep and specific U-Net layers.
-ProSpect 30zhang2023prospect recommends optimizing multiple token embeddings tailored to different denoising timesteps based on the observation that different types of prompts, like layout, color, structure, and texture, are activated at different stages of the denoising process.
-Similarly, a study by 79agarwal2023image shows layered activation insight to learn distinct attributes by selectively activating the tokens within their respective scopes.
-Later, HiFiTuner 48wang2023hifi integrates multiple techniques into the learnable token, including mask-guided loss function, parameter regularization, time-dependent embedding, and generation refinement assisted by the nearest reference.
-Alternative approaches, like DreamArtist 68dong2022dreamartist , choose to optimize both negative and positive prompt embeddings to refine the detail preservation.
-In addition to these token-level refinement approaches, the field continues to evolve with novel techniques such as InstructBooth’s reinforcement learning framework 56chae2023instructbooth and gradient-free evolutionary optimization 73fei2023gradient .
+DVAR [90] improves training efficiency by proposing a clear stopping criterion by removing all randomness to indicate the convergence.
+For enhanced representation, P+ [18] introduces distinct learnable tokens across different layers of the U-Net architecture, thereby offering better attribute control through additional learnable parameters.
+NeTI [30] advances this concept by proposing a neural mapper that adaptively outputs token embeddings based on the denoising timestep and specific U-Net layers.
+ProSpect [31] recommends optimizing multiple token embeddings tailored to different denoising timesteps based on the observation that different types of prompts, like layout, color, structure, and texture, are activated at different stages of the denoising process.
+Similarly, a study by [91] shows layered activation insight to learn distinct attributes by selectively activating the tokens within their respective scopes.
+Later, HiFiTuner [32] integrates multiple techniques into the learnable token, including mask-guided loss function, parameter regularization, time-dependent embedding, and generation refinement assisted by the nearest reference.
+Alternative approaches, like DreamArtist [92], choose to optimize both negative and positive prompt embeddings to refine the detail preservation.
+In addition to these token-level refinement approaches, the field continues to evolve with novel techniques such as InstructBooth’s reinforcement learning framework [93] and gradient-free evolutionary optimization [94].
 In summary, recent developments following the foundational work of Textual Inversion focus on reducing training times and enhancing the visual quality of generated images.
 In the realm of TTF methods for PCS, there is a clear shift towards fine-tuning model weights rather than just the token embeddings.
-This approach often addresses the limitations where token embeddings alone struggle to capture complex semantics uncovered in the pre-training data 46gu2024mix ; 80zhao2023catversion .
-DreamBooth 2ruiz2023dreambooth proposes to use a unique modifier by a rare token to represent the SoI and fine-tune the whole parameters of the diffusion model.
+This approach often addresses the limitations where token embeddings alone struggle to capture complex semantics uncovered in the pre-training data [25, 95].
+DreamBooth [4] proposes to use a unique modifier by a rare token to represent the SoI and fine-tune the whole parameters of the diffusion model.
 Besides, a regularization dataset containing 20-30 images with the same category as SoI is adopted to overcome the overfitting problem.
@@
 However, fine-tuning the entire model for each new object causes considerable storage costs, potentially rendering widespread application.
-To address this, Custom Diffusion 6kumari2023multi focuses on identifying and fine-tuning critical parameters, particularly the key-value projections in cross-attention layers, to achieve a balance of visual fidelity and storage efficiency.
-Further approach, Perfusion 7tewel2023key , also adopts the cross-attention fine-tuning and proposes to regularize the update direction of the K (key) projection towards the super-category token embedding and the V (value) projection towards the learnable token embedding.
-COMCAT 75xiao2023comcat introduces a low-rank approximation of attention matrices, which drastically reduces storage requirements to 6 MB while maintaining high fidelity in the outputs.
-Additionally, methods like adapters 72xiang2023closer ; 5sohn2024styledrop and LoRA variants 46gu2024mix ; 58ruiz2023hyperdreambooth ; 115kong2024omg ; 51lu2024object ; 95achlioptas2023stellar ; 76yeh2023navigating are increasingly utilized in personalized generation for parameter-efficient fine-tuning.
+To address this, Custom Diffusion [20] focuses on identifying and fine-tuning critical parameters, particularly the key-value projections in cross-attention layers, to achieve a balance of visual fidelity and storage efficiency.
+Further approach, Perfusion [21], also adopts the cross-attention fine-tuning and proposes to regularize the update direction of the K (key) projection towards the super-category token embedding and the V (value) projection towards the learnable token embedding.
+COMCAT [96] introduces a low-rank approximation of attention matrices, which drastically reduces storage requirements to 6 MB while maintaining high fidelity in the outputs.
+Additionally, methods like adapters [23, 24] and LoRA variants [25, 26, 27, 34, 35, 36] are increasingly utilized in personalized generation for parameter-efficient fine-tuning.
 It is worth noting that the token embedding fine-tuning is compatible with diffusion weight fine-tuning.
-Multiple methods 90hyung2023magicapture ; 74avrahami2023break ; 86hua2023dreamtuner have started using combined weight fine-tuning.
+Multiple methods [37, 38, 39] have started using combined weight fine-tuning.
 PTA framework.
@@
 PTA methods enable real-time generation (less than 10 sec/subject) by leveraging large-scale pre-training to avoid per-subject optimization during the inference phase.
-Re-Imagen 67chen2022re introduces a retrieval-augmented generative approach, which leverages features from text-image pairs retrieved via a specific prompt.
+Re-Imagen [97] introduces a retrieval-augmented generative approach, which leverages features from text-image pairs retrieved via a specific prompt.
 While it is not specifically tailored for object personalization, it demonstrates the feasibility of training reference-conditioned frameworks.
-Later, ELITE 19wei2023elite specifically targets image personalization by combining the global reference features with text embedding while incorporating local features that exclude irrelevant backgrounds.
+Later, ELITE [47] specifically targets image personalization by combining the global reference features with text embedding while incorporating local features that exclude irrelevant backgrounds.
 Both fused features and local features serve as conditions for the denoising process.
-Similarly, InstantBooth 22shi2023instantbooth retrains CLIP models to extract image features and patch features, which are injected into the diffusion model via the attention mechanism and learnable adapter, respectively.
-Additionally, UMM-Diffusion 71ma2023unified designs a multi-modal encoder that produces fused features based on the reference image and text prompt.
+Similarly, InstantBooth [98] retrains CLIP models to extract image features and patch features, which are injected into the diffusion model via the attention mechanism and learnable adapter, respectively.
+Additionally, UMM-Diffusion [61] designs a multi-modal encoder that produces fused features based on the reference image and text prompt.
 The text features and multi-modal hidden state are seen as guidance signals to predict a mixed noise.
-Another work, SuTI 11chen2024subject , adopts the same architecture as Re-Imagen.
+Another work, SuTI [52], adopts the same architecture as Re-Imagen.
 The difference lies in the training samples which are produced by a massive number of TTF models, each tuned on a particular subject set.
 This strategy promotes a more precise alignment with personalization at an instance level rather than the class level of Re-Imagen.
-Moreover, Domain-Agnostic 17arar2023domain combines a contrastive-based regularization technique to push the pseudo embedding produced by the image encoder towards the existing nearest pre-trained token.
+Moreover, Domain-Agnostic [89] combines a contrastive-based regularization technique to push the pseudo embedding produced by the image encoder towards the existing nearest pre-trained token.
 Besides, They introduce a dual-path attention module separately conditioned on the nearest token and pseudo embedding.
 Compared to the methods that use separate encoders to process a single modality, some works have explored the usage of pre-trained multi-modal large language models (MLLM) that can process text and image modality within a unified framework.
-For example, BLIP-Diffusion 10li2024blip utilizes the pre-trained BLIP2 li2023blip that encodes multimodal inputs including the SoI reference and a class noun.
+For example, BLIP-Diffusion [57] utilizes the pre-trained BLIP2 [62] that encodes multimodal inputs including the SoI reference and a class noun.
 The output embedding is then concatenated with context description and serves as a condition to generate images.
-Further, Customization Assistant 84zhou2023customization and KOSMOS-G 77pan2023kosmos replace the text encoder of Stable Diffusion with a pre-trained MLLM to output a fused feature based on the reference and context description.
+Further, Customization Assistant [58] and KOSMOS-G [60] replace the text encoder of Stable Diffusion with a pre-trained MLLM to output a fused feature based on the reference and context description.
 Meanwhile, to meet the standard format of Stable Diffusion, a network is trained to align the dimension of the output embedding.
@@
 TTF framework.
-In this field, StyleDrop 5sohn2024styledrop leverages adapter tuning to efficiently capture the style from a single reference image.
+In this field, StyleDrop [24] leverages adapter tuning to efficiently capture the style from a single reference image.
 This method demonstrates the effectiveness through iterative training, utilizing synthesized images refined by feedback mechanisms, like human evaluations and CLIP scores.
 This approach not only enhances style learning but also ensures that the generated styles align closely with human aesthetic judgments.
-Later, GAL 151zhang2024generative proposes an uncertainty-based evaluation strategy to filter high-quality synthetic-style data and uses a weighted schema to balance the contribution of the additional samples and the original reference.
-Furthermore, StyleAligned 98hertz2023style focuses on maintaining stylistic consistency across a batch of images.
+Later, GAL [19] proposes an uncertainty-based evaluation strategy to filter high-quality synthetic-style data and uses a weighted schema to balance the contribution of the additional samples and the original reference.
+Furthermore, StyleAligned [99] focuses on maintaining stylistic consistency across a batch of images.
 This is achieved by using the first image as a reference, which acts as an additional key and value in the self-attention layers, ensuring that all subsequent images in the batch adhere to the same stylistic guidelines.
-Style-friendly 211choi2024style introduces a novel diffusion model fine-tuning approach that enhances personalized artistic style generation by adaptively biasing noise sampling toward higher noise levels, where stylistic features emerge.
+Style-friendly [100] introduces a novel diffusion model fine-tuning approach that enhances personalized artistic style generation by adaptively biasing noise sampling toward higher noise levels, where stylistic features emerge.
 PTA framework.
-For the PTA framework, StyleAdapter 26wang2023styleadapter employs a dual-path cross-attention mechanism within the PTA framework.
+For the PTA framework, StyleAdapter [82] employs a dual-path cross-attention mechanism within the PTA framework.
 This model introduces a specialized embedding module designed to extract and integrate global features from multiple style references.
-Diptych Prompting 209shin2024large utilizes an inpainting mechanism to draw another image with the same style of the reference part.
+Diptych Prompting [101] utilizes an inpainting mechanism to draw another image with the same style of the reference part.
 5.3 Personalized Face Generation
-Personalized face generation aims to generate diverse ID images that adhere to text prompt specifications, utilizing only a few initial face images.
+Personalized face generation aims to generate diverse identity images that adhere to text prompt specifications, utilizing only a few initial face images.
 Compared to general object personalization, the scope is narrowed to a specific class, humans.
-It is easy to acquire large-scale human-centric datasets zheng2022general ; yang2020facescape ; cao2018vggface2 and utilize pre-trained models in well-developed areas, like face landmark detection wu2019facial ; khabarlak2021fast and face recognition adjabi2020past .
+An obvious benefit is that one can readily leverage large-scale human-centric datasets [67, 102, 103] and utilize pre-trained models in well-developed areas, like face landmark detection [104] and face recognition [105].
 TTF framework.
-As for the TTF methods, PromptNet 55zhou2023enhancing trains a diffusion-based network that encodes the input image and noisy latent to a word embedding.
-To alleviate the overfitting problem, the noises predicted by the word embedding and context description are balanced through fusion sampling through classifier-free guidance.
-HyperDreamBooth 58ruiz2023hyperdreambooth proposes a second-time fine-tuning strategy to further enhance the fidelity after training a PTA model on a large-scale dataset.
-Additionally, Celeb Basis 60yuan2023inserting provides a novel idea that the personalized ID can be viewed as the composition of celebrity’s face which has been learned by the pre-trained diffusion model.
-Based on this hypothesis, a simple MLP is optimized to transform face features into the celeb embedding space.
+Regarding TTF methods, PromptNet [106] trains a diffusion-based network that encodes the input image and noisy latent to a word embedding.
+To alleviate the overfitting problem, the noises predicted by the word embedding and context description are balanced through fusion sampling in classifier-free guidance.
+Additionally, Celeb Basis [65] provides a novel idea that the personalized ID can be viewed as the composition of celebrity’s face, which has been learned by the pre-trained diffusion model.
+Based on this hypothesis, a simple MLP is optimized at test time to transform face features into the weighting of different celebrity name embeddings.
 PTA framework.
 Thanks to the abundance of available datasets featuring the same individual in various contexts, which provide valuable data for pre-training PTA methods, the number of works in PTA frameworks is rapidly increasing.
-Face0 88valevski2023face0 crops the face region to extract refined embeddings and concatenates them with text features.
+Face0 [107] crops the face region to extract refined embeddings and concatenates them with text features.
 During the sampling phase, the output of classifier-free guidance is replaced by a weighted combination of the noise patterns predicted by face-only embedding, text-only embedding, and concatenated face-text embedding.
-The $\mathcal{W}+$ Adapter 91li2023stylegan constructs a mapping network and residual cross-attention modules to transform the facial features from the StyleGAN karras2019style $\mathcal{W}+$ space into the text embedding space of Stable Diffusion.
-FaceStudio 31yan2023facestudio adapts the cross-attention layer to support hybrid guidance including stylized images, facial images, and textual prompts.
-Moreover, PhotoMaker 28li2023photomaker constructs a high-quality dataset through a meticulous data collection and filtering pipeline.
+The $\mathcal{W}+$ Adapter [53] constructs a mapping network and residual cross-attention modules to transform the facial features from the StyleGAN [108] $\mathcal{W}+$ space into the text embedding space of Stable Diffusion.
+FaceStudio [109] adapts the cross-attention layer to support hybrid guidance including stylized images, facial images, and textual prompts.
+Moreover, PhotoMaker [46] constructs a high-quality dataset through a meticulous data collection and filtering pipeline.
 They use a two-layer MLP to fuse ID features and class embeddings for an overall representation of human portrait.
-PortraitBooth 96peng2023portraitbooth also employs a simple MLP, which fuses the text condition and shallow features of a pre-trained face recognition model.
+PortraitBooth [50] also employs a simple MLP, which fuses the text condition and shallow features of a pre-trained face recognition model.
 To ensure expression manipulation and facial fidelity, they add another expression token and incorporate the identity preservation loss and mask-based cross-attention loss.
-InstantID 37wang2024instantid additionally introduces a variant of ControlNet that takes facial landmarks as input, providing stronger guiding signals compared to the methods that solely rely on attention fusion.
+InstantID [6] additionally introduces a variant of ControlNet that takes facial landmarks as input, providing stronger guiding signals compared to the methods that solely rely on attention fusion.
 5.4 Multiple Subject Composition
@@
 Some works focus on the one-for-one generation following a fusion mechanism.
-For instance, Custom Diffusion 6kumari2023multi proposes a constrained optimization method to merge the cross-attention key-value projection weights with the goal of maximizing reconstruction performance for each subject.
-Mix-of-Show 46gu2024mix fuses the LoRA hu2021lora weights with the same optimization objective.
-StyleDrop 5sohn2024styledrop dynamically summarizes noise predictions from each personalized diffusion model.
-In OMG 115kong2024omg , the latent predicted by each LoRA-tuned model is spatially composited using the subject mask.
+For instance, Custom Diffusion [20] proposes a constrained optimization method to merge the cross-attention key-value projection weights with the goal of maximizing reconstruction performance for each subject.
+Mix-of-Show [25] fuses the LoRA [110] weights with the same optimization objective.
+StyleDrop [24] dynamically summarizes noise predictions from each personalized diffusion model.
+In OMG [27], the latent predicted by each LoRA-tuned model is spatially composited using the subject mask.
 Joint training is another strategy to cover all expected subjects.
-SVDiff 8han2023svdiff employs a data augmentation method called Cut-Mix to compose several subjects together and applies a location loss to regularize attention maps, ensuring alignment between each subject and its corresponding token.
-Similar strategies are found in other works 6kumari2023multi ; 60yuan2023inserting which train a single model by reconstructing the appearance of every SoI.
+SVDiff [22] employs a data augmentation method called Cut-Mix to compose several subjects together and applies a location loss to regularize attention maps, ensuring alignment between each subject and its corresponding token.
+Similar strategies are found in other works [20, 65] which train a single model by reconstructing the appearance of every SoI.
 There are advanced control mechanisms designed to manage multiple subjects.
-Cones 4liu2023cones proposes to find a small cluster of neurons that preserve the most information about SoI.
+Cones [111] proposes to find a small cluster of neurons that preserve the most information about SoI.
 The neurons belonging to different SoI will be simultaneously activated to generate the combination.
-Compositional Inversion 9zhang2023compositional introduces spatial region assignment to different subjects to improve the composition success rate.
+Compositional Inversion [87] introduces spatial region assignment to different subjects to improve the composition success rate.
 PTA framework.
 For PTA frameworks, multi-subject generation is achieved through specialized architectural designs.
-Fastcomposer 54xiao2023fastcomposer , Subject-Diffusion 16ma2023subject , and $\lambda$-eclipse 222patel2024lambda place each subject feature in its corresponding placeholder in the text embedding, ensuring a seamless and efficient combination.
-CustomNet 29yuan2023customnet and MIGC 174zhou2024migc train a PTA network that supports location control for each subject.
-SSR-Encoder 221zhang2024ssr implements an encoder to selectively preserve the desired subject feature and a cross-attention module to support multi-subject feature fusion.
+Fastcomposer [5], Subject-Diffusion [63], and $\lambda$-eclipse [112] place each subject feature in its corresponding placeholder in the text embedding, ensuring a seamless and efficient combination.
+CustomNet [113] and MIGC [114] train a PTA network that supports location control for each subject.
+SSR-Encoder [55] implements an encoder to selectively preserve the desired subject feature and a cross-attention module to support multi-subject feature fusion.
 5.5 High-level Semantic Personalization
@@
 For now, all researches in this field are based on the TTF framework.
-ReVersion 63huang2023reversion intends to invert object relations from references.
+ReVersion [115] intends to invert object relations from references.
 Specifically, they use a contrastive loss to guide the optimization of the token embedding towards specific clusters of Part-of-Speech tags, such as prepositions, nouns, and verbs.
 Meanwhile, they also increase the likelihood of adding noise at larger timesteps during the training process to emphasize the extraction of high-level semantic features.
-Lego 45motamed2023lego focuses on more general concepts, such as adjectives, which are frequently intertwined with the subject appearance.
+Lego [116] focuses on more general concepts, such as adjectives, which are frequently intertwined with the subject appearance.
 The concept can be learned from a contrastive loss applied to the dataset comprising clean subject images and images that embody the desired adjectives.
-Moreover, ADI 43huang2023learning aims to learn the action-specific identifier from the references.
+Moreover, ADI [117] aims to learn the action-specific identifier from the references.
 To ensure the inversion only focuses on the desired action, ADI extracts gradient invariance from a constructed triplet sample and applies a threshold to mask out the irrelevant feature channels.
@@
 Recent personalization tasks tend to include additional conditions for diverse content customization.
-One common application is to condition on a fixed source image.
-For example, PhotoSwap 15gu2024photoswap introduces a new task that replaces the subject in the source image with the SoI from the reference image.
+One common application is to customize the subject into a fixed source image.
+For example, PhotoSwap [118] introduces a new task that replaces the subject in the source image with the SoI from the reference image.
 To address this requirement, they first fine-tunes a diffusion model on the references to obtain a personalized model.
-To preserve the background of the source image, they initialize the noise with DDIM inversion song2020denoising and replace the intermediate feature maps with those derived from source image generation during inference time.
-Later, MagiCapture 91li2023stylegan broadens the scope to face customization.
+To preserve the background of the source image, they initialize the noise with DDIM inversion [119] and replace the intermediate feature maps with those derived from source image generation during inference time.
+Later, MagiCapture [53] broadens the scope to face customization.
 Another similar application can be found in Virtual Try-on, which aims to fit selected clothing onto a target person.
-The complexities of this task have been thoroughly analyzed in another survey islam2024deep .
-Additional conditions in personalization tasks may include adjusting the layout 152chen2024training , transforming sketches 94chen2023democaricature , controlling viewpoint 29yuan2023customnet ; 163kumari2024customizing , or modifying poses 37wang2024instantid .
+The complexities of this task have been thoroughly analyzed in another survey [120].
+Additional conditions in personalization tasks may include adjusting the layout [69], transforming sketches [121], controlling viewpoint [113, 122], or modifying poses [6].
 Each of these conditions presents unique challenges and requires specialized approaches to integrate these elements seamlessly into the personalized content.
 6.2 Personalized Video Generation
-With the rising popularity of video generation xing2024survey , video personalization has also begun to attract attention.
+With the rising popularity of video generation [123], video personalization has also begun to attract attention.
 In video personalization, the SoI can be classified into three distinct categories: appearance, motion, and the combination of both appearance and motion.
@@
 The standard TTF pipeline utilizes reference images as appearance anchors and fine-tunes video diffusion models (VDM) for temporal synthesis.
-The process involves leveraging sophisticated methods from 2D personalization, such as parameter-efficient fine-tuning 145wang2024customvideo , data augmentation 138chen2023videodreamer ; 145wang2024customvideo , and attention manipulation 83jiang2023videobooth ; 212liu2024stylecrafter ; 145wang2024customvideo .
-Additionally, several studies 82zhao2023videoassembler ; 83jiang2023videobooth ; 212liu2024stylecrafter ; 105feng2023dreamoving have explored the PTA framework.
+The process involves leveraging sophisticated methods from 2D personalization, such as parameter-efficient fine-tuning [81], data augmentation [80, 81], and attention manipulation [124, 54, 81].
+Additionally, several studies [125, 124, 54, 126] have explored the PTA framework.
 These diffusion models are specifically tailored to synthesize videos based on the image references.
@@
 In this task, the reference input switches to a video clip containing a consistent action.
-A common approach is to fine-tune the video diffusion model by reconstructing the action clip 150wu2023tune ; 103song2023save ; 140jeong2023vmc ; 148esser2023structure ; 136he2023animate ; 102wu2023lamp ; 104materzynska2023customizing .
+A common approach is to fine-tune the video diffusion model by reconstructing the action clip [127, 128, 129, 130, 131, 132, 133].
 However, distinguishing between appearance and motion within the reference video can be challenging.
-To solve this problem, SAVE 103song2023save applies appearance learning to ensure that appearance is excluded from the motion learning phase.
-Additionally, VMC 140jeong2023vmc removes the background information during training prompt construction.
+To solve this problem, SAVE [128] applies appearance learning to ensure that appearance is excluded from the motion learning phase.
+Additionally, VMC [129] removes the background information during training prompt construction.
 Appearance and motion personalization.
 When integrating both subject appearance and motion, innovative methods are employed to address the complexities of learning both aspects simultaneously.
-MotionDirector 101zhao2023motiondirector utilizes spatial and temporal losses to facilitate learning across these dimensions.
-Another approach, DreamVideo 142wei2023dreamvideo , incorporates residual features from a randomly selected frame to emphasize subject information.
+MotionDirector [134] utilizes spatial and temporal losses to facilitate learning across these dimensions.
+Another approach, DreamVideo [135], incorporates residual features from a randomly selected frame to emphasize subject information.
 This technique enables the fine-tuned module to primarily focus on learning motion dynamics.
@@
 Basically, the pipeline begins by fine-tuning a 2D diffusion model using TTF methods.
-This tuned model then guides the optimization of a 3D Neural Radiance Field (NeRF) model mildenhall2021nerf for each specific prompt 120lin2023magic3d ; 126shi2023mvdream .
-In the second phase, DreamFusion poole2022dreamfusion is the key technique, which introduces Score Distillation Sampling (SDS) to train a 3D model capable of rendering images aligned with 2D diffusion model.
+This tuned model then utilize Score Distillation Sampling (SDS) [136] to train a 3D Neural Radiance Field (NeRF) model [137] for each specific prompt [138, 139].
 Building on this foundation, several methods have been developed to improve the workflow.
-DreamBooth3D 121raj2023dreambooth3d structures the process into three phases: initializing and optimizing a NeRF from a DreamBooth model, rendering multi-view images, and fine-tuning a secondary DreamBooth for the final 3D NeRF refinement.
-Consist3D 127ouyang2023chasing enhances text embeddings by training two distinct tokens, a semantic and a geometric token, during 3D model optimization.
-TextureDreamer 133yeh2024texturedreamer focuses on extracting texture maps from optimized spatially-varying bidirectional reflectance distribution (BRDF) fields for rendering texture on wide-range 3D subjects.
+DreamBooth3D [140] structures the process into three phases: initializing and optimizing a NeRF from a DreamBooth model, rendering multi-view images, and fine-tuning a secondary DreamBooth for the final 3D NeRF refinement.
+Consist3D [141] enhances text embeddings by training two distinct tokens, a semantic and a geometric token, during 3D model optimization.
+TextureDreamer [142] focuses on extracting texture maps from optimized spatially-varying bidirectional reflectance distribution (BRDF) fields for rendering texture on wide-range 3D subjects.
 Additionally, advancements extend to 3D avatar rendering and dynamic scenes.
-Animate124 128zhao2023animate124 and Dream-in-4D 130zheng2023unified integrate video diffusion for 4D dynamic scene support within the 3D optimization process.
-In avatar rendering, PAS 122azadi2023text generates 3D body poses configurable by avatar settings, StyleAvatar3D 123zhang2023styleavatar3d facilitates 3D avatar generation based on images, and AvatarBooth 124zeng2023avatarbooth employs dual fine-tuned diffusion models for separate face and body generation.
+Animate124 [143] and Dream-in-4D [144] integrate video diffusion for 4D dynamic scene support within the 3D optimization process.
+In avatar rendering, PAS [145] generates 3D body poses configurable by avatar settings, StyleAvatar3D [146] facilitates 3D avatar generation based on images, and AvatarBooth [147] employs dual fine-tuned diffusion models for separate face and body generation.
 6.4 Attack and Defense
@@
 There is an increased risk that individuals or organizations may exploit them to produce misleading content or manipulate public perception.
-To mitigate this, Anti-DreamBooth 13van2023anti aims to add a subtle noise perturbation to the references so that any personalized model trained on these samples only produces terrible results.
+To mitigate this, Anti-DreamBooth [148] aims to add a subtle noise perturbation to the references so that any personalized model trained on these samples only produces terrible results.
 The basic idea is to maximize the reconstruction loss of the surrogate model.
-Additionally, 50zhang2023backdooring suggests to predefine a collection of trigger words and meaningless images.
+Additionally, [149] suggests to predefine a collection of trigger words and meaningless images.
 These data are paired and incorporated during the training phase.
@@
 Several works are exploring different personalization extensions.
-For example, SVG personalization is introduced by 36zhang2023text , in which a parameter-efficient fine-tuning method is applied to create SVGs.
+For example, SVG personalization is introduced by [150], in which a parameter-efficient fine-tuning method is applied to create SVGs.
 After first-step generation, the SVGs are refined through a process that includes semantic alignment and a dual optimization approach, which utilizes both image-level and vector-level losses to enhance the final output.
-Another application, 360-degree panorama customization 25wang2024customizing , is also emerging as a potential tool for personalization in the digital imaging realm.
+Another application, 360-degree panorama customization [151], is also emerging as a potential tool for personalization in the digital imaging realm.
 7 Evaluation
@@
 To assess the performance of personalized models, various datasets have been developed:
-DreamBench 2ruiz2023dreambooth serves as the primary evaluation benchmark for DreamBooth 2ruiz2023dreambooth , containing 30 diverse subjects (e.g., backpacks, animals, vehicles, and toys) with 25 unique prompts per subject.
-DreamBench-v2 11chen2024subject expands the evaluation scope of DreamBench by adding 220 test prompts for each subject.
-Custom-10 6kumari2023multi used in Custom Diffusion 6kumari2023multi evaluates 10 subjects, each with 20 specific test prompts, and includes tests for multi-subject composition with 5 pairs of subjects and 8 prompts for each pair.
-Custom-101 6kumari2023multi is the latest released dataset by the authors of Custom Diffusion 6kumari2023multi , which comprises 101 subjects to provide a broader scope of evaluation.
-Stellar 95achlioptas2023stellar specifically targets human-centric evaluation, featuring 20,000 prompts on 400 human identities.
+DreamBench [4] serves as the primary evaluation benchmark for DreamBooth [4], containing 30 diverse subjects (e.g., backpacks, animals, vehicles, and toys) with 25 unique prompts per subject.
+DreamBench-v2 [52] expands the evaluation scope of DreamBench by adding 220 test prompts for each subject.
+Custom-10 [20] used in Custom Diffusion [20] evaluates 10 subjects, each with 20 specific test prompts, and includes tests for multi-subject composition with 5 pairs of subjects and 8 prompts for each pair.
+Custom-101 [20] is the latest released dataset by the authors of Custom Diffusion [20], which comprises 101 subjects to provide a broader scope of evaluation.
+Stellar [35] specifically targets human-centric evaluation, featuring 20,000 prompts on 400 human identities.
 Despite these contributions, they remain fragmented across different research groups and the research community still lacks a benchmark tested on a large number of personalized generation tasks.
@@
 The text alignment metrics quantify how precisely generated outputs reflect prompt semantics:
-CLIP-T measures semantic alignment using the cosine similarity between CLIP radford2021learning embeddings of generated images and their text prompts.
-ImageReward xu2023imagereward , HPS Score (v1/v2) wu2023human ; wu2023humanv2 , and PickScore kirstain2023pick employ learned models trained on human judgments to better correlate with perceptual quality to reflect human preference better.
+CLIP-T measures semantic alignment using the cosine similarity between CLIP [42] embeddings of generated images and their text prompts.
+ImageReward [152], HPS Score (v1/v2) [153, 154], and PickScore [155] employ learned models trained on human judgments to better correlate with perceptual quality to reflect human preference better.
 To determine how closely the generated subject resembles the SoI, visual fidelity can be assessed via the following metrics:
 CLIP-I evaluates subject preservation through CLIP image embedding similarity between generations and references.
-Optimal values balance fidelity (high scores) against overfitting (excessive scores that ignore text guidance).
-DINO-I caron2021emerging provides complementary assessment using DINO’s instance-aware features, particularly effective for object-level similarity.
-Fréchet Inception Distance (FID) FIDheusel2017gans quantifies the statistical similarity between generated and real image distributions through Inception-V3.
-Inception Score (IS) ISsalimans2016improved evaluates both image quality and diversity through classifier confidence.
-In addition to these commonly adopted metrics, there are some discussions of specialized metrics for PCS system evaluation: LyCORIS 76yeh2023navigating introduces a 5-dimensional assessment covering fidelity, controllability, diversity, base model preservation, and image quality.
-Stellar 95achlioptas2023stellar develops six human-centric metrics including soft-penalized CLIP text score, Identity Preservation Score, Attribute Preservation Score, Stability of Identity Score, Grounding Objects Accuracy, and Relation Fidelity Score.
+Optimal values should balance fidelity (high scores) against overfitting (excessive scores that ignore text guidance).
+DINO-I [156] provides a complementary assessment using DINO’s instance-aware features, particularly effective for object-level similarity.
+Fréchet Inception Distance (FID) [157] quantifies the statistical similarity between generated and real image distributions through Inception-V3 [158].
+In addition to these commonly adopted metrics, there are some discussions of specialized metrics for PCS system evaluation: LyCORIS [36] introduces a 5-dimensional assessment covering fidelity, controllability, diversity, base model preservation, and image quality.
+Stellar [35] develops six human-centric metrics including soft-penalized CLIP text score, Identity Preservation Score, Attribute Preservation Score, Stability of Identity Score, Grounding Objects Accuracy, and Relation Fidelity Score.
 These metrics ensure a structured and detailed evaluation of personalized models.
@@
 7.3 New Benchmark on SoTA Methods
-Although multiple evaluation datasets have been proposed in this area, there is still an urgent need for a standardized benchmark that systematically assesses performance across diverse PCS methodologies.To address this limitation, we present a comprehensive evaluation built upon existing works.
+Although multiple evaluation datasets have been proposed in this area, there is still an urgent need for a standardized benchmark that systematically assesses performance across diverse PCS methodologies.
+To address this limitation, we present a new test dataset called Persona with a comprehensive evaluation built upon existing works.
 For Object.
-We collect a dataset of 47 subjects from available resources 1gal2022image ; 2ruiz2023dreambooth ; 6kumari2023multi .
-Following the methodology of DreamBooth 2ruiz2023dreambooth , we categorize the subjects into two classes: objects and live pets, based on whether the subject is a living entity.
+Persona includes 47 subjects from available resources [7, 4, 20].
+Following the methodology of DreamBooth [4], we categorize the subjects into two classes: objects and live pets, based on whether the subject is a living entity.
 Specifically, 10 out of 47 subjects are pets and the remaining 37 are various objects.
-To evaluate the performance, we utilize the text prompts from DreamBooth 2ruiz2023dreambooth .
+To evaluate the performance, we utilize the text prompts from DreamBooth [4].
 This includes 20 recontextualization prompts and 5 property modification prompts for objects, along with 10 recontextualization, 10 accessorization, and 5 property modification prompts for pets, totally 25 prompts per category.
-For Face. We collect 15 subjects from Celeb-A liu2018large dataset. We use 40 prompts for evaluation, including 10 accessory prompts, 10 style prompts, 10 action prompts, and 10 context prompts.
+For Face.
+We also collected 15 subjects from Celeb-A [159] into our Persona dataset.
+We use 40 prompts for evaluation, including 10 accessory prompts, 10 style prompts, 10 action prompts, and 10 context prompts.
 Settings.
@@
 We calculate CLIP-I for subject fidelity in object generation.
-For subject fidelity in face generation, we detect faces in both the generated and target images using MTCNN zhang2016mtcnn and calculate the pairwise identity similarity using FaceNet schroff2015facenet .
+For subject fidelity in face generation, we detect faces in both the generated and target images using MTCNN [160] and calculate the pairwise identity similarity using FaceNet [161].
 To uniformly evaluate and compare existing personalized generation methods, we select 22 representative PCS methods for evaluation.
@@
 Results.
-The evaluation results are shown in Tab.1.
+The evaluation results are shown in Table 1.
 It is evident that no method excels simultaneously in both visual fidelity and text alignment metrics.
@@
 Consequently, the visual similarity metric yields higher results based on the mirrored output and the reference, rather than providing an accurate representation of the performance.
-Table 1: Evaluation results of representative methods on our evaluation dataset.
+Table 1: Evaluation results of representative methods on our Persona evaluation dataset.
 Type Methods Framework Backbone CLIP-T CLIP-I
-Object Textual Inversion 1gal2022image TTF SD 1.5 0.199 0.749
-Dreambooth 2ruiz2023dreambooth TTF SD 1.5 0.286 0.772
-P+ 23voynov2023p+ TTF SD 1.4 0.244 0.643
-Custom Diffusion 6kumari2023multi TTF SD 1.4 0.307 0.722
-NeTI 61alaluf2023neural TTF SD 1.4 0.283 0.801
-SVDiff 8han2023svdiff TTF SD 1.5 0.282 0.776
-Perfusion 7tewel2023key TTF SD 1.5 0.273 0.691
-ELITE 19wei2023elite PTA SD 1.4 0.292 0.765
-BLIP-Diffusion 10li2024blip PTA SD 1.5 0.292 0.772
-IP-Adapter 40ye2023ip PTA SD 1.5 0.2722 0.825
-SSR Encoder 221zhang2024ssr PTA SD 1.5 0.288 0.792
-MoMA song2024moma PTA SD 1.5 0.322 0.748
-Diptych Prompting 209shin2024large PTA FLUX 1.0 dev 0.327 0.722
-$\lambda$-eclipse 222patel2024lambda PTA Kandinsky 2.2 0.272 0.824
-MS-Diffusion 220personalizationms PTA SDXL 0.298 0.777
-Face CrossInitialization 66pang2023cross TTF SD 2.1 0.261 0.469
-Face2Diffusion 170shiohara2024face2diffusion PTA SD 1.4 0.265 0.588
-SSR Encoder 221zhang2024ssr PTA SD 1.5 0.233 0.490
-FastComposer 54xiao2023fastcomposer PTA SD 1.5 0.230 0.516
-IP-Adapter 40ye2023ip PTA SD 1.5 0.292 0.462
-IP-Adapter 40ye2023ip PTA SDXL 0.292 0.642
-PhotoMaker 28li2023photomaker PTA SDXL 0.311 0.547
-InstantID 37wang2024instantid PTA SDXL 0.278 0.707
+Object Textual Inversion [7] TTF SD 1.5 0.199 0.749
+Dreambooth [4] TTF SD 1.5 0.286 0.772
+P+ [18] TTF SD 1.4 0.244 0.643
+Custom Diffusion [20] TTF SD 1.4 0.307 0.722
+NeTI [30] TTF SD 1.4 0.283 0.801
+SVDiff [22] TTF SD 1.5 0.282 0.776
+Perfusion [21] TTF SD 1.5 0.273 0.691
+ELITE [47] PTA SD 1.4 0.292 0.765
+BLIP-Diffusion [57] PTA SD 1.5 0.292 0.772
+IP-Adapter [51] PTA SD 1.5 0.2722 0.825
+SSR Encoder [55] PTA SD 1.5 0.288 0.792
+MoMA [59] PTA SD 1.5 0.322 0.748
+Diptych Prompting [101] PTA FLUX 1.0 dev 0.327 0.722
+$\lambda$-eclipse [112] PTA Kandinsky 2.2 0.272 0.824
+MS-Diffusion [56] PTA SDXL 0.298 0.777
+Face CrossInitialization [33] TTF SD 2.1 0.261 0.469
+Face2Diffusion [162] PTA SD 1.4 0.265 0.588
+SSR Encoder [55] PTA SD 1.5 0.233 0.490
+FastComposer [5] PTA SD 1.5 0.230 0.516
+IP-Adapter [51] PTA SD 1.5 0.292 0.462
+IP-Adapter [51] PTA SDXL 0.292 0.642
+PhotoMaker [46] PTA SDXL 0.311 0.547
+InstantID [6] PTA SDXL 0.278 0.707
 8 Challenge and Outlook
 8.1 Overfitting Problem
-As discussed in Sec.
-7.3, current PCS systems face a critical challenge of overfitting because of a limited set of reference images.
+As discussed in Section 7.3, current PCS systems face a critical challenge of overfitting because of a limited set of reference images.
 This overfitting problem manifests in two ways: 1) Loss of SoI editability.
@@
 The irrelevant elements in the references are generated in the output, such as backgrounds or objects not pertinent to the current context.
-To investigate the rationale behind, Compositional Inversion 9zhang2023compositional observes that the learned token embedding is located in an out-of-distribution area compared to the center distribution formed by pre-trained words.
-This is also found in another work 66pang2023cross that the learnable token embeddings deviate significantly from the distribution of the initial embedding.
-In addition, there is evidence 9zhang2023compositional ; 24jia2023taming ; 20hao2023vico suggesting that the unique modifier dominates in the cross-attention layers compared to the other context tokens, leading to the absence of other semantic appearances.
+To investigate the rationale behind, Compositional Inversion [87] observes that the learned token embedding is located in an out-of-distribution area compared to the center distribution formed by pre-trained words.
+This is also found in another work [33] that the learnable token embeddings deviate significantly from the distribution of the initial embedding.
+In addition, there is evidence [87, 76, 71] suggesting that the unique modifier dominates in the cross-attention layers compared to the other context tokens, leading to the absence of other semantic appearances.
 To address this issue, many solutions have been proposed.
-Most methods discussed in Sec.
-4 contribute to the alleviation of the overfitting problem, such as the exclusion of redundant background, attention manipulation, regularization of the learnable parameters, and data augmentation.
-However, it has not been solved yet, especially in the cases where the SoI has a non-rigid appearance 9zhang2023compositional or the context prompt has a similar semantic correlation with the irrelevant elements in the reference 151zhang2024generative .
+Most methods discussed in Section 4 contribute to the alleviation of the overfitting problem, such as the exclusion of redundant background, attention manipulation, regularization of the learnable parameters, and data augmentation.
+However, it has not been solved yet, especially in the cases where the SoI has a non-rigid appearance [87] or the context prompt has a similar semantic correlation with the irrelevant elements in the reference [19].
 It is clear that addressing overfitting in PCS is not merely a technical challenge but a necessity for ensuring the practical deployment and scalability of these systems in varied and dynamic real-world environments.
@@
 As a result, it becomes challenging to achieve flexible adaptation in different contexts while simultaneously pushing the model to capture fine-grained details.
-To address this inherent conflict, Perfusion 7tewel2023key proposes to regularize the attention projections by these two items.
-35cao2023concept decouples the conditional guidance into two separate processes, which allows for the distinct handling of subject fidelity and textual alignment.
+To address this inherent conflict, Perfusion [21] proposes to regularize the attention projections by these two items.
+[163] decouples the conditional guidance into two separate processes, which allows for the distinct handling of subject fidelity and textual alignment.
 Despite these efforts, there still remains room for further exploration and refinement of this issue.
@@
 Recent advancements in multimodal autoregressive models present novel solutions for PCS by unifying cross-modal understanding and generation.
-Models like Emu3 wang2024emu3 demonstrate that autoregressive architectures can natively handle image-text-video sequences through discrete tokenization and joint transformer training.
+Models like Emu3 [164] demonstrate that autoregressive architectures can natively handle image-text-video sequences through discrete tokenization and joint transformer training.
 This paradigm enables seamless integration of user-provided multimodal references (e.g., text descriptions paired with SoI images) while maintaining contextual coherence across generation steps.
-Besides, this framework natively support for subject editing via multi-round chat, effectively addressing the overfitting limitations observed in diffusion-based PCS.
+Besides, this framework natively support for subject editing via multi-round chat, effectively addressing the overfitting limitations commonly observed in diffusion-based models.
 8.5 Interactive Personalization Workflow
-The evolution of interactive generation systems presents valuable opportunities for PCS, particularly in addressing core challenges like the overfitting problem while enhancing user control.
-Recent advancements in conversational PCS systems, such as Gemini-2.0-flash’s interactive image editing team2023gemini , demonstrate how iterative refinement through natural language can optimize both subject fidelity and prompt alignment.
-This aligns with PCS’s fundamental goal of balancing precise subject representation with flexible editing capability.
+The evolution of interactive generation systems has unlocked new frontiers for PCS, particularly through the integration of multi-round interactive generation.
+This capability allows users to iteratively refine and accurately define the SoI, addressing the challenge of translating vague or complex requirements into precise content generation.
+For instance, conversational PCS systems like Gemini-2.0-flash [165] exemplify this progress, leveraging natural language dialogue to iteratively optimize both subject fidelity and prompt alignment.
+By enabling users to provide real-time feedback and adjust parameters through chat-like interactions, these systems bridge the gap between abstract intent and concrete outputs, aligning with PCS’s core objective of balancing faithful subject representation with flexible editability.
 9 Conclusion
@@
 Paper Scope Framework Backbone Technique
-Textual Inversion 1gal2022image Object TTF LDM Token Embedding Fine-tuning
-DreamBooth 2ruiz2023dreambooth Object TTF Imagen Diffusion Model Fine-tuning; Regularization Dataset
-DreamArtist 68dong2022dreamartist Object TTF LDM Negative Prompt Fine-tuning
-DVAR 69voronov2024loss Object TTF SD 1.5 Randomness Erasing
-HiPer 33han2023highly Object TTF SD Token Embedding Enhancement
-P+ 23voynov2023p+ Object TTF SD 1.4 Token Embedding Enhancement
-Unet-finetune 72xiang2023closer Object TTF SD Parameter-efficient Fine-tuning
-Jia et al. 24jia2023taming Object TTF Imagen Mask-assisted Generation; Regularization
-COTI 3yang2023controllable Object TTF SD 2.0 Data Augmentation
-Gradient-Free TI 73fei2023gradient Object TTF SD Evolutionary Strategy
-PerSAM 21zhang2023personalize Object TTF SD Mask-assisted Generation
-DisenBooth 18chen2023disenbooth Object TTF SD 2.1 Regularization
-NeTI 61alaluf2023neural Object TTF SD 1.4 Token Embedding Enhancement
-Prospect 30zhang2023prospect Object TTF SD 1.4 Token Embedding Enhancement
-Break-a-Scene 74avrahami2023break Object TTF SD 2.1 Attention-based Operation; Attention-based Operation; Mask-assisted Generation
-COMCAT 75xiao2023comcat Object TTF SD 1.4 Attention-based Operation
-ViCo 20hao2023vico Object TTF SD Attention-based Operation; Regularization
-OFT 59qiu2024controlling Object TTF SD 1.5 Finetuning Strategy
-LyCORIS 76yeh2023navigating Object TTF SD 1.5 Attention-based Operation
-He et al. 57he2023data Object TTF SD, SDXL Data Augmentation
-DIFFNAT 78roy2023diffnat Object TTF SD Regularization
-MATTE 79agarwal2023image Object; Style; High-level Semantic TTF SD Regularization
-LEGO 45motamed2023lego Object TTF SD Regularization
-Catversion 80zhao2023catversion Object TTF SD 1.5 Embedding Concatenation
-CLiC 81safaee2023clic Object TTF SD 1.4 Attention-based Operation; Mask-assisted Generation
-HiFi Tuner 48wang2023hifi Object TTF SD 1.4 Attention-based Operation; Mask-assisted Generation; Regularization
-InstructBooth 56chae2023instructbooth Object TTF SD 1.5 Reinforcement Learning
-DETEX 85cai2023decoupled Object TTF SD 1.4 Mask-assisted Generation
-DreamDistribution 114zhao2023dreamdistribution Object TTF SD 2.1 Token Embedding Enhancement
-DreamTuner 86hua2023dreamtuner Object TTF SD Attention-based Operation
-Xie et al. 51lu2024object Object TTF SD Regularization; Mask-assisted Generation
-Infusion 157zeng2024infusion Object TTF SD 1.5 Attention-based Operation
-Pair Customization 160jones2024customizing Object TTF SDXL Disentanglement Approach
-DisenDreamer 171chen2024disendreamer Object TTF SD 2.1 Disentanglement Approach
-DreamBooth+++++ + 219fan2024dreambooth Object TTF SD Attention-based Operation
-CoRe 218wu2024core Object TTF SD 2.1 Regularization
-DCI 216jin2025customized Object TTF SD Token Embedding Optimization
-Re-Imagen 67chen2022re Object PTA Imagen Retrieval-augmented Paradigm
-Versatile Diffusion 109xu2023versatile Object PTA SD 1.4 Architecture Design
-Tuning-Encoder 12gal2023encoder Object PTA SD Regularization
-ELITE 19wei2023elite Object PTA SD 1.4 Attention-based Operation; Mask-assisted Generation
-UMM-Diffusion 71ma2023unified Object PTA SD 1.5 Reference Feature Injection
-SuTI 11chen2024subject Object PTA Imagen Data Augmentation
-InstantBooth 22shi2023instantbooth Object PTA SD 1.4 Patch Feature Extraction
-BLIP-Diffusion 10li2024blip Object PTA SD 1.5 Data Augmentation; Mask-assisted Generation
-Domain-Agnostic 17arar2023domain Object PTA SD Regularization; Attention-based Operation
-IP-Adapter 40ye2023ip Object PTA SD 1.5 Reference Feature Injection
-Kosmos-G 77pan2023kosmos Object; Face PTA SD 1.5 Multimodal Language Modeling
-Kim et al. 47kim2023personalized Object PTA SD 1.4 Reference Feature Injection
-CAFÉ 84zhou2023customization Object; Face PTA SD 2.0 Multimodal Language Modeling
-SAG 99pan2023towards Object; Face; Style PTA SD 1.5 Gradient-based Guidance
-BootPIG 41purushwalkam2024bootpig Object PTA SD Data Augmentation
-CustomContrast 217chen2024customcontrast Object PTA SD 1.5 Contrastive Learning
-MoMA song2024moma Object PTA SD 1.5 Multimodal LLM Adapter
-Diptych Prompting 209shin2024large Object; Style PTA FLUX Attention-based Operation; Mask-assisted Generation
+Textual Inversion [7] Object TTF LDM Token Embedding Fine-tuning
+DreamBooth [4] Object TTF Imagen Diffusion Model Fine-tuning; Regularization Dataset
+DreamArtist [92] Object TTF LDM Negative Prompt Fine-tuning
+DVAR [90] Object TTF SD 1.5 Randomness Erasing
+HiPer [166] Object TTF SD Token Embedding Enhancement
+P+ [18] Object TTF SD 1.4 Token Embedding Enhancement
+Unet-finetune [23] Object TTF SD Parameter-efficient Fine-tuning
+Jia et al. [76] Object TTF Imagen Mask-assisted Generation; Regularization
+COTI [83] Object TTF SD 2.0 Data Augmentation
+Gradient-Free TI [94] Object TTF SD Evolutionary Strategy
+PerSAM [44] Object TTF SD Mask-assisted Generation
+DisenBooth [77] Object TTF SD 2.1 Regularization
+NeTI [30] Object TTF SD 1.4 Token Embedding Enhancement
+Prospect [31] Object TTF SD 1.4 Token Embedding Enhancement
+Break-a-Scene [38] Object TTF SD 2.1 Attention-based Operation; Attention-based Operation; Mask-assisted Generation
+COMCAT [96] Object TTF SD 1.4 Attention-based Operation
+ViCo [71] Object TTF SD Attention-based Operation; Regularization
+OFT [167] Object TTF SD 1.5 Finetuning Strategy
+LyCORIS [36] Object TTF SD 1.5 Attention-based Operation
+He et al. [40] Object TTF SD, SDXL Data Augmentation
+DIFFNAT [168] Object TTF SD Regularization
+MATTE [91] Object; Style; High-level Semantic TTF SD Regularization
+LEGO [116] Object TTF SD Regularization
+Catversion [95] Object TTF SD 1.5 Embedding Concatenation
+CLiC [45] Object TTF SD 1.4 Attention-based Operation; Mask-assisted Generation
+HiFi Tuner [32] Object TTF SD 1.4 Attention-based Operation; Mask-assisted Generation; Regularization
+InstructBooth [93] Object TTF SD 1.5 Reinforcement Learning
+DETEX [49] Object TTF SD 1.4 Mask-assisted Generation
+DreamDistribution [169] Object TTF SD 2.1 Token Embedding Enhancement
+DreamTuner [39] Object TTF SD Attention-based Operation
+Xie et al. [34] Object TTF SD Regularization; Mask-assisted Generation
+Infusion [170] Object TTF SD 1.5 Attention-based Operation
+Pair Customization [171] Object TTF SDXL Disentanglement Approach
+DisenDreamer [172] Object TTF SD 2.1 Disentanglement Approach
+DreamBooth+++++ + [173] Object TTF SD Attention-based Operation
+CoRe [174] Object TTF SD 2.1 Regularization
+DCI [175] Object TTF SD Token Embedding Optimization
+Re-Imagen [97] Object PTA Imagen Retrieval-augmented Paradigm
+Versatile Diffusion [176] Object PTA SD 1.4 Architecture Design
+Tuning-Encoder [88] Object PTA SD Regularization
+ELITE [47] Object PTA SD 1.4 Attention-based Operation; Mask-assisted Generation
+UMM-Diffusion [61] Object PTA SD 1.5 Reference Feature Injection
+SuTI [52] Object PTA Imagen Data Augmentation
+InstantBooth [98] Object PTA SD 1.4 Patch Feature Extraction
+BLIP-Diffusion [57] Object PTA SD 1.5 Data Augmentation; Mask-assisted Generation
+Domain-Agnostic [89] Object PTA SD Regularization; Attention-based Operation
+IP-Adapter [51] Object PTA SD 1.5 Reference Feature Injection
+Kosmos-G [60] Object; Face PTA SD 1.5 Multimodal Language Modeling
+Kim et al. [177] Object PTA SD 1.4 Reference Feature Injection
+CAFÉ [58] Object; Face PTA SD 2.0 Multimodal Language Modeling
+SAG [178] Object; Face; Style PTA SD 1.5 Gradient-based Guidance
+BootPIG [179] Object PTA SD Data Augmentation
+CustomContrast [180] Object PTA SD 1.5 Contrastive Learning
+MoMA [59] Object PTA SD 1.5 Multimodal LLM Adapter
+Diptych Prompting [101] Object; Style PTA FLUX Attention-based Operation; Mask-assisted Generation
+StyleDrop [24] Style; Multiple Subjects TTF Muse Data Augmentation
+StyleBoost [84] Style TTF SD 1.5 Regularization Dataset
+StyleAligned [99] Style TTF SDXL Regularization
+GAL [19] Style; Object TTF SD 1.5 Data Augmentation
+StyleForge [181] Style TTF SD 1.5 Parameter-efficient Fine-tuning
+FineStyle [182] Style TTF Muse Mask-assisted Generation
+Style-Friendly [100] Style TTF FLUX-dev; SD 3.5 SNR Sampler
+StyleAdapter [82] Style PTA SD Data Augmentation; Reference Feature Fxtraction
+ArtAdapter [183] Style PTA SD 1.5 Data Augmentation
+ProFusion [106] Face TTF SD 2.0 Fusion Sampling
+Celeb Basis [65] Face TTF SD Face Feature Representation
+Banerjee et al. [184] Face TTF SD 1.4 Regularization
+Magicapture [37] Face; Multiple Subjects TTF SD 1.5 Attention-based Operation; Mask-assisted Generation; Identity Loss
+Concept-centric [163] Face TTF SD 1.5 Modified Classifier-free Guidance
+Cross Initialization [33] Face TTF SD 2.1 Regularization
 Table 3: Paper summary on personalization.
 Paper Scope Framework Technique Backbone
-StyleDrop 5sohn2024styledrop Style; Multiple Subjects TTF Muse Data Augmentation
-StyleBoost 53park2023styleboost Style TTF SD 1.5 Regularization Dataset
-StyleAligned 98hertz2023style Style TTF SDXL Regularization
-GAL 151zhang2024generative Style; Object TTF SD 1.5 Data Augmentation
-StyleForge 153park2024text Style TTF SD 1.5 Parameter-efficient Fine-tuning
-FineStyle 213zhangfinestyle Style TTF Muse Mask-assisted Generation
-Style-Friendly 211choi2024style Style TTF FLUX-dev; SD 3.5 SNR Sampler
-StyleAdapter 26wang2023styleadapter Style PTA SD Data Augmentation; Reference Feature Fxtraction
-ArtAdapter 42chen2023artadapter Style PTA SD 1.5 Data Augmentation
-ProFusion 55zhou2023enhancing Face TTF SD 2.0 Fusion Sampling
-Celeb Basis 60yuan2023inserting Face TTF SD Face Feature Representation
-HyperDreamBooth 58ruiz2023hyperdreambooth Face TTF; PTA SD 1.5 Pretraining and Fast Fine-tuning
-Banerjee et al. 39banerjee2023identity Face TTF SD 1.4 Regularization
-Magicapture 90hyung2023magicapture Face; Multiple Subjects TTF SD 1.5 Attention-based Operation; Mask-assisted Generation; Identity Loss
-Concept-centric 35cao2023concept Face TTF SD 1.5 Modified Classifier-free Guidance
-Cross Initialization 66pang2023cross Face TTF SD 2.1 Regularization
-OMG 115kong2024omg Face; Multiple Subjects TTF SDXL Attention-based Operation; Mask-assisted Generation; Identity Loss
-InstantFamily 161kim2024instantfamily Face; Multiple Subjects; Extra Conditions TTF SD 1.5 Mask-assisted Generation
-PersonaMagic 203li2024personamagic Face TTF SD 1.4 Attention-based Operation
-Su et al. 49su2023identity Face PTA LDM; StyleGAN Identity Loss; Multi-task Learning
-FastComposer 54xiao2023fastcomposer Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
-Face0 88valevski2023face0 Face PTA SD 1.4 Fusion Sampling
-DreamIdentity 89chen2023dreamidentity Face PTA SD 2.1 Attention-based Operation
-Face-Diffuser 52wang2023high Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
-W-Plus-Adapter 91li2023stylegan Face PTA SD 1.5; StyleGAN Face Feature Representation
-Portrait Diffusion 32liu2023portrait Face PTA SD 1.5 Mask-assisted Generation
-RetriNet 92tang2023retrieving Face PTA SD Mask-assisted Generation
-FaceStudio 31yan2023facestudio Face PTA SD Attention-based Operation
-PVA 27xu2024personalized Face PTA LDM Mask-assisted Generation
-DemoCaricature 94chen2023democaricature Face; Extra Conditions PTA SD 1.5 Regularization
-PhotoMaker 28li2023photomaker Face PTA SDXL Data Collection; Reference Feature Extraction
-Stellar 95achlioptas2023stellar Face PTA SDXL Evaluation Prompts and Metrics
-PortraitBooth 96peng2023portraitbooth Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
-InstantID 37wang2024instantid Face; Extra Conditions PTA SDXL Multi-feature Injection
-ID-Aligner 162chen2024id Face PTA SD 1.5; SDXL Feedback Learning
-MoA 164ostashev2024moa Face; Multiple Subjects PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
-IDAdapter 166cui2024idadapter Face PTA SD 2.1 Mixed Facial Features
-Infinite-ID 169wu2024infinite Face PTA SDXL Reference Feature Injection
-Face2Diffusion 170shiohara2024face2diffusion Face PTA SD Multi-feature Injection
-FreeCure 210cai2024foundation Face PTA SD 1.5; SDXL Mask-assisted Generation
-Omni-ID 205qian2024omni Face PTA FLUX Face representation enhancement
+OMG [27] Face; Multiple Subjects TTF SDXL Attention-based Operation; Mask-assisted Generation; Identity Loss
+InstantFamily [185] Face; Multiple Subjects; Extra Conditions TTF SD 1.5 Mask-assisted Generation
+PersonaMagic [186] Face TTF SD 1.4 Attention-based Operation
+HyperDreamBooth [26] Face TTF; PTA SD 1.5 Pretraining and Fast Fine-tuning
+Su et al. [187] Face PTA LDM; StyleGAN Identity Loss; Multi-task Learning
+FastComposer [5] Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
+Face0 [107] Face PTA SD 1.4 Fusion Sampling
+DreamIdentity [48] Face PTA SD 2.1 Attention-based Operation
+Face-Diffuser [79] Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
+W-Plus-Adapter [53] Face PTA SD 1.5; StyleGAN Face Feature Representation
+Portrait Diffusion [188] Face PTA SD 1.5 Mask-assisted Generation
+RetriNet [74] Face PTA SD Mask-assisted Generation
+FaceStudio [109] Face PTA SD Attention-based Operation
+PVA [189] Face PTA LDM Mask-assisted Generation
+DemoCaricature [121] Face; Extra Conditions PTA SD 1.5 Regularization
+PhotoMaker [46] Face PTA SDXL Data Collection; Reference Feature Extraction
+Stellar [35] Face PTA SDXL Evaluation Prompts and Metrics
+PortraitBooth [50] Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
+InstantID [6] Face; Extra Conditions PTA SDXL Multi-feature Injection
+ID-Aligner [190] Face PTA SD 1.5; SDXL Feedback Learning
+MoA [191] Face; Multiple Subjects PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
+IDAdapter [192] Face PTA SD 2.1 Mixed Facial Features
+Infinite-ID [193] Face PTA SDXL Reference Feature Injection
+Face2Diffusion [162] Face PTA SD Multi-feature Injection
+FreeCure [194] Face PTA SD 1.5; SDXL Mask-assisted Generation
+Omni-ID [195] Face PTA FLUX Face representation enhancement
+Custom Diffusion [20] Multiple Subjects TTF SD 1.4 Parameter-efficient Fine-tuning; Constrained Optimization
+Cones [111] Multiple Subjects TTF SD 1.4 Concept Neurons Activation
+SVDiff [22] Multiple Subjects TTF SD Data Augmentation; Attention-based Operation
+Perfusion [21] Multiple Subjects TTF SD 1.5 Regularization
+Mix-of-Show [25] Multiple Subjects TTF SD 1.5 Attention-based Operation; Mask-assisted Generation
+Cones-2 [70] Multiple Subjects TTF SD 2.1 Attention-based Operation; Regularization; Mask-assisted Generation
+PACGen [75] Multiple Subjects TTF SD 1.4 Mask-assisted Generation
+Compositional Inversion [87] Multiple Subjects TTF SD Attention-based Operation; Regularization
+EM-style [72] Multiple Subjects TTF SD Attention-based Operation; Mask-assisted Generation
+MC2 [196] Multiple Subjects TTF SD 1.5 Attention-based Operation
+MultiBooth [64] Multiple Subjects TTF SD 1.5 Mask-assisted Generation
+Matsuda et al. [197] Multiple Subjects TTF LDM Mask-assisted Generation
+MagicTailor [198] Multiple Subjects TTF SD 2.1 Mask-assisted Generation; Attention-based Operation
+AnyDoor [78] Multiple Subjects PTA SD 2.1 Mask-assisted Generation
+Subject-Diffusion [63] Multiple Subjects PTA SD 2.0 Attention-based Operation; Data Augmentation; Mask-assisted Generation
+CustomNet [113] Multiple Subjects; Extra Conditions PTA SD Multi-condition Integration
+MIGC [114] Multiple Subjects PTA SD Attention-based Operation; Mask-assisted Generation
+Emu2 [199] Multiple Subjects PTA SDXL In-Context Learning
+SSR-Encoder [55] Multiple Subjects PTA SD 1.5 Attention-based Operation
+$\lambda$-eclipse [112] Multiple Subjects PTA Kandinsky v2.2 Contrasstive Learning
+MS-Diffusion [56] Multiple Subjects PTA SD Attention-based Operation
+ReVersion [115] High-level Semantic TTF SD Regularization
+Inv-ReVersion [200] High-level Semantic TTF SD 1.5 Regularization
+CusConcept [201] High-level Semantic TTF SD 2.1 Regularization
+ADI [117] High-level Semantic TTF SD 2.1 Mask-assisted Generation
+PhotoSwap [118] Extra Conditions TTF SD 2.1 Attention-based Operation
+PE-VITON [202] Extra Conditions TTF SD Shape and Texture Control
+Layout-Control [69] Extra Conditions TTF SD Attention-based Operation
+SwapAnything [203] Extra Conditions TTF SD 2.1 Mask-assisted Generation
+PE-VITON [202] Extra Conditions TTF SD Shape and Texture Control
+Viewpoint Control [122] Extra Conditions TTF SDXL 3D Feature Incorporation
+Prompt-Free Diffusion [204] Extra Conditions PTA SD 2.0 Reference Feature Injection
+Uni-ControlNet [205] Extra Conditions PTA SD Multi-feature Injection
+ViscoNet [206] Extra Conditions PTA SD Multi-feature Injection
+Context Diffusion [207] Extra Conditions PTA LDM Multi-feature Injection
+FreeControl [208] Extra Conditions PTA SD 1.5; SD 2.1; SDXL Fusion Guidance
+Li et al. [209] Extra Conditions PTA SD Attention-based Operation; Mask-assisted Generation
 Table 4: Paper summary on personalization.
 Paper Scope Framework Technique Backbone
-Custom Diffusion 6kumari2023multi Multiple Subjects TTF SD 1.4 Parameter-efficient Fine-tuning; Constrained Optimization
-Cones 4liu2023cones Multiple Subjects TTF SD 1.4 Concept Neurons Activation
-SVDiff 8han2023svdiff Multiple Subjects TTF SD Data Augmentation; Attention-based Operation
-Perfusion 7tewel2023key Multiple Subjects TTF SD 1.5 Regularization
-Mix-of-Show 46gu2024mix Multiple Subjects TTF SD 1.5 Attention-based Operation; Mask-assisted Generation
-Cones-2 62liu2023cones Multiple Subjects TTF SD 2.1 Attention-based Operation; Regularization; Mask-assisted Generation
-PACGen 34li2023generate Multiple Subjects TTF SD 1.4 Mask-assisted Generation
-Compositional Inversion 9zhang2023compositional Multiple Subjects TTF SD Attention-based Operation; Regularization
-EM-style 44rahman2024visual Multiple Subjects TTF SD Attention-based Operation; Mask-assisted Generation
-MC2 154jiang2024mc Multiple Subjects TTF SD 1.5 Attention-based Operation
-MultiBooth 156zhu2024multibooth Multiple Subjects TTF SD 1.5 Mask-assisted Generation
-Matsuda et al. 165matsuda2024multi Multiple Subjects TTF LDM Mask-assisted Generation
-MagicTailor zhou2024magictailor Multiple Subjects TTF SD 2.1 Mask-assisted Generation; Attention-based Operation
-AnyDoor 14chen2023anydoor Multiple Subjects PTA SD 2.1 Mask-assisted Generation
-Subject-Diffusion 16ma2023subject Multiple Subjects PTA SD 2.0 Attention-based Operation; Data Augmentation; Mask-assisted Generation
-CustomNet 29yuan2023customnet Multiple Subjects; Extra Conditions PTA SD Multi-condition Integration
-MIGC 174zhou2024migc Multiple Subjects PTA SD Attention-based Operation; Mask-assisted Generation
-Emu2 223sun2024generative Multiple Subjects PTA SDXL In-Context Learning
-SSR-Encoder 221zhang2024ssr Multiple Subjects PTA SD 1.5 Attention-based Operation
-$\lambda$-eclipse 222patel2024lambda Multiple Subjects PTA Kandinsky v2.2 Contrasstive Learning
-MS-Diffusion 220personalizationms Multiple Subjects PTA SD Attention-based Operation
-ReVersion 63huang2023reversion High-level Semantic TTF SD Regularization
-Inv-ReVersion 159zhang2024inv High-level Semantic TTF SD 1.5 Regularization
-CusConcept 215xu2024cusconcept High-level Semantic TTF SD 2.1 Regularization
-ADI 43huang2023learning High-level Semantic TTF SD 2.1 Mask-assisted Generation
-PhotoSwap 15gu2024photoswap Extra Conditions TTF SD 2.1 Attention-based Operation
-PE-VITON 38zhang2023two Extra Conditions TTF SD Shape and Texture Control
-Layout-Control 152chen2024training Extra Conditions TTF SD Attention-based Operation
-SwapAnything 155gu2024swapanything Extra Conditions TTF SD 2.1 Mask-assisted Generation
-PE-VITON 38zhang2023two Extra Conditions TTF SD Shape and Texture Control
-Viewpoint Control 163kumari2024customizing Extra Conditions TTF SDXL 3D Feature Incorporation
-Prompt-Free Diffusion 110xu2023prompt Extra Conditions PTA SD 2.0 Reference Feature Injection
-Uni-ControlNet 111zhao2024uni Extra Conditions PTA SD Multi-feature Injection
-ViscoNet 93cheong2023visconet Extra Conditions PTA SD Multi-feature Injection
-Context Diffusion 112najdenkoska2023context Extra Conditions PTA LDM Multi-feature Injection
-FreeControl 113mo2023freecontrol Extra Conditions PTA SD 1.5; SD 2.1; SDXL Fusion Guidance
-Li et al. 168li2024tuning Extra Conditions PTA SD Attention-based Operation; Mask-assisted Generation
-Table 5: Paper summary on personalization.
-Paper Scope Framework Technique Backbone
-Tune-A-Video 150wu2023tune Video TTF SD 1.4 Parameter-efficient Fine-tuning
-Gen-1 148esser2023structure Video TTF LDM Multi-feature Injection
-Make-A-Protagonist 135zhao2023make Video TTF LDM Attention-based Operation
-Animate-A-Story 136he2023animate Video TTF SD Retrieval-augmented Paradigm
-MotionDirector 101zhao2023motiondirector Video TTF ZeroScope; ModelScope Disentanglement Approach
-LAMP 102wu2023lamp Video TTF SDXL; SD 1.4 First-frame Conditioned Pipeline
-VMC 140jeong2023vmc Video TTF Show-1 Parameter-efficient Fine-tuning
-SAVE 103song2023save Video TTF VDM Attention-based Operation
-Customizing Motion 104materzynska2023customizing Video TTF ZeroScope Regularization
-DreamVideo 142wei2023dreamvideo Video TTF ModelScope Disentanglement Approach
-MotionCrafter 106zhang2023motioncrafter Video TTF VDM Disentanglement Approach
-Customize-A-Video 172ren2024customize Video TTF ModelScope Multi-stage Refinement
-Magic-Me 173ma2024magic Video TTF SD 1.5; AnimateDiff Multi-stage Refinement
-CustomTTT 202bi2024customttt Video TTF VDM Fine-tuning and Alignment
-PersonalVideo 208li2024personalvideo Video; Face TTF AnimateDiff Identity Injection
-CustomVideo 145wang2024customvideo Video; Multiple Subjects TTF ZeroScope Data Augmentation; Attention-based Operation
-VideoDreamer 138chen2023videodreamer Video; Multiple Subjects TTF SD 2.1 Data Augmentation
-VideoAssembler 82zhao2023videoassembler Video PTA VidRD Reference Feature Injection
-VideoBooth 83jiang2023videobooth Video PTA VDM Reference Feature Injection
-VideoMaker 201wu2024videomaker Video PTA VDM Reference Feature Injection
-SUGAR 204zhou2024sugar Video PTA VDM Attention-based Operation
-StyleCrafter 212liu2024stylecrafter Video; Style PTA SD 1.5 Reference Feature Injection
-ID-Animator 158he2024id Video; Face PTA AnimateDiff Reference Feature Injection
-ConsisID 207yuan2024identity Video; Face PTA CogVideoX frequency decomposition
-MotionCharacter 206fang2024motioncharacter Video; Face PTA VDM ID-Preserving Optimization; Motion Control Enhancement
-DreaMoving 105feng2023dreamoving Video; Face; Extra Conditions PTA SD Multi-feature Injection
-Magic3D 120lin2023magic3d 3D TTF Imagen; LDM Score Distillation Sampling
-DreamBooth3D 121raj2023dreambooth3d 3D TTF Imagen Multi-stage Refinement
-PAS 122azadi2023text 3D PTA SD Text-to-3D-Pose
-StyleAvatar3D 123zhang2023styleavatar3d 3D PTA LDM Multi-view Alignment
-AvatarBooth 124zeng2023avatarbooth 3D TTF SD Dual Model Fine-tuning
-MVDream 126shi2023mvdream 3D TTF SD 2.1 Score Distillation Sampling
-Consist3D 127ouyang2023chasing 3D TTF SD Token Embedding Enhancement
-Animate124 128zhao2023animate124 3D TTF SD 1.5 Multi-stage Refinement
-Dream-in-4D 130zheng2023unified 3D TTF SD 2.1 Multi-stage Refinement
-TextureDreamer 133yeh2024texturedreamer 3D TTF LDM Texture Extraction
-TIP-Editor 134zhuang2024tip 3D TTF SD Multi-stage Refinement; Mask-assisted Generation
-Anti-DreamBooth 13van2023anti Attack and Defense TTF SD 2.1 Perturbation Learning
-Concept Censorship 50zhang2023backdooring Attack and Defense TTF SD 1.4 Trigger Injection
-Huang et al. 167huang2024personalization Attack and Defense TTF SD backdoor attack
-Continual Diffusion 64smith2023continual Others TTF SD Continual Learning
-SVGCustomization 36zhang2023text Others TTF SD 1.5 Fine-tuning and Alignment
-StitchDiffusion 25wang2024customizing Others TTF LDM Parameter-efficient Fine-tuning
-MC-TI 214wang2024multi Others TTF SD 1.5 Regularization
+Tune-A-Video [127] Video TTF SD 1.4 Parameter-efficient Fine-tuning
+Gen-1 [130] Video TTF LDM Multi-feature Injection
+Make-A-Protagonist [210] Video TTF LDM Attention-based Operation
+Animate-A-Story [131] Video TTF SD Retrieval-augmented Paradigm
+MotionDirector [134] Video TTF ZeroScope; ModelScope Disentanglement Approach
+LAMP [132] Video TTF SDXL; SD 1.4 First-frame Conditioned Pipeline
+VMC [129] Video TTF Show-1 Parameter-efficient Fine-tuning
+SAVE [128] Video TTF VDM Attention-based Operation
+Customizing Motion [133] Video TTF ZeroScope Regularization
+DreamVideo [135] Video TTF ModelScope Disentanglement Approach
+MotionCrafter [211] Video TTF VDM Disentanglement Approach
+Customize-A-Video [212] Video TTF ModelScope Multi-stage Refinement
+Magic-Me [213] Video TTF SD 1.5; AnimateDiff Multi-stage Refinement
+CustomTTT [214] Video TTF VDM Fine-tuning and Alignment
+PersonalVideo [215] Video; Face TTF AnimateDiff Identity Injection
+CustomVideo [81] Video; Multiple Subjects TTF ZeroScope Data Augmentation; Attention-based Operation
+VideoDreamer [80] Video; Multiple Subjects TTF SD 2.1 Data Augmentation
+VideoAssembler [125] Video PTA VidRD Reference Feature Injection
+VideoBooth [124] Video PTA VDM Reference Feature Injection
+VideoMaker [216] Video PTA VDM Reference Feature Injection
+SUGAR [217] Video PTA VDM Attention-based Operation
+StyleCrafter [54] Video; Style PTA SD 1.5 Reference Feature Injection
+ID-Animator [218] Video; Face PTA AnimateDiff Reference Feature Injection
+ConsisID [219] Video; Face PTA CogVideoX frequency decomposition
+MotionCharacter [220] Video; Face PTA VDM ID-Preserving Optimization; Motion Control Enhancement
+DreaMoving [126] Video; Face; Extra Conditions PTA SD Multi-feature Injection
+Magic3D [138] 3D TTF Imagen; LDM Score Distillation Sampling
+DreamBooth3D [140] 3D TTF Imagen Multi-stage Refinement
+PAS [145] 3D PTA SD Text-to-3D-Pose
+StyleAvatar3D [146] 3D PTA LDM Multi-view Alignment
+AvatarBooth [147] 3D TTF SD Dual Model Fine-tuning
+MVDream [139] 3D TTF SD 2.1 Score Distillation Sampling
+Consist3D [141] 3D TTF SD Token Embedding Enhancement
+Animate124 [143] 3D TTF SD 1.5 Multi-stage Refinement
+Dream-in-4D [144] 3D TTF SD 2.1 Multi-stage Refinement
+TextureDreamer [142] 3D TTF LDM Texture Extraction
+TIP-Editor [221] 3D TTF SD Multi-stage Refinement; Mask-assisted Generation
+Anti-DreamBooth [148] Attack and Defense TTF SD 2.1 Perturbation Learning
+Concept Censorship [149] Attack and Defense TTF SD 1.4 Trigger Injection
+Huang et al. [222] Attack and Defense TTF SD backdoor attack
+Continual Diffusion [223] Others TTF SD Continual Learning
+SVGCustomization [150] Others TTF SD 1.5 Fine-tuning and Alignment
+StitchDiffusion [151] Others TTF LDM Parameter-efficient Fine-tuning
+MC-TI [224] Others TTF SD 1.5 Regularization
 Acknowledgments

### v5

@@
 Recently, generative models have shown remarkable progress in the field of natural language processing and computer vision.
-Notable examples, such as ChatGPT [1] and text-to-image diffusion models [2], have showcased impressive capabilities in content creation.
+Notable examples, such as ChatGPT [wu2023brief] and text-to-image diffusion models [croitoru2023diffusion], have showcased impressive capabilities in content creation.
 However, these advanced models often struggle to fulfill specific requirements, such as answering domain-specific queries or accurately depicting user’s portraits.
 This limitation highlights the importance of Personalized Content Synthesis (PCS), which enables users to customize models for their unique tasks and requirements.
-As a critical area of research, PCS is increasingly recognized as essential for achieving Artificial General Intelligence (AGI), evidenced by the growing number of companies releasing products that support personalized content creation, like utilizing Reinforcement Learning to fine-tune language models [3].
+As a critical area of research, PCS is increasingly recognized as essential for achieving Artificial General Intelligence (AGI), evidenced by the growing number of companies releasing products that support personalized content creation, like utilizing Reinforcement Learning to fine-tune language models [uc2023survey].
 In this paper, we focus on PCS within the context of diffusion models in computer vision.
@@
 1, PCS can customize a unique cat into various scenarios, such as “wearing pink sunglasses” and “in the snow”.
-Figure 1: Given a few reference images of a subject (e.g., a cat [4] or face [5]), PCS aims to generate new renditions of the subject that align with user-defined textual prompts.
+Figure 1: Given a few reference images of a subject (e.g., a cat [2ruiz2023dreambooth] or face [54xiao2023fastcomposer]), PCS aims to generate new renditions of the subject that align with user-defined textual prompts.
 The task requires preserving the subject’s identity while adapting to diverse contexts.
-The examples are generated by this survey using DreamBooth [4] and InstantID [6].
+The examples are generated by this survey using DreamBooth [2ruiz2023dreambooth] and InstantID [37wang2024instantid].
 The emergence of diffusion models has significantly facilitated text-guided content generation, leading to a rapid expansion of PCS.
@@
 2, the number of research papers on PCS has surged over time.
-Following the release of key innovations such as DreamBooth [4] and Textual Inversion [7] in August 2022, over 150 methods have been proposed in a remarkably short period.
+Following the release of key innovations such as DreamBooth [2ruiz2023dreambooth] and Textual Inversion [1gal2022image] in August 2022, over 150 methods have been proposed in a remarkably short period.
 These methods can be classified using several criteria.
@@
 We divide PCS methods with 3 different criteria: training strategy, personalization scope, and technique.
-Figure 3: The trade-off between text alignment and visual fidelity in personalized image synthesis, illustrated through DreamBooth-generated [4] examples of a customized cat wearing sunglasses.
+Figure 3: The trade-off between text alignment and visual fidelity in personalized image synthesis, illustrated through DreamBooth-generated [2ruiz2023dreambooth] examples of a customized cat wearing sunglasses.
 Overfitting occurs when the model focuses solely on reconstructing the cat, disregarding the sunglasses context.
@@
 2 Fundamentals
-Modern diffusion modeling has evolved into a sophisticated framework that unifies discrete and continuous generative paradigms through stochastic differential equations (SDEs) [8], ordinary differential equations (ODEs) [8, 9].
+Modern diffusion modeling has evolved into a sophisticated framework that unifies discrete and continuous generative paradigms through stochastic differential equations (SDEs) [song2020score], ordinary differential equations (ODEs) [song2020score, karras2022elucidating].
 This section introduces the core mathematical foundations, emphasizing innovations in the diffusion process and conditional mechanisms.
@@
 The architecture of diffusion models comprises two complementary processes: a forward diffusion that systematically perturbs data distributions, and a learned reverse process that reconstructs signals through iterative refinement.
-Originating from discrete-time Markov chains in Denoising Diffusion Probabilistic Model (DDPM) [10], the forward process applies Gaussian noise corruption over $T$ steps:
+Originating from discrete-time Markov chains in Denoising Diffusion Probabilistic Model (DDPM) [ho2020denoising], the forward process applies Gaussian noise corruption over $T$ steps:
 $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},\quad\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ (1)
@@
 Although DDPMs establish basic denoising dynamics, this discrete formulation imposes constraints on flexible noise scheduling and computational efficiency.
-To overcome these limitations, the framework was generalized through continuous-time SDE [8], which provide a unified perspective for modeling diffusion dynamics:
+To overcome these limitations, the framework was generalized through continuous-time SDE [song2020score], which provide a unified perspective for modeling diffusion dynamics:
 $d\mathbf{x}=\mathbf{f}(\mathbf{x},t)dt+g(t)d\mathbf{w}$ (4)
 where $\mathbf{f}(\mathbf{x},t)$ encodes deterministic drift components, $g(t)$ modulates stochastic diffusion via Wiener process $\mathbf{w}$.
-This continuous perspective subsumes DDPMs as special cases while enabling adaptive noise scheduling strategies such as variance-preserving (VP) and variance-exploding (VE) schedules through strategic choices of $\mathbf{f}$ and $g$ [9].
-For the reverse process, a crucial result [11] shows that the reverse diffusion process also follows an SDE:
+This continuous perspective subsumes DDPMs as special cases while enabling adaptive noise scheduling strategies such as variance-preserving (VP) and variance-exploding (VE) schedules through strategic choices of $\mathbf{f}$ and $g$ [karras2022elucidating].
+For the reverse process, a crucial result [anderson1982reverse] shows that the reverse diffusion process also follows an SDE:
 $d\mathbf{x}=\left[\mathbf{f}(\mathbf{x},t)-g(t)^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right]dt+g(t)d\mathbf{\bar{w}}$ (5)
@@
 Furthermore, stochastic trajectories exhibit erratic curvature profiles, forcing fixed-step solvers to adopt conservative discretization schemes that often necessitate 1,000+ steps for high-fidelity generation.
-These limitations motivate the derivation of deterministic sampling trajectories through probability flow ordinary differential equations (ODEs) [8, 9], obtained by eliminating the stochastic term from the SDE formulation:
+These limitations motivate the derivation of deterministic sampling trajectories through probability flow ordinary differential equations (ODEs) [song2020score, karras2022elucidating], obtained by eliminating the stochastic term from the SDE formulation:
 $d\mathbf{x}=\left[\mathbf{f}(\mathbf{x},t)-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})\right]dt$ (6)
@@
 Crucially, the resulting trajectories exhibit intrinsic geometric regularity, allowing the complex denoising process to operate in a simplified mathematical space.
-This enabling 5-10x faster sampling than SDEs through adaptive ODE solvers like DPM-Solver [12] and DPM-Solver++ [13].
-Recent work [14] further identifies strong geometric regularity in these ODE-based sampling trajectories, demonstrating that they inherently follow a linear-nonlinear-linear structure regardless of generated content.
+This enabling 5-10x faster sampling than SDEs through adaptive ODE solvers like DPM-Solver [lu2022dpm] and DPM-Solver++ [lu2022dpmplus].
+Recent work [chen2024trajectory] further identifies strong geometric regularity in these ODE-based sampling trajectories, demonstrating that they inherently follow a linear-nonlinear-linear structure regardless of generated content.
 This trajectory regularity enables dynamic programming-based time scheduling optimization with negligible computational overhead.
@@
 $\mathcal{L}=\mathbb{E}_{\mathbf{x}_{0},t,\mathbf{\epsilon}}\left[\|\mathbf{\epsilon}-\mathbf{\epsilon}_{\theta}(\mathbf{x}_{t},t,c)\|_{2}^{2}\right]$ (7)
-where the conditioning signal $c$ can be integrated via multiple synergistic mechanisms, such as cross-modal attention [15] and spatial modulation [16].
+where the conditioning signal $c$ can be integrated via multiple synergistic mechanisms, such as cross-modal attention [rombach2022high] and spatial modulation [zhang2023adding].
 This conditional paradigm directly enables the applications of PCS.
-Leading text-to-image systems, such as Stable Diffusion (SD) [15] and DALLE [17], are widely adopted to empower users in controlling customized content through text instructions.
+Leading text-to-image systems, such as Stable Diffusion (SD) [rombach2022high] and DALLE [ramesh2022hierarchical], are widely adopted to empower users in controlling customized content through text instructions.
 3 Generic Framework
@@
 Compared to the large-scale pre-training described in Equation 7, key differences lie in the training data and learnable parameters.
-The training samples are typically restricted to the SoI references, sometimes supplemented with a regularization dataset to mitigate overfitting [4].
-For the selection of learnable parameters $\mathbf{\theta^{\prime}}$, the commonly adopted options include token embeddings [7, 18], the entire diffusion model [4, 19], specific subsets of parameters [20, 21, 22], or introducing new parameters such as adapters [23, 24], and LoRA [25, 26, 27], which will be discussed in Section 3.1.3.
+The training samples are typically restricted to the SoI references, sometimes supplemented with a regularization dataset to mitigate overfitting [2ruiz2023dreambooth].
+For the selection of learnable parameters $\mathbf{\theta^{\prime}}$, the commonly adopted options include token embeddings [1gal2022image, 23voynov2023p+], the entire diffusion model [2ruiz2023dreambooth, 151zhang2024generative], specific subsets of parameters [6kumari2023multi, 7tewel2023key, 8han2023svdiff], or introducing new parameters such as adapters [72xiang2023closer, 5sohn2024styledrop], and LoRA [46gu2024mix, 58ruiz2023hyperdreambooth, 115kong2024omg], which will be discussed in Section 3.1.3.
 Test-time generation.
@@
 Normally, the construction of the unique modifier can be divided into three categories:
-Plain text [4, 20, 21, 24, 19].
+Plain text [2ruiz2023dreambooth, 6kumari2023multi, 7tewel2023key, 5sohn2024styledrop, 151zhang2024generative].
 This approach utilizes an explicit text description to represent the SoI.
 For example, words such as cat could directly represent the user’s cat in the references.
-This setting typically requires fine-tuning the parameters of the diffusion model components, such as UNet [28] or Transformer blocks [29], to enable the model to associate the SoI’s visual features with the plain‐text token.
+This setting typically requires fine-tuning the parameters of the diffusion model components, such as UNet [ronneberger2015unet] or Transformer blocks [vaswani2017attention], to enable the model to associate the SoI’s visual features with the plain‐text token.
 The plain text offers user-friendly prompt construction and injects subject prior information to ease tuning difficulty.
 However, this technique may over-specialize common terms, limiting their broader applicability as the model learns to associate general words with specific SoI characteristics.
-Rare token [4].
+Rare token [2ruiz2023dreambooth].
 This technique employs infrequently used tokens to minimize their impact on commonly used vocabulary.
@@
 However, these rare tokens often fail to provide useful subject prior information and still exhibit weak interference with unrelated vocabulary, potentially leading to ambiguity between the original meaning and the intended SoI reference.
-Learnable token embedding [7, 18, 30, 31, 32, 33].
+Learnable token embedding [1gal2022image, 23voynov2023p+, 61alaluf2023neural, 30zhang2023prospect, 48wang2023hifi, 66pang2023cross].
 This method adds a new token and its corresponding embedding vector to the dictionary of the tokenizer.
@@
 Token embedding.
-As introduced in Section 3.1.2, token embedding optimization [7] introduces learnable tokens as the unique modifier to represent SoI through noise-to-image reconstruction.
+As introduced in Section 3.1.2, token embedding optimization [1gal2022image] introduces learnable tokens as the unique modifier to represent SoI through noise-to-image reconstruction.
 While achieving remarkable parameter efficiency, this approach faces challenges in detail preservation and prolonged training time (typically larger than 20 minutes) due to the inherent compression of complex features into low-dimensional embeddings.
-Subsequent works [18, 30, 31, 32, 33] aim to address these limitations through different strategies, which are summarized in Section 5.1.
+Subsequent works [23voynov2023p+, 61alaluf2023neural, 30zhang2023prospect, 48wang2023hifi, 66pang2023cross] aim to address these limitations through different strategies, which are summarized in Section 5.1.
 Existing model parameters.
-This paradigm directly optimizes pre-trained model components, such as the text encoder, U-Net blocks, and transformer layers [4, 19, 20, 21, 22].
+This paradigm directly optimizes pre-trained model components, such as the text encoder, U-Net blocks, and transformer layers [2ruiz2023dreambooth, 151zhang2024generative, 6kumari2023multi, 7tewel2023key, 8han2023svdiff].
 Benefitting from the advanced representation capacity of these modules, the fine-tuning phase can achieve faster convergence (5-10 minutes) and superior visual fidelity compared to token-only methods, though at the cost of significant storage overhead.
@@
 Parameter-efficient extensions.
-Recent advanced methods have introduced parameter-efficient techniques into PCS, such as LoRA [25, 26, 27, 34, 35, 36] and adapter modules [23, 24], which inject small, trainable components into the base model.
+Recent advanced methods have introduced parameter-efficient techniques into PCS, such as LoRA [46gu2024mix, 58ruiz2023hyperdreambooth, 115kong2024omg, 51lu2024object, 95achlioptas2023stellar, 76yeh2023navigating] and adapter modules [72xiang2023closer, 5sohn2024styledrop], which inject small, trainable components into the base model.
 These methods achieve comparable performance to full parameter fine-tuning while dramatically reducing storage requirements.
@@
 Since the aforementioned strategies are not conflicting, some methods allocate different learning rates and training phases to each component type to achieve optimal balance between fidelity and efficiency.
-For instance, the fine-tuned token embedding can be regarded as an effective initialization for the subsequent model weight fine-tuning [37].
-Also, these two parts can be simultaneously optimized with different learning rates [38, 39].
+For instance, the fine-tuned token embedding can be regarded as an effective initialization for the subsequent model weight fine-tuning [90hyung2023magicapture].
+Also, these two parts can be simultaneously optimized with different learning rates [74avrahami2023break, 86hua2023dreamtuner].
 3.1.4 Prompt Engineering
@@
 A simplest example is “Photo of V*”.
-However, DreamBooth [4] noted that such a simple description causes a long training time and unsatisfactory performance.
+However, DreamBooth [2ruiz2023dreambooth] noted that such a simple description causes a long training time and unsatisfactory performance.
 To address this, they incorporate the unique modifier with a class noun to describe the SoI in the references (e.g., “Photo of V* cat”).
-Also, the training caption for each training reference can be more precious for better disentanglement of SoI and irrelevant concepts [40], such as “Photo of V* cat on the chair”.
-This follows the tending that high-quality captions in the training set could assist in further improvement of accurate text control [41].
+Also, the training caption for each training reference can be more precious for better disentanglement of SoI and irrelevant concepts [57he2023data], such as “Photo of V* cat on the chair”.
+This follows the tending that high-quality captions in the training set could assist in further improvement of accurate text control [BetkerImprovingIG].
 The PTA framework has emerged as a breakthrough approach for PCS, aiming to eliminate the computational burden of per-request fine-tuning while maintaining high-quality, subject-specific generation capabilities.
@@
 Extracting representative features of the SoI is crucial in the creation of personalized content.
-A common approach is to employ an encoder, leveraging pre-trained models such as CLIP [42] and BLIP [43].
+A common approach is to employ an encoder, leveraging pre-trained models such as CLIP [radford2021learning] and BLIP [li2022blip].
 While these models excel at capturing global features, they often include irrelevant information that can detract from the fidelity, potentially compromising the quality of the personalized output, such as including the same background in the generation.
 To mitigate this issue, some studies incorporate additional prior knowledge to guide the learning process so as to focus on the targeted SoI.
-For instance, the SoI-specific mask [44, 32, 38, 45, 34, 35] contributes to the effective exclusion of the influence of the background.
-Moreover, using facial landmarks [6] in the context of human face customization helps improve identity preservation.
+For instance, the SoI-specific mask [21zhang2023personalize, 48wang2023hifi, 74avrahami2023break, 81safaee2023clic, 51lu2024object, 95achlioptas2023stellar] contributes to the effective exclusion of the influence of the background.
+Moreover, using facial landmarks [37wang2024instantid] in the context of human face customization helps improve identity preservation.
 We will discuss a more detailed technical summary of mask-assistant generation in Section 4.2.
@@
 Yet, the majority of current PTA systems are limited to supporting one reference input.
-Some research works [6, 46] propose to average or stack features extracted from multiple references to form a composite SoI representation.
+Some research works [37wang2024instantid, 28li2023photomaker] propose to average or stack features extracted from multiple references to form a composite SoI representation.
 3.2.2 Subject Feature Fusion
@@
 Current methodologies can be categorized into four primary approaches:
-Concatenation-based Fusion [47, 46, 5, 48, 49, 50].
+Concatenation-based Fusion [19wei2023elite, 28li2023photomaker, 54xiao2023fastcomposer, 89chen2023dreamidentity, 85cai2023decoupled, 96peng2023portraitbooth].
 This method makes the unique modifier a placeholder token to encapsulate visual subject characteristics.
@@
 This combined representation subsequently guides generation through standard cross-attention layers in the diffusion process, enabling basic subject-text alignment while maintaining architectural simplicity.
-Cross-attention Fusion [51, 52, 6, 53, 54, 55, 56].
+Cross-attention Fusion [40ye2023ip, 11chen2024subject, 37wang2024instantid, 91li2023stylegan, 212liu2024stylecrafter, 221zhang2024ssr, 220personalizationms].
 This paradigm extends the U-Net architecture with specialized attention mechanisms that jointly process visual and textual conditions.
-For example, IP-Adapter [51] introduces decoupled cross-attention layers that maintain separate query projections for image and text features.
+For example, IP-Adapter [40ye2023ip] introduces decoupled cross-attention layers that maintain separate query projections for image and text features.
 In this case, the unique modifier directly displays the plain text.
-Multimodal Encoder Fusion [57, 58, 59, 60, 61].
-This approach leverages powerful multimodal encoder architectures (e.g., BLIP-2 [62]) to jointly embed visual and textual subject descriptors.
-BLIP-Diffusion [57] exemplifies this strategy by learning a compact subject prompt embedding that fuses image patches with textual names through Q-Former modules.
-Hybrid Fusion [63, 64].
+Multimodal Encoder Fusion [10li2024blip, 84zhou2023customization, song2024moma, 77pan2023kosmos, 71ma2023unified].
+This approach leverages powerful multimodal encoder architectures (e.g., BLIP-2 [li2023blip]) to jointly embed visual and textual subject descriptors.
+BLIP-Diffusion [10li2024blip] exemplifies this strategy by learning a compact subject prompt embedding that fuses image patches with textual names through Q-Former modules.
+Hybrid Fusion [16ma2023subject, 156zhu2024multibooth].
 Moreover, some systems integrate multiple fusion strategies.
-For example, Subject-Diffusion [63] integrate both concatenation and cross-attention fusion, taking advantage of the strengths of each approach to enhance the overall personalization capability.
+For example, Subject-Diffusion [16ma2023subject] integrate both concatenation and cross-attention fusion, taking advantage of the strengths of each approach to enhance the overall personalization capability.
 3.2.3 Training Data
@@
 Several strategies have been proposed to mitigate this issue: 1) Data Augmentation.
-Techniques such as foreground segmentation followed by placement in a different background are used to construct triplet data [57].
+Techniques such as foreground segmentation followed by placement in a different background are used to construct triplet data [10li2024blip].
 2) Synthetic Sample Generation.
-Methods like SuTI [52] utilize multiple TTF models to generate synthetic samples, which are then paired with original references.
+Methods like SuTI [11chen2024subject] utilize multiple TTF models to generate synthetic samples, which are then paired with original references.
 3) Utilizing Recognizable SoIs.
-Collecting images of easily recognizable subjects, such as celebrities, significantly facilitates face personalization [65].
+Collecting images of easily recognizable subjects, such as celebrities, significantly facilitates face personalization [60yuan2023inserting].
 Dual Data (Reference Image, Reference Caption).
 This dataset is essentially a simplified version of the triplet format, where the personalized content is the original image itself.
-Such datasets are more accessible, including collections like LAION [66] and LAION-FACE [67].
+Such datasets are more accessible, including collections like LAION [schuhmann2022laion] and LAION-FACE [zheng2022general].
 However, a notable drawback is that training tends to focus more on reconstructing the reference image rather than incorporating the text prompts.
@@
 Recently, some works have started to explore the combination of TTF and PTA methods.
-HyperDreamBooth [26] states that PTA methods provide a general framework capable of handling a wide range of common objects, while TTF techniques enable fine-tuning to specific instances, which can improve the preservation of fine-grained details.
+HyperDreamBooth [58ruiz2023hyperdreambooth] states that PTA methods provide a general framework capable of handling a wide range of common objects, while TTF techniques enable fine-tuning to specific instances, which can improve the preservation of fine-grained details.
 They first develop a PTA network, followed by a subject-driven fine-tuning.
-Similarly, DreamTuner [39] pre-trains a subject encoder that outputs diffusion conditions for accurate reconstruction.
+Similarly, DreamTuner [86hua2023dreamtuner] pre-trains a subject encoder that outputs diffusion conditions for accurate reconstruction.
 Then an additional fine-tuning stage is conducted for fine identity preservation.
-In contrast, SuTI [52] first applies TTF methods to generate synthetic pairwise samples, which can be used for training the PTA network.
+In contrast, SuTI [11chen2024subject] first applies TTF methods to generate synthetic pairwise samples, which can be used for training the PTA network.
 4 Techniques in Personalized Content Synthesis
@@
 4.1 Attention-based Operation
-Attention-based operations have become a crucial technique in model learning, particularly for processing features effectively [68].
+Attention-based operations have become a crucial technique in model learning, particularly for processing features effectively [hassanin2024visual].
 In diffusion models, these operations generally involve manipulating the way a model focuses on different parts of data, often through a method known as the Query-Key-Value (QKV) scheme.
@@
 A cluster of studies focuses on restricting the influence of the SoI token within the attention layers.
-For example, Mix-of-Show [25] designs region-aware cross-attention where the feature map is initially generated by the global prompt and replaced with distinct regional features corresponding to each entity.
+For example, Mix-of-Show [46gu2024mix] designs region-aware cross-attention where the feature map is initially generated by the global prompt and replaced with distinct regional features corresponding to each entity.
 This avoids the misalignment between the words and visual regions.
-DreamTuner [39] designs a self-subject-attention layer to further refine the subject identity.
+DreamTuner [86hua2023dreamtuner] designs a self-subject-attention layer to further refine the subject identity.
 This attention module takes the features of the generated image as query, the concatenation of generated features as key, and reference features as value.
-Layout-Control [69] adjusts attention weights specifically around the layout without additional training.
-Cones 2 [70] also defines some negative attention areas to penalize the illegal occupation to allow multiple object generation.
-VICO [71] inserts a new attention layer where a binary mask is deployed to selectively obscure the attention map between the noisy latent and the reference image features.
+Layout-Control [152chen2024training] adjusts attention weights specifically around the layout without additional training.
+Cones 2 [62liu2023cones] also defines some negative attention areas to penalize the illegal occupation to allow multiple object generation.
+VICO [20hao2023vico] inserts a new attention layer where a binary mask is deployed to selectively obscure the attention map between the noisy latent and the reference image features.
 Implicit attention guidance.
-In addition to these explicit attention weights modification methods, many researchers [5, 22, 38, 63, 45, 72, 37] employ localization supervision in the cross-attention module.
+In addition to these explicit attention weights modification methods, many researchers [54xiao2023fastcomposer, 8han2023svdiff, 74avrahami2023break, 16ma2023subject, 81safaee2023clic, 44rahman2024visual, 90hyung2023magicapture] employ localization supervision in the cross-attention module.
 Specifically, they train cross-attention modules using coordinate-aware loss functions, forcing attention maps to align with annotated subject positions.
-DreamTuner [39] further refines this approach by designing an attention layer that effectively integrates features from different parts of the image.
+DreamTuner [86hua2023dreamtuner] further refines this approach by designing an attention layer that effectively integrates features from different parts of the image.
 4.2 Mask-guided Generation
@@
 Pixel-level mask.
-Benefitting from advanced segmentation methods like SAM [73], the SoI can be precisely isolated from the background.
-Based on this strategy, plenty of studies [44, 32, 38, 45, 34, 35, 37, 74, 75] choose to discard the pixels of the background area so that the reconstruction loss can focus on the targeted object and exclude irrelevant disturbances.
-Another technique [49] further adds the masked background reconstruction for better disentanglement.
-In addition, the layout indicated by the pixel mask can be incorporated into the attention modules as a supervision signal [5, 22, 38, 63, 45, 72, 37] to adjust the attention’s concentration adaptively.
-Moreover, the mask can stitch specific feature maps to construct more informative semantic patterns [47, 76, 63].
+Benefitting from advanced segmentation methods like SAM [kirillov2023segment], the SoI can be precisely isolated from the background.
+Based on this strategy, plenty of studies [21zhang2023personalize, 48wang2023hifi, 74avrahami2023break, 81safaee2023clic, 51lu2024object, 95achlioptas2023stellar, 90hyung2023magicapture, 92tang2023retrieving, 34li2023generate] choose to discard the pixels of the background area so that the reconstruction loss can focus on the targeted object and exclude irrelevant disturbances.
+Another technique [85cai2023decoupled] further adds the masked background reconstruction for better disentanglement.
+In addition, the layout indicated by the pixel mask can be incorporated into the attention modules as a supervision signal [54xiao2023fastcomposer, 8han2023svdiff, 74avrahami2023break, 16ma2023subject, 81safaee2023clic, 44rahman2024visual, 90hyung2023magicapture] to adjust the attention’s concentration adaptively.
+Moreover, the mask can stitch specific feature maps to construct more informative semantic patterns [19wei2023elite, 24jia2023taming, 16ma2023subject].
 Feature-level mask.
 In addition to the pixel-level manipulation, masks can be extended to feature-level operations.
-DisenBooth [77] defines an identity-irrelevant embedding with a learnable mask.
+DisenBooth [18chen2023disenbooth] defines an identity-irrelevant embedding with a learnable mask.
 By maximizing the cosine similarity between the identity-preservation embedding and identity-irrelevant embedding, the mask will adaptively exclude the redundant information, and thus the subject appearance can be better preserved.
-AnyDoor [78] defines a high-frequency mask that stores detailed SoI features as a condition for the image generation process.
-Face-Diffuser [79] determines the mask through augmentation from the noise predicted by both a pre-trained text-to-image diffusion model and a PTA personalized model.
+AnyDoor [14chen2023anydoor] defines a high-frequency mask that stores detailed SoI features as a condition for the image generation process.
+Face-Diffuser [52wang2023high] determines the mask through augmentation from the noise predicted by both a pre-trained text-to-image diffusion model and a PTA personalized model.
 Each model makes its own noise prediction, and the final noise output is a composite created through mask-guided concatenation.
@@
 Some methods enhance data diversity through classical image augmentation like blending and spatial rearrangement.
-SVDiff [22] manually constructs mixed images of multiple SoI as new training data, thereby enhancing the model’s exposure to complex scenarios.
-Such concept composition is also used in other works [63, 80, 81].
-BLIP-Diffusion [57] segments the foreground subject and composes it in a random background so that the original text-image pairs are expanded to a larger dataset.
-StyleAdapter [82] chooses to shuffle the image patches to break the irrelevant subject and preserve the desired style.
-PACGen [75] shows that the spatial position entangles with the identity information.
+SVDiff [8han2023svdiff] manually constructs mixed images of multiple SoI as new training data, thereby enhancing the model’s exposure to complex scenarios.
+Such concept composition is also used in other works [16ma2023subject, 138chen2023videodreamer, 145wang2024customvideo].
+BLIP-Diffusion [10li2024blip] segments the foreground subject and composes it in a random background so that the original text-image pairs are expanded to a larger dataset.
+StyleAdapter [26wang2023styleadapter] chooses to shuffle the image patches to break the irrelevant subject and preserve the desired style.
+PACGen [34li2023generate] shows that the spatial position entangles with the identity information.
 Thus rescaling, center crop, and relocation are effective augmentation solutions.
@@
 Generated synthetic data can provide a large amount of training resources when coupled with quality assurance mechanisms.
-SuTI [52] establishes a cascaded pipeline where TTF models first generate diverse variations of each SoI.
+SuTI [11chen2024subject] establishes a cascaded pipeline where TTF models first generate diverse variations of each SoI.
 These synthetic samples then train the target PTA model.
-Similarly, DreamIdentity [48] leverages the existing knowledge of celebrities embedded in large-scale pre-trained diffusion model to generate both the source image and the edited face image.
-StyleDrop [24] and GAL [19] implement iterative refinement pipelines where high-quality synthetic outputs from early training phases are incorporated into subsequent rounds.
+Similarly, DreamIdentity [89chen2023dreamidentity] leverages the existing knowledge of celebrities embedded in large-scale pre-trained diffusion model to generate both the source image and the edited face image.
+StyleDrop [5sohn2024styledrop] and GAL [151zhang2024generative] implement iterative refinement pipelines where high-quality synthetic outputs from early training phases are incorporated into subsequent rounds.
 External sources.
 It is intuitive to leveraging web resources to expand training datasets.
-COTI [83] adopts a scorer network to progressively expand the training set by selecting semantic-relevant samples with high aesthetic quality from a large web-crawled data pool.
+COTI [3yang2023controllable] adopts a scorer network to progressively expand the training set by selecting semantic-relevant samples with high aesthetic quality from a large web-crawled data pool.
 4.4 Regularization
@@
 Auxiliary data regularization.
-To mitigate the overfitting problem where the PCS system consistently produces identical outputs as the references, studies start to use an additional dataset composed of images with the same category of the SoI [4].
+To mitigate the overfitting problem where the PCS system consistently produces identical outputs as the references, studies start to use an additional dataset composed of images with the same category of the SoI [2ruiz2023dreambooth].
 By reconstructing these images, the personalized model is required to generate diverse instances of the class while adapting to the target subject.
-Building on this strategy, StyleBoost [84] introduces an auxiliary style-specific data to separate content and aesthetic adaptation.
-Later, a dataset [40] is curated with detailed textual prompts (specifying attributes/contexts) to improve disentanglement between subject characteristics and background features.
+Building on this strategy, StyleBoost [53park2023styleboost] introduces an auxiliary style-specific data to separate content and aesthetic adaptation.
+Later, a dataset [57he2023data] is curated with detailed textual prompts (specifying attributes/contexts) to improve disentanglement between subject characteristics and background features.
 Text embedding constraints.
@@
 By strategically constraining how subject-specific representations interact with established linguistic concepts in the embedding space, these approaches can achieve better generalization ability.
-For example, Perfusion [21] constrains key projections toward class noun embeddings while learning value projections from subject images.
-Inspired by coached active learning [85, 86], which uses anchor concepts for optimization guidance, Compositional Inversion [87] employs a set of semantically related tokens as anchors to constrain the token embedding search.
-In addition, some works [88, 34] regularize learnable token offsets relative to pre-trained CLIP embeddings.
+For example, Perfusion [7tewel2023key] constrains key projections toward class noun embeddings while learning value projections from subject images.
+Inspired by coached active learning [wei2012coaching, wei2011coached], which uses anchor concepts for optimization guidance, Compositional Inversion [9zhang2023compositional] employs a set of semantically related tokens as anchors to constrain the token embedding search.
+In addition, some works [12gal2023encoder, 51lu2024object] regularize learnable token offsets relative to pre-trained CLIP embeddings.
 By minimizing the offset, the final word embedding is able to achieve better text alignment.
-Similarly, Cones 2 [70] minimizes the offset by reconstructing the features of 1,000 sentences containing the class noun.
-And [33] optimizes the learnable token towards the mean textual embedding of 691 well-known names.
-Domain-Agnostic [89] proposes to use a contrastive loss to guide the SoI text embedding close to its nearest CLIP tokens pre-trained on large-scale samples.
-On the other hand, VICO [71] empirically finds that the end-of-text token <|EOT|> keeps the semantic consistency of SoI.
+Similarly, Cones 2 [62liu2023cones] minimizes the offset by reconstructing the features of 1,000 sentences containing the class noun.
+And [66pang2023cross] optimizes the learnable token towards the mean textual embedding of 691 well-known names.
+Domain-Agnostic [17arar2023domain] proposes to use a contrastive loss to guide the SoI text embedding close to its nearest CLIP tokens pre-trained on large-scale samples.
+On the other hand, VICO [20hao2023vico] empirically finds that the end-of-text token <|EOT|> keeps the semantic consistency of SoI.
 To leverage this discovery, an L2 loss is leveraged to reduce the difference of attention similarity logits between the SoI token and <|EOT|>.
@@
 An important branch in TTF methods is to optimize the learnable token embeddings.
-The first work started from Textual Inversion [7], which applies a simple yet effective method that introduces the unique modifier as a new token to represent the SoI.
+The first work started from Textual Inversion [1gal2022image], which applies a simple yet effective method that introduces the unique modifier as a new token to represent the SoI.
 One of the significant benefits of this method is its minimal storage requirement as the new tokens consumes just a few kilobytes.
@@
 Recent work aims to address these limitations through several directions.
-DVAR [90] improves training efficiency by proposing a clear stopping criterion by removing all randomness to indicate the convergence.
-For enhanced representation, P+ [18] introduces distinct learnable tokens across different layers of the U-Net architecture, thereby offering better attribute control through additional learnable parameters.
-NeTI [30] advances this concept by proposing a neural mapper that adaptively outputs token embeddings based on the denoising timestep and specific U-Net layers.
-ProSpect [31] recommends optimizing multiple token embeddings tailored to different denoising timesteps based on the observation that different types of prompts, like layout, color, structure, and texture, are activated at different stages of the denoising process.
-Similarly, a study by [91] shows layered activation insight to learn distinct attributes by selectively activating the tokens within their respective scopes.
-Later, HiFiTuner [32] integrates multiple techniques into the learnable token, including mask-guided loss function, parameter regularization, time-dependent embedding, and generation refinement assisted by the nearest reference.
-Alternative approaches, like DreamArtist [92], choose to optimize both negative and positive prompt embeddings to refine the detail preservation.
-In addition to these token-level refinement approaches, the field continues to evolve with novel techniques such as InstructBooth’s reinforcement learning framework [93] and gradient-free evolutionary optimization [94].
+DVAR [69voronov2024loss] improves training efficiency by proposing a clear stopping criterion by removing all randomness to indicate the convergence.
+For enhanced representation, P+ [23voynov2023p+] introduces distinct learnable tokens across different layers of the U-Net architecture, thereby offering better attribute control through additional learnable parameters.
+NeTI [61alaluf2023neural] advances this concept by proposing a neural mapper that adaptively outputs token embeddings based on the denoising timestep and specific U-Net layers.
+ProSpect [30zhang2023prospect] recommends optimizing multiple token embeddings tailored to different denoising timesteps based on the observation that different types of prompts, like layout, color, structure, and texture, are activated at different stages of the denoising process.
+Similarly, a study by [79agarwal2023image] shows layered activation insight to learn distinct attributes by selectively activating the tokens within their respective scopes.
+Later, HiFiTuner [48wang2023hifi] integrates multiple techniques into the learnable token, including mask-guided loss function, parameter regularization, time-dependent embedding, and generation refinement assisted by the nearest reference.
+Alternative approaches, like DreamArtist [68dong2022dreamartist], choose to optimize both negative and positive prompt embeddings to refine the detail preservation.
+In addition to these token-level refinement approaches, the field continues to evolve with novel techniques such as InstructBooth’s reinforcement learning framework [56chae2023instructbooth] and gradient-free evolutionary optimization [73fei2023gradient].
 In summary, recent developments following the foundational work of Textual Inversion focus on reducing training times and enhancing the visual quality of generated images.
 In the realm of TTF methods for PCS, there is a clear shift towards fine-tuning model weights rather than just the token embeddings.
-This approach often addresses the limitations where token embeddings alone struggle to capture complex semantics uncovered in the pre-training data [25, 95].
-DreamBooth [4] proposes to use a unique modifier by a rare token to represent the SoI and fine-tune the whole parameters of the diffusion model.
+This approach often addresses the limitations where token embeddings alone struggle to capture complex semantics uncovered in the pre-training data [46gu2024mix, 80zhao2023catversion].
+DreamBooth [2ruiz2023dreambooth] proposes to use a unique modifier by a rare token to represent the SoI and fine-tune the whole parameters of the diffusion model.
 Besides, a regularization dataset containing 20-30 images with the same category as SoI is adopted to overcome the overfitting problem.
@@
 However, fine-tuning the entire model for each new object causes considerable storage costs, potentially rendering widespread application.
-To address this, Custom Diffusion [20] focuses on identifying and fine-tuning critical parameters, particularly the key-value projections in cross-attention layers, to achieve a balance of visual fidelity and storage efficiency.
-Further approach, Perfusion [21], also adopts the cross-attention fine-tuning and proposes to regularize the update direction of the K (key) projection towards the super-category token embedding and the V (value) projection towards the learnable token embedding.
-COMCAT [96] introduces a low-rank approximation of attention matrices, which drastically reduces storage requirements to 6 MB while maintaining high fidelity in the outputs.
-Additionally, methods like adapters [23, 24] and LoRA variants [25, 26, 27, 34, 35, 36] are increasingly utilized in personalized generation for parameter-efficient fine-tuning.
+To address this, Custom Diffusion [6kumari2023multi] focuses on identifying and fine-tuning critical parameters, particularly the key-value projections in cross-attention layers, to achieve a balance of visual fidelity and storage efficiency.
+Further approach, Perfusion [7tewel2023key], also adopts the cross-attention fine-tuning and proposes to regularize the update direction of the K (key) projection towards the super-category token embedding and the V (value) projection towards the learnable token embedding.
+COMCAT [75xiao2023comcat] introduces a low-rank approximation of attention matrices, which drastically reduces storage requirements to 6 MB while maintaining high fidelity in the outputs.
+Additionally, methods like adapters [72xiang2023closer, 5sohn2024styledrop] and LoRA variants [46gu2024mix, 58ruiz2023hyperdreambooth, 115kong2024omg, 51lu2024object, 95achlioptas2023stellar, 76yeh2023navigating] are increasingly utilized in personalized generation for parameter-efficient fine-tuning.
 It is worth noting that the token embedding fine-tuning is compatible with diffusion weight fine-tuning.
-Multiple methods [37, 38, 39] have started using combined weight fine-tuning.
+Multiple methods [90hyung2023magicapture, 74avrahami2023break, 86hua2023dreamtuner] have started using combined weight fine-tuning.
 PTA framework.
@@
 PTA methods enable real-time generation (less than 10 sec/subject) by leveraging large-scale pre-training to avoid per-subject optimization during the inference phase.
-Re-Imagen [97] introduces a retrieval-augmented generative approach, which leverages features from text-image pairs retrieved via a specific prompt.
+Re-Imagen [67chen2022re] introduces a retrieval-augmented generative approach, which leverages features from text-image pairs retrieved via a specific prompt.
 While it is not specifically tailored for object personalization, it demonstrates the feasibility of training reference-conditioned frameworks.
-Later, ELITE [47] specifically targets image personalization by combining the global reference features with text embedding while incorporating local features that exclude irrelevant backgrounds.
+Later, ELITE [19wei2023elite] specifically targets image personalization by combining the global reference features with text embedding while incorporating local features that exclude irrelevant backgrounds.
 Both fused features and local features serve as conditions for the denoising process.
-Similarly, InstantBooth [98] retrains CLIP models to extract image features and patch features, which are injected into the diffusion model via the attention mechanism and learnable adapter, respectively.
-Additionally, UMM-Diffusion [61] designs a multi-modal encoder that produces fused features based on the reference image and text prompt.
+Similarly, InstantBooth [22shi2023instantbooth] retrains CLIP models to extract image features and patch features, which are injected into the diffusion model via the attention mechanism and learnable adapter, respectively.
+Additionally, UMM-Diffusion [71ma2023unified] designs a multi-modal encoder that produces fused features based on the reference image and text prompt.
 The text features and multi-modal hidden state are seen as guidance signals to predict a mixed noise.
-Another work, SuTI [52], adopts the same architecture as Re-Imagen.
+Another work, SuTI [11chen2024subject], adopts the same architecture as Re-Imagen.
 The difference lies in the training samples which are produced by a massive number of TTF models, each tuned on a particular subject set.
 This strategy promotes a more precise alignment with personalization at an instance level rather than the class level of Re-Imagen.
-Moreover, Domain-Agnostic [89] combines a contrastive-based regularization technique to push the pseudo embedding produced by the image encoder towards the existing nearest pre-trained token.
+Moreover, Domain-Agnostic [17arar2023domain] combines a contrastive-based regularization technique to push the pseudo embedding produced by the image encoder towards the existing nearest pre-trained token.
 Besides, They introduce a dual-path attention module separately conditioned on the nearest token and pseudo embedding.
 Compared to the methods that use separate encoders to process a single modality, some works have explored the usage of pre-trained multi-modal large language models (MLLM) that can process text and image modality within a unified framework.
-For example, BLIP-Diffusion [57] utilizes the pre-trained BLIP2 [62] that encodes multimodal inputs including the SoI reference and a class noun.
+For example, BLIP-Diffusion [10li2024blip] utilizes the pre-trained BLIP2 [li2023blip] that encodes multimodal inputs including the SoI reference and a class noun.
 The output embedding is then concatenated with context description and serves as a condition to generate images.
-Further, Customization Assistant [58] and KOSMOS-G [60] replace the text encoder of Stable Diffusion with a pre-trained MLLM to output a fused feature based on the reference and context description.
+Further, Customization Assistant [84zhou2023customization] and KOSMOS-G [77pan2023kosmos] replace the text encoder of Stable Diffusion with a pre-trained MLLM to output a fused feature based on the reference and context description.
 Meanwhile, to meet the standard format of Stable Diffusion, a network is trained to align the dimension of the output embedding.
@@
 TTF framework.
-In this field, StyleDrop [24] leverages adapter tuning to efficiently capture the style from a single reference image.
+In this field, StyleDrop [5sohn2024styledrop] leverages adapter tuning to efficiently capture the style from a single reference image.
 This method demonstrates the effectiveness through iterative training, utilizing synthesized images refined by feedback mechanisms, like human evaluations and CLIP scores.
 This approach not only enhances style learning but also ensures that the generated styles align closely with human aesthetic judgments.
-Later, GAL [19] proposes an uncertainty-based evaluation strategy to filter high-quality synthetic-style data and uses a weighted schema to balance the contribution of the additional samples and the original reference.
-Furthermore, StyleAligned [99] focuses on maintaining stylistic consistency across a batch of images.
+Later, GAL [151zhang2024generative] proposes an uncertainty-based evaluation strategy to filter high-quality synthetic-style data and uses a weighted schema to balance the contribution of the additional samples and the original reference.
+Furthermore, StyleAligned [98hertz2023style] focuses on maintaining stylistic consistency across a batch of images.
 This is achieved by using the first image as a reference, which acts as an additional key and value in the self-attention layers, ensuring that all subsequent images in the batch adhere to the same stylistic guidelines.
-Style-friendly [100] introduces a novel diffusion model fine-tuning approach that enhances personalized artistic style generation by adaptively biasing noise sampling toward higher noise levels, where stylistic features emerge.
+Style-friendly [211choi2024style] introduces a novel diffusion model fine-tuning approach that enhances personalized artistic style generation by adaptively biasing noise sampling toward higher noise levels, where stylistic features emerge.
 PTA framework.
-For the PTA framework, StyleAdapter [82] employs a dual-path cross-attention mechanism within the PTA framework.
+For the PTA framework, StyleAdapter [26wang2023styleadapter] employs a dual-path cross-attention mechanism within the PTA framework.
 This model introduces a specialized embedding module designed to extract and integrate global features from multiple style references.
-Diptych Prompting [101] utilizes an inpainting mechanism to draw another image with the same style of the reference part.
+Diptych Prompting [209shin2024large] utilizes an inpainting mechanism to draw another image with the same style of the reference part.
 5.3 Personalized Face Generation
@@
 Compared to general object personalization, the scope is narrowed to a specific class, humans.
-An obvious benefit is that one can readily leverage large-scale human-centric datasets [67, 102, 103] and utilize pre-trained models in well-developed areas, like face landmark detection [104] and face recognition [105].
+An obvious benefit is that one can readily leverage large-scale human-centric datasets [zheng2022general, yang2020facescape, cao2018vggface2] and utilize pre-trained models in well-developed areas, like face landmark detection [wu2019facial] and face recognition [adjabi2020past].
 TTF framework.
-Regarding TTF methods, PromptNet [106] trains a diffusion-based network that encodes the input image and noisy latent to a word embedding.
+Regarding TTF methods, PromptNet [55zhou2023enhancing] trains a diffusion-based network that encodes the input image and noisy latent to a word embedding.
 To alleviate the overfitting problem, the noises predicted by the word embedding and context description are balanced through fusion sampling in classifier-free guidance.
-Additionally, Celeb Basis [65] provides a novel idea that the personalized ID can be viewed as the composition of celebrity’s face, which has been learned by the pre-trained diffusion model.
+Additionally, Celeb Basis [60yuan2023inserting] provides a novel idea that the personalized ID can be viewed as the composition of celebrity’s face, which has been learned by the pre-trained diffusion model.
 Based on this hypothesis, a simple MLP is optimized at test time to transform face features into the weighting of different celebrity name embeddings.
@@
 Thanks to the abundance of available datasets featuring the same individual in various contexts, which provide valuable data for pre-training PTA methods, the number of works in PTA frameworks is rapidly increasing.
-Face0 [107] crops the face region to extract refined embeddings and concatenates them with text features.
+Face0 [88valevski2023face0] crops the face region to extract refined embeddings and concatenates them with text features.
 During the sampling phase, the output of classifier-free guidance is replaced by a weighted combination of the noise patterns predicted by face-only embedding, text-only embedding, and concatenated face-text embedding.
-The $\mathcal{W}+$ Adapter [53] constructs a mapping network and residual cross-attention modules to transform the facial features from the StyleGAN [108] $\mathcal{W}+$ space into the text embedding space of Stable Diffusion.
-FaceStudio [109] adapts the cross-attention layer to support hybrid guidance including stylized images, facial images, and textual prompts.
-Moreover, PhotoMaker [46] constructs a high-quality dataset through a meticulous data collection and filtering pipeline.
+The $\mathcal{W}+$ Adapter [91li2023stylegan] constructs a mapping network and residual cross-attention modules to transform the facial features from the StyleGAN [karras2019style] $\mathcal{W}+$ space into the text embedding space of Stable Diffusion.
+FaceStudio [31yan2023facestudio] adapts the cross-attention layer to support hybrid guidance including stylized images, facial images, and textual prompts.
+Moreover, PhotoMaker [28li2023photomaker] constructs a high-quality dataset through a meticulous data collection and filtering pipeline.
 They use a two-layer MLP to fuse ID features and class embeddings for an overall representation of human portrait.
-PortraitBooth [50] also employs a simple MLP, which fuses the text condition and shallow features of a pre-trained face recognition model.
+PortraitBooth [96peng2023portraitbooth] also employs a simple MLP, which fuses the text condition and shallow features of a pre-trained face recognition model.
 To ensure expression manipulation and facial fidelity, they add another expression token and incorporate the identity preservation loss and mask-based cross-attention loss.
-InstantID [6] additionally introduces a variant of ControlNet that takes facial landmarks as input, providing stronger guiding signals compared to the methods that solely rely on attention fusion.
+InstantID [37wang2024instantid] additionally introduces a variant of ControlNet that takes facial landmarks as input, providing stronger guiding signals compared to the methods that solely rely on attention fusion.
 5.4 Multiple Subject Composition
@@
 Some works focus on the one-for-one generation following a fusion mechanism.
-For instance, Custom Diffusion [20] proposes a constrained optimization method to merge the cross-attention key-value projection weights with the goal of maximizing reconstruction performance for each subject.
-Mix-of-Show [25] fuses the LoRA [110] weights with the same optimization objective.
-StyleDrop [24] dynamically summarizes noise predictions from each personalized diffusion model.
-In OMG [27], the latent predicted by each LoRA-tuned model is spatially composited using the subject mask.
+For instance, Custom Diffusion [6kumari2023multi] proposes a constrained optimization method to merge the cross-attention key-value projection weights with the goal of maximizing reconstruction performance for each subject.
+Mix-of-Show [46gu2024mix] fuses the LoRA [hu2021lora] weights with the same optimization objective.
+StyleDrop [5sohn2024styledrop] dynamically summarizes noise predictions from each personalized diffusion model.
+In OMG [115kong2024omg], the latent predicted by each LoRA-tuned model is spatially composited using the subject mask.
 Joint training is another strategy to cover all expected subjects.
-SVDiff [22] employs a data augmentation method called Cut-Mix to compose several subjects together and applies a location loss to regularize attention maps, ensuring alignment between each subject and its corresponding token.
-Similar strategies are found in other works [20, 65] which train a single model by reconstructing the appearance of every SoI.
+SVDiff [8han2023svdiff] employs a data augmentation method called Cut-Mix to compose several subjects together and applies a location loss to regularize attention maps, ensuring alignment between each subject and its corresponding token.
+Similar strategies are found in other works [6kumari2023multi, 60yuan2023inserting] which train a single model by reconstructing the appearance of every SoI.
 There are advanced control mechanisms designed to manage multiple subjects.
-Cones [111] proposes to find a small cluster of neurons that preserve the most information about SoI.
+Cones [4liu2023cones] proposes to find a small cluster of neurons that preserve the most information about SoI.
 The neurons belonging to different SoI will be simultaneously activated to generate the combination.
-Compositional Inversion [87] introduces spatial region assignment to different subjects to improve the composition success rate.
+Compositional Inversion [9zhang2023compositional] introduces spatial region assignment to different subjects to improve the composition success rate.
 PTA framework.
 For PTA frameworks, multi-subject generation is achieved through specialized architectural designs.
-Fastcomposer [5], Subject-Diffusion [63], and $\lambda$-eclipse [112] place each subject feature in its corresponding placeholder in the text embedding, ensuring a seamless and efficient combination.
-CustomNet [113] and MIGC [114] train a PTA network that supports location control for each subject.
-SSR-Encoder [55] implements an encoder to selectively preserve the desired subject feature and a cross-attention module to support multi-subject feature fusion.
+Fastcomposer [54xiao2023fastcomposer], Subject-Diffusion [16ma2023subject], and $\lambda$-eclipse [222patel2024lambda] place each subject feature in its corresponding placeholder in the text embedding, ensuring a seamless and efficient combination.
+CustomNet [29yuan2023customnet] and MIGC [174zhou2024migc] train a PTA network that supports location control for each subject.
+SSR-Encoder [221zhang2024ssr] implements an encoder to selectively preserve the desired subject feature and a cross-attention module to support multi-subject feature fusion.
 5.5 High-level Semantic Personalization
@@
 For now, all researches in this field are based on the TTF framework.
-ReVersion [115] intends to invert object relations from references.
+ReVersion [63huang2023reversion] intends to invert object relations from references.
 Specifically, they use a contrastive loss to guide the optimization of the token embedding towards specific clusters of Part-of-Speech tags, such as prepositions, nouns, and verbs.
 Meanwhile, they also increase the likelihood of adding noise at larger timesteps during the training process to emphasize the extraction of high-level semantic features.
-Lego [116] focuses on more general concepts, such as adjectives, which are frequently intertwined with the subject appearance.
+Lego [45motamed2023lego] focuses on more general concepts, such as adjectives, which are frequently intertwined with the subject appearance.
 The concept can be learned from a contrastive loss applied to the dataset comprising clean subject images and images that embody the desired adjectives.
-Moreover, ADI [117] aims to learn the action-specific identifier from the references.
+Moreover, ADI [43huang2023learning] aims to learn the action-specific identifier from the references.
 To ensure the inversion only focuses on the desired action, ADI extracts gradient invariance from a constructed triplet sample and applies a threshold to mask out the irrelevant feature channels.
@@
 One common application is to customize the subject into a fixed source image.
-For example, PhotoSwap [118] introduces a new task that replaces the subject in the source image with the SoI from the reference image.
+For example, PhotoSwap [15gu2024photoswap] introduces a new task that replaces the subject in the source image with the SoI from the reference image.
 To address this requirement, they first fine-tunes a diffusion model on the references to obtain a personalized model.
-To preserve the background of the source image, they initialize the noise with DDIM inversion [119] and replace the intermediate feature maps with those derived from source image generation during inference time.
-Later, MagiCapture [53] broadens the scope to face customization.
+To preserve the background of the source image, they initialize the noise with DDIM inversion [song2020denoising] and replace the intermediate feature maps with those derived from source image generation during inference time.
+Later, MagiCapture [91li2023stylegan] broadens the scope to face customization.
 Another similar application can be found in Virtual Try-on, which aims to fit selected clothing onto a target person.
-The complexities of this task have been thoroughly analyzed in another survey [120].
-Additional conditions in personalization tasks may include adjusting the layout [69], transforming sketches [121], controlling viewpoint [113, 122], or modifying poses [6].
+The complexities of this task have been thoroughly analyzed in another survey [islam2024deep].
+Additional conditions in personalization tasks may include adjusting the layout [152chen2024training], transforming sketches [94chen2023democaricature], controlling viewpoint [29yuan2023customnet, 163kumari2024customizing], or modifying poses [37wang2024instantid].
 Each of these conditions presents unique challenges and requires specialized approaches to integrate these elements seamlessly into the personalized content.
 6.2 Personalized Video Generation
-With the rising popularity of video generation [123], video personalization has also begun to attract attention.
+With the rising popularity of video generation [xing2024survey], video personalization has also begun to attract attention.
 In video personalization, the SoI can be classified into three distinct categories: appearance, motion, and the combination of both appearance and motion.
@@
 The standard TTF pipeline utilizes reference images as appearance anchors and fine-tunes video diffusion models (VDM) for temporal synthesis.
-The process involves leveraging sophisticated methods from 2D personalization, such as parameter-efficient fine-tuning [81], data augmentation [80, 81], and attention manipulation [124, 54, 81].
-Additionally, several studies [125, 124, 54, 126] have explored the PTA framework.
+The process involves leveraging sophisticated methods from 2D personalization, such as parameter-efficient fine-tuning [145wang2024customvideo], data augmentation [138chen2023videodreamer, 145wang2024customvideo], and attention manipulation [83jiang2023videobooth, 212liu2024stylecrafter, 145wang2024customvideo].
+Additionally, several studies [82zhao2023videoassembler, 83jiang2023videobooth, 212liu2024stylecrafter, 105feng2023dreamoving] have explored the PTA framework.
 These diffusion models are specifically tailored to synthesize videos based on the image references.
@@
 In this task, the reference input switches to a video clip containing a consistent action.
-A common approach is to fine-tune the video diffusion model by reconstructing the action clip [127, 128, 129, 130, 131, 132, 133].
+A common approach is to fine-tune the video diffusion model by reconstructing the action clip [150wu2023tune, 103song2023save, 140jeong2023vmc, 148esser2023structure, 136he2023animate, 102wu2023lamp, 104materzynska2023customizing].
 However, distinguishing between appearance and motion within the reference video can be challenging.
-To solve this problem, SAVE [128] applies appearance learning to ensure that appearance is excluded from the motion learning phase.
-Additionally, VMC [129] removes the background information during training prompt construction.
+To solve this problem, SAVE [103song2023save] applies appearance learning to ensure that appearance is excluded from the motion learning phase.
+Additionally, VMC [140jeong2023vmc] removes the background information during training prompt construction.
 Appearance and motion personalization.
 When integrating both subject appearance and motion, innovative methods are employed to address the complexities of learning both aspects simultaneously.
-MotionDirector [134] utilizes spatial and temporal losses to facilitate learning across these dimensions.
-Another approach, DreamVideo [135], incorporates residual features from a randomly selected frame to emphasize subject information.
+MotionDirector [101zhao2023motiondirector] utilizes spatial and temporal losses to facilitate learning across these dimensions.
+Another approach, DreamVideo [142wei2023dreamvideo], incorporates residual features from a randomly selected frame to emphasize subject information.
 This technique enables the fine-tuned module to primarily focus on learning motion dynamics.
@@
 Basically, the pipeline begins by fine-tuning a 2D diffusion model using TTF methods.
-This tuned model then utilize Score Distillation Sampling (SDS) [136] to train a 3D Neural Radiance Field (NeRF) model [137] for each specific prompt [138, 139].
+This tuned model then utilize Score Distillation Sampling (SDS) [poole2022dreamfusion] to train a 3D Neural Radiance Field (NeRF) model [mildenhall2021nerf] for each specific prompt [120lin2023magic3d, 126shi2023mvdream].
 Building on this foundation, several methods have been developed to improve the workflow.
-DreamBooth3D [140] structures the process into three phases: initializing and optimizing a NeRF from a DreamBooth model, rendering multi-view images, and fine-tuning a secondary DreamBooth for the final 3D NeRF refinement.
-Consist3D [141] enhances text embeddings by training two distinct tokens, a semantic and a geometric token, during 3D model optimization.
-TextureDreamer [142] focuses on extracting texture maps from optimized spatially-varying bidirectional reflectance distribution (BRDF) fields for rendering texture on wide-range 3D subjects.
+DreamBooth3D [121raj2023dreambooth3d] structures the process into three phases: initializing and optimizing a NeRF from a DreamBooth model, rendering multi-view images, and fine-tuning a secondary DreamBooth for the final 3D NeRF refinement.
+Consist3D [127ouyang2023chasing] enhances text embeddings by training two distinct tokens, a semantic and a geometric token, during 3D model optimization.
+TextureDreamer [133yeh2024texturedreamer] focuses on extracting texture maps from optimized spatially-varying bidirectional reflectance distribution (BRDF) fields for rendering texture on wide-range 3D subjects.
 Additionally, advancements extend to 3D avatar rendering and dynamic scenes.
-Animate124 [143] and Dream-in-4D [144] integrate video diffusion for 4D dynamic scene support within the 3D optimization process.
-In avatar rendering, PAS [145] generates 3D body poses configurable by avatar settings, StyleAvatar3D [146] facilitates 3D avatar generation based on images, and AvatarBooth [147] employs dual fine-tuned diffusion models for separate face and body generation.
+Animate124 [128zhao2023animate124] and Dream-in-4D [130zheng2023unified] integrate video diffusion for 4D dynamic scene support within the 3D optimization process.
+In avatar rendering, PAS [122azadi2023text] generates 3D body poses configurable by avatar settings, StyleAvatar3D [123zhang2023styleavatar3d] facilitates 3D avatar generation based on images, and AvatarBooth [124zeng2023avatarbooth] employs dual fine-tuned diffusion models for separate face and body generation.
 6.4 Attack and Defense
@@
 There is an increased risk that individuals or organizations may exploit them to produce misleading content or manipulate public perception.
-To mitigate this, Anti-DreamBooth [148] aims to add a subtle noise perturbation to the references so that any personalized model trained on these samples only produces terrible results.
+To mitigate this, Anti-DreamBooth [13van2023anti] aims to add a subtle noise perturbation to the references so that any personalized model trained on these samples only produces terrible results.
 The basic idea is to maximize the reconstruction loss of the surrogate model.
-Additionally, [149] suggests to predefine a collection of trigger words and meaningless images.
+Additionally, [50zhang2023backdooring] suggests to predefine a collection of trigger words and meaningless images.
 These data are paired and incorporated during the training phase.
@@
 Several works are exploring different personalization extensions.
-For example, SVG personalization is introduced by [150], in which a parameter-efficient fine-tuning method is applied to create SVGs.
+For example, SVG personalization is introduced by [36zhang2023text], in which a parameter-efficient fine-tuning method is applied to create SVGs.
 After first-step generation, the SVGs are refined through a process that includes semantic alignment and a dual optimization approach, which utilizes both image-level and vector-level losses to enhance the final output.
-Another application, 360-degree panorama customization [151], is also emerging as a potential tool for personalization in the digital imaging realm.
+Another application, 360-degree panorama customization [25wang2024customizing], is also emerging as a potential tool for personalization in the digital imaging realm.
 7 Evaluation
@@
 To assess the performance of personalized models, various datasets have been developed:
-DreamBench [4] serves as the primary evaluation benchmark for DreamBooth [4], containing 30 diverse subjects (e.g., backpacks, animals, vehicles, and toys) with 25 unique prompts per subject.
-DreamBench-v2 [52] expands the evaluation scope of DreamBench by adding 220 test prompts for each subject.
-Custom-10 [20] used in Custom Diffusion [20] evaluates 10 subjects, each with 20 specific test prompts, and includes tests for multi-subject composition with 5 pairs of subjects and 8 prompts for each pair.
-Custom-101 [20] is the latest released dataset by the authors of Custom Diffusion [20], which comprises 101 subjects to provide a broader scope of evaluation.
-Stellar [35] specifically targets human-centric evaluation, featuring 20,000 prompts on 400 human identities.
+DreamBench [2ruiz2023dreambooth] serves as the primary evaluation benchmark for DreamBooth [2ruiz2023dreambooth], containing 30 diverse subjects (e.g., backpacks, animals, vehicles, and toys) with 25 unique prompts per subject.
+DreamBench-v2 [11chen2024subject] expands the evaluation scope of DreamBench by adding 220 test prompts for each subject.
+Custom-10 [6kumari2023multi] used in Custom Diffusion [6kumari2023multi] evaluates 10 subjects, each with 20 specific test prompts, and includes tests for multi-subject composition with 5 pairs of subjects and 8 prompts for each pair.
+Custom-101 [6kumari2023multi] is the latest released dataset by the authors of Custom Diffusion [6kumari2023multi], which comprises 101 subjects to provide a broader scope of evaluation.
+Stellar [95achlioptas2023stellar] specifically targets human-centric evaluation, featuring 20,000 prompts on 400 human identities.
 Despite these contributions, they remain fragmented across different research groups and the research community still lacks a benchmark tested on a large number of personalized generation tasks.
@@
 The text alignment metrics quantify how precisely generated outputs reflect prompt semantics:
-CLIP-T measures semantic alignment using the cosine similarity between CLIP [42] embeddings of generated images and their text prompts.
-ImageReward [152], HPS Score (v1/v2) [153, 154], and PickScore [155] employ learned models trained on human judgments to better correlate with perceptual quality to reflect human preference better.
+CLIP-T measures semantic alignment using the cosine similarity between CLIP [radford2021learning] embeddings of generated images and their text prompts.
+ImageReward [xu2023imagereward], HPS Score (v1/v2) [wu2023human, wu2023humanv2], and PickScore [kirstain2023pick] employ learned models trained on human judgments to better correlate with perceptual quality to reflect human preference better.
 To determine how closely the generated subject resembles the SoI, visual fidelity can be assessed via the following metrics:
@@
 Optimal values should balance fidelity (high scores) against overfitting (excessive scores that ignore text guidance).
-DINO-I [156] provides a complementary assessment using DINO’s instance-aware features, particularly effective for object-level similarity.
-Fréchet Inception Distance (FID) [157] quantifies the statistical similarity between generated and real image distributions through Inception-V3 [158].
-In addition to these commonly adopted metrics, there are some discussions of specialized metrics for PCS system evaluation: LyCORIS [36] introduces a 5-dimensional assessment covering fidelity, controllability, diversity, base model preservation, and image quality.
-Stellar [35] develops six human-centric metrics including soft-penalized CLIP text score, Identity Preservation Score, Attribute Preservation Score, Stability of Identity Score, Grounding Objects Accuracy, and Relation Fidelity Score.
+DINO-I [caron2021emerging] provides a complementary assessment using DINO’s instance-aware features, particularly effective for object-level similarity.
+Fréchet Inception Distance (FID) [FIDheusel2017gans] quantifies the statistical similarity between generated and real image distributions through Inception-V3 [inception].
+In addition to these commonly adopted metrics, there are some discussions of specialized metrics for PCS system evaluation: LyCORIS [76yeh2023navigating] introduces a 5-dimensional assessment covering fidelity, controllability, diversity, base model preservation, and image quality.
+Stellar [95achlioptas2023stellar] develops six human-centric metrics including soft-penalized CLIP text score, Identity Preservation Score, Attribute Preservation Score, Stability of Identity Score, Grounding Objects Accuracy, and Relation Fidelity Score.
 These metrics ensure a structured and detailed evaluation of personalized models.
@@
 For Object.
-Persona includes 47 subjects from available resources [7, 4, 20].
-Following the methodology of DreamBooth [4], we categorize the subjects into two classes: objects and live pets, based on whether the subject is a living entity.
+Persona includes 47 subjects from available resources [1gal2022image, 2ruiz2023dreambooth, 6kumari2023multi].
+Following the methodology of DreamBooth [2ruiz2023dreambooth], we categorize the subjects into two classes: objects and live pets, based on whether the subject is a living entity.
 Specifically, 10 out of 47 subjects are pets and the remaining 37 are various objects.
-To evaluate the performance, we utilize the text prompts from DreamBooth [4].
+To evaluate the performance, we utilize the text prompts from DreamBooth [2ruiz2023dreambooth].
 This includes 20 recontextualization prompts and 5 property modification prompts for objects, along with 10 recontextualization, 10 accessorization, and 5 property modification prompts for pets, totally 25 prompts per category.
 For Face.
-We also collected 15 subjects from Celeb-A [159] into our Persona dataset.
+We also collected 15 subjects from Celeb-A [liu2018large] into our Persona dataset.
 We use 40 prompts for evaluation, including 10 accessory prompts, 10 style prompts, 10 action prompts, and 10 context prompts.
@@
 We calculate CLIP-I for subject fidelity in object generation.
-For subject fidelity in face generation, we detect faces in both the generated and target images using MTCNN [160] and calculate the pairwise identity similarity using FaceNet [161].
+For subject fidelity in face generation, we detect faces in both the generated and target images using MTCNN [zhang2016mtcnn] and calculate the pairwise identity similarity using FaceNet [schroff2015facenet].
 To uniformly evaluate and compare existing personalized generation methods, we select 22 representative PCS methods for evaluation.
@@
 Type Methods Framework Backbone CLIP-T CLIP-I
-Object Textual Inversion [7] TTF SD 1.5 0.199 0.749
-Dreambooth [4] TTF SD 1.5 0.286 0.772
-P+ [18] TTF SD 1.4 0.244 0.643
-Custom Diffusion [20] TTF SD 1.4 0.307 0.722
-NeTI [30] TTF SD 1.4 0.283 0.801
-SVDiff [22] TTF SD 1.5 0.282 0.776
-Perfusion [21] TTF SD 1.5 0.273 0.691
-ELITE [47] PTA SD 1.4 0.292 0.765
-BLIP-Diffusion [57] PTA SD 1.5 0.292 0.772
-IP-Adapter [51] PTA SD 1.5 0.2722 0.825
-SSR Encoder [55] PTA SD 1.5 0.288 0.792
-MoMA [59] PTA SD 1.5 0.322 0.748
-Diptych Prompting [101] PTA FLUX 1.0 dev 0.327 0.722
-$\lambda$-eclipse [112] PTA Kandinsky 2.2 0.272 0.824
-MS-Diffusion [56] PTA SDXL 0.298 0.777
-Face CrossInitialization [33] TTF SD 2.1 0.261 0.469
-Face2Diffusion [162] PTA SD 1.4 0.265 0.588
-SSR Encoder [55] PTA SD 1.5 0.233 0.490
-FastComposer [5] PTA SD 1.5 0.230 0.516
-IP-Adapter [51] PTA SD 1.5 0.292 0.462
-IP-Adapter [51] PTA SDXL 0.292 0.642
-PhotoMaker [46] PTA SDXL 0.311 0.547
-InstantID [6] PTA SDXL 0.278 0.707
+Object Textual Inversion [1gal2022image] TTF SD 1.5 0.199 0.749
+Dreambooth [2ruiz2023dreambooth] TTF SD 1.5 0.286 0.772
+P+ [23voynov2023p+] TTF SD 1.4 0.244 0.643
+Custom Diffusion [6kumari2023multi] TTF SD 1.4 0.307 0.722
+NeTI [61alaluf2023neural] TTF SD 1.4 0.283 0.801
+SVDiff [8han2023svdiff] TTF SD 1.5 0.282 0.776
+Perfusion [7tewel2023key] TTF SD 1.5 0.273 0.691
+ELITE [19wei2023elite] PTA SD 1.4 0.292 0.765
+BLIP-Diffusion [10li2024blip] PTA SD 1.5 0.292 0.772
+IP-Adapter [40ye2023ip] PTA SD 1.5 0.2722 0.825
+SSR Encoder [221zhang2024ssr] PTA SD 1.5 0.288 0.792
+MoMA [song2024moma] PTA SD 1.5 0.322 0.748
+Diptych Prompting [209shin2024large] PTA FLUX 1.0 dev 0.327 0.722
+$\lambda$-eclipse [222patel2024lambda] PTA Kandinsky 2.2 0.272 0.824
+MS-Diffusion [220personalizationms] PTA SDXL 0.298 0.777
+Face CrossInitialization [66pang2023cross] TTF SD 2.1 0.261 0.469
+Face2Diffusion [170shiohara2024face2diffusion] PTA SD 1.4 0.265 0.588
+SSR Encoder [221zhang2024ssr] PTA SD 1.5 0.233 0.490
+FastComposer [54xiao2023fastcomposer] PTA SD 1.5 0.230 0.516
+IP-Adapter [40ye2023ip] PTA SD 1.5 0.292 0.462
+IP-Adapter [40ye2023ip] PTA SDXL 0.292 0.642
+PhotoMaker [28li2023photomaker] PTA SDXL 0.311 0.547
+InstantID [37wang2024instantid] PTA SDXL 0.278 0.707
 8 Challenge and Outlook
@@
 The irrelevant elements in the references are generated in the output, such as backgrounds or objects not pertinent to the current context.
-To investigate the rationale behind, Compositional Inversion [87] observes that the learned token embedding is located in an out-of-distribution area compared to the center distribution formed by pre-trained words.
-This is also found in another work [33] that the learnable token embeddings deviate significantly from the distribution of the initial embedding.
-In addition, there is evidence [87, 76, 71] suggesting that the unique modifier dominates in the cross-attention layers compared to the other context tokens, leading to the absence of other semantic appearances.
+To investigate the rationale behind, Compositional Inversion [9zhang2023compositional] observes that the learned token embedding is located in an out-of-distribution area compared to the center distribution formed by pre-trained words.
+This is also found in another work [66pang2023cross] that the learnable token embeddings deviate significantly from the distribution of the initial embedding.
+In addition, there is evidence [9zhang2023compositional, 24jia2023taming, 20hao2023vico] suggesting that the unique modifier dominates in the cross-attention layers compared to the other context tokens, leading to the absence of other semantic appearances.
 To address this issue, many solutions have been proposed.
 Most methods discussed in Section 4 contribute to the alleviation of the overfitting problem, such as the exclusion of redundant background, attention manipulation, regularization of the learnable parameters, and data augmentation.
-However, it has not been solved yet, especially in the cases where the SoI has a non-rigid appearance [87] or the context prompt has a similar semantic correlation with the irrelevant elements in the reference [19].
+However, it has not been solved yet, especially in the cases where the SoI has a non-rigid appearance [9zhang2023compositional] or the context prompt has a similar semantic correlation with the irrelevant elements in the reference [151zhang2024generative].
 It is clear that addressing overfitting in PCS is not merely a technical challenge but a necessity for ensuring the practical deployment and scalability of these systems in varied and dynamic real-world environments.
@@
 As a result, it becomes challenging to achieve flexible adaptation in different contexts while simultaneously pushing the model to capture fine-grained details.
-To address this inherent conflict, Perfusion [21] proposes to regularize the attention projections by these two items.
-[163] decouples the conditional guidance into two separate processes, which allows for the distinct handling of subject fidelity and textual alignment.
+To address this inherent conflict, Perfusion [7tewel2023key] proposes to regularize the attention projections by these two items.
+[35cao2023concept] decouples the conditional guidance into two separate processes, which allows for the distinct handling of subject fidelity and textual alignment.
 Despite these efforts, there still remains room for further exploration and refinement of this issue.
@@
 Recent advancements in multimodal autoregressive models present novel solutions for PCS by unifying cross-modal understanding and generation.
-Models like Emu3 [164] demonstrate that autoregressive architectures can natively handle image-text-video sequences through discrete tokenization and joint transformer training.
+Models like Emu3 [wang2024emu3] demonstrate that autoregressive architectures can natively handle image-text-video sequences through discrete tokenization and joint transformer training.
 This paradigm enables seamless integration of user-provided multimodal references (e.g., text descriptions paired with SoI images) while maintaining contextual coherence across generation steps.
@@
 This capability allows users to iteratively refine and accurately define the SoI, addressing the challenge of translating vague or complex requirements into precise content generation.
-For instance, conversational PCS systems like Gemini-2.0-flash [165] exemplify this progress, leveraging natural language dialogue to iteratively optimize both subject fidelity and prompt alignment.
+For instance, conversational PCS systems like Gemini-2.0-flash [team2023gemini] exemplify this progress, leveraging natural language dialogue to iteratively optimize both subject fidelity and prompt alignment.
 By enabling users to provide real-time feedback and adjust parameters through chat-like interactions, these systems bridge the gap between abstract intent and concrete outputs, aligning with PCS’s core objective of balancing faithful subject representation with flexible editability.
@@
 Paper Scope Framework Backbone Technique
-Textual Inversion [7] Object TTF LDM Token Embedding Fine-tuning
-DreamBooth [4] Object TTF Imagen Diffusion Model Fine-tuning; Regularization Dataset
-DreamArtist [92] Object TTF LDM Negative Prompt Fine-tuning
-DVAR [90] Object TTF SD 1.5 Randomness Erasing
-HiPer [166] Object TTF SD Token Embedding Enhancement
-P+ [18] Object TTF SD 1.4 Token Embedding Enhancement
-Unet-finetune [23] Object TTF SD Parameter-efficient Fine-tuning
-Jia et al. [76] Object TTF Imagen Mask-assisted Generation; Regularization
-COTI [83] Object TTF SD 2.0 Data Augmentation
-Gradient-Free TI [94] Object TTF SD Evolutionary Strategy
-PerSAM [44] Object TTF SD Mask-assisted Generation
-DisenBooth [77] Object TTF SD 2.1 Regularization
-NeTI [30] Object TTF SD 1.4 Token Embedding Enhancement
-Prospect [31] Object TTF SD 1.4 Token Embedding Enhancement
-Break-a-Scene [38] Object TTF SD 2.1 Attention-based Operation; Attention-based Operation; Mask-assisted Generation
-COMCAT [96] Object TTF SD 1.4 Attention-based Operation
-ViCo [71] Object TTF SD Attention-based Operation; Regularization
-OFT [167] Object TTF SD 1.5 Finetuning Strategy
-LyCORIS [36] Object TTF SD 1.5 Attention-based Operation
-He et al. [40] Object TTF SD, SDXL Data Augmentation
-DIFFNAT [168] Object TTF SD Regularization
-MATTE [91] Object; Style; High-level Semantic TTF SD Regularization
-LEGO [116] Object TTF SD Regularization
-Catversion [95] Object TTF SD 1.5 Embedding Concatenation
-CLiC [45] Object TTF SD 1.4 Attention-based Operation; Mask-assisted Generation
-HiFi Tuner [32] Object TTF SD 1.4 Attention-based Operation; Mask-assisted Generation; Regularization
-InstructBooth [93] Object TTF SD 1.5 Reinforcement Learning
-DETEX [49] Object TTF SD 1.4 Mask-assisted Generation
-DreamDistribution [169] Object TTF SD 2.1 Token Embedding Enhancement
-DreamTuner [39] Object TTF SD Attention-based Operation
-Xie et al. [34] Object TTF SD Regularization; Mask-assisted Generation
-Infusion [170] Object TTF SD 1.5 Attention-based Operation
-Pair Customization [171] Object TTF SDXL Disentanglement Approach
-DisenDreamer [172] Object TTF SD 2.1 Disentanglement Approach
-DreamBooth+++++ + [173] Object TTF SD Attention-based Operation
-CoRe [174] Object TTF SD 2.1 Regularization
-DCI [175] Object TTF SD Token Embedding Optimization
-Re-Imagen [97] Object PTA Imagen Retrieval-augmented Paradigm
-Versatile Diffusion [176] Object PTA SD 1.4 Architecture Design
-Tuning-Encoder [88] Object PTA SD Regularization
-ELITE [47] Object PTA SD 1.4 Attention-based Operation; Mask-assisted Generation
-UMM-Diffusion [61] Object PTA SD 1.5 Reference Feature Injection
-SuTI [52] Object PTA Imagen Data Augmentation
-InstantBooth [98] Object PTA SD 1.4 Patch Feature Extraction
-BLIP-Diffusion [57] Object PTA SD 1.5 Data Augmentation; Mask-assisted Generation
-Domain-Agnostic [89] Object PTA SD Regularization; Attention-based Operation
-IP-Adapter [51] Object PTA SD 1.5 Reference Feature Injection
-Kosmos-G [60] Object; Face PTA SD 1.5 Multimodal Language Modeling
-Kim et al. [177] Object PTA SD 1.4 Reference Feature Injection
-CAFÉ [58] Object; Face PTA SD 2.0 Multimodal Language Modeling
-SAG [178] Object; Face; Style PTA SD 1.5 Gradient-based Guidance
-BootPIG [179] Object PTA SD Data Augmentation
-CustomContrast [180] Object PTA SD 1.5 Contrastive Learning
-MoMA [59] Object PTA SD 1.5 Multimodal LLM Adapter
-Diptych Prompting [101] Object; Style PTA FLUX Attention-based Operation; Mask-assisted Generation
-StyleDrop [24] Style; Multiple Subjects TTF Muse Data Augmentation
-StyleBoost [84] Style TTF SD 1.5 Regularization Dataset
-StyleAligned [99] Style TTF SDXL Regularization
-GAL [19] Style; Object TTF SD 1.5 Data Augmentation
-StyleForge [181] Style TTF SD 1.5 Parameter-efficient Fine-tuning
-FineStyle [182] Style TTF Muse Mask-assisted Generation
-Style-Friendly [100] Style TTF FLUX-dev; SD 3.5 SNR Sampler
-StyleAdapter [82] Style PTA SD Data Augmentation; Reference Feature Fxtraction
-ArtAdapter [183] Style PTA SD 1.5 Data Augmentation
-ProFusion [106] Face TTF SD 2.0 Fusion Sampling
-Celeb Basis [65] Face TTF SD Face Feature Representation
-Banerjee et al. [184] Face TTF SD 1.4 Regularization
-Magicapture [37] Face; Multiple Subjects TTF SD 1.5 Attention-based Operation; Mask-assisted Generation; Identity Loss
-Concept-centric [163] Face TTF SD 1.5 Modified Classifier-free Guidance
-Cross Initialization [33] Face TTF SD 2.1 Regularization
+Textual Inversion [1gal2022image] Object TTF LDM Token Embedding Fine-tuning
+DreamBooth [2ruiz2023dreambooth] Object TTF Imagen Diffusion Model Fine-tuning; Regularization Dataset
+DreamArtist [68dong2022dreamartist] Object TTF LDM Negative Prompt Fine-tuning
+DVAR [69voronov2024loss] Object TTF SD 1.5 Randomness Erasing
+HiPer [33han2023highly] Object TTF SD Token Embedding Enhancement
+P+ [23voynov2023p+] Object TTF SD 1.4 Token Embedding Enhancement
+Unet-finetune [72xiang2023closer] Object TTF SD Parameter-efficient Fine-tuning
+Jia et al. [24jia2023taming] Object TTF Imagen Mask-assisted Generation; Regularization
+COTI [3yang2023controllable] Object TTF SD 2.0 Data Augmentation
+Gradient-Free TI [73fei2023gradient] Object TTF SD Evolutionary Strategy
+PerSAM [21zhang2023personalize] Object TTF SD Mask-assisted Generation
+DisenBooth [18chen2023disenbooth] Object TTF SD 2.1 Regularization
+NeTI [61alaluf2023neural] Object TTF SD 1.4 Token Embedding Enhancement
+Prospect [30zhang2023prospect] Object TTF SD 1.4 Token Embedding Enhancement
+Break-a-Scene [74avrahami2023break] Object TTF SD 2.1 Attention-based Operation; Attention-based Operation; Mask-assisted Generation
+COMCAT [75xiao2023comcat] Object TTF SD 1.4 Attention-based Operation
+ViCo [20hao2023vico] Object TTF SD Attention-based Operation; Regularization
+OFT [59qiu2024controlling] Object TTF SD 1.5 Finetuning Strategy
+LyCORIS [76yeh2023navigating] Object TTF SD 1.5 Attention-based Operation
+He et al. [57he2023data] Object TTF SD, SDXL Data Augmentation
+DIFFNAT [78roy2023diffnat] Object TTF SD Regularization
+MATTE [79agarwal2023image] Object; Style; High-level Semantic TTF SD Regularization
+LEGO [45motamed2023lego] Object TTF SD Regularization
+Catversion [80zhao2023catversion] Object TTF SD 1.5 Embedding Concatenation
+CLiC [81safaee2023clic] Object TTF SD 1.4 Attention-based Operation; Mask-assisted Generation
+HiFi Tuner [48wang2023hifi] Object TTF SD 1.4 Attention-based Operation; Mask-assisted Generation; Regularization
+InstructBooth [56chae2023instructbooth] Object TTF SD 1.5 Reinforcement Learning
+DETEX [85cai2023decoupled] Object TTF SD 1.4 Mask-assisted Generation
+DreamDistribution [114zhao2023dreamdistribution] Object TTF SD 2.1 Token Embedding Enhancement
+DreamTuner [86hua2023dreamtuner] Object TTF SD Attention-based Operation
+Xie et al. [51lu2024object] Object TTF SD Regularization; Mask-assisted Generation
+Infusion [157zeng2024infusion] Object TTF SD 1.5 Attention-based Operation
+Pair Customization [160jones2024customizing] Object TTF SDXL Disentanglement Approach
+DisenDreamer [171chen2024disendreamer] Object TTF SD 2.1 Disentanglement Approach
+DreamBooth++++ [219fan2024dreambooth] Object TTF SD Attention-based Operation
+CoRe [218wu2024core] Object TTF SD 2.1 Regularization
+DCI [216jin2025customized] Object TTF SD Token Embedding Optimization
+Re-Imagen [67chen2022re] Object PTA Imagen Retrieval-augmented Paradigm
+Versatile Diffusion [109xu2023versatile] Object PTA SD 1.4 Architecture Design
+Tuning-Encoder [12gal2023encoder] Object PTA SD Regularization
+ELITE [19wei2023elite] Object PTA SD 1.4 Attention-based Operation; Mask-assisted Generation
+UMM-Diffusion [71ma2023unified] Object PTA SD 1.5 Reference Feature Injection
+SuTI [11chen2024subject] Object PTA Imagen Data Augmentation
+InstantBooth [22shi2023instantbooth] Object PTA SD 1.4 Patch Feature Extraction
+BLIP-Diffusion [10li2024blip] Object PTA SD 1.5 Data Augmentation; Mask-assisted Generation
+Domain-Agnostic [17arar2023domain] Object PTA SD Regularization; Attention-based Operation
+IP-Adapter [40ye2023ip] Object PTA SD 1.5 Reference Feature Injection
+Kosmos-G [77pan2023kosmos] Object; Face PTA SD 1.5 Multimodal Language Modeling
+Kim et al. [47kim2023personalized] Object PTA SD 1.4 Reference Feature Injection
+CAFÉ [84zhou2023customization] Object; Face PTA SD 2.0 Multimodal Language Modeling
+SAG [99pan2023towards] Object; Face; Style PTA SD 1.5 Gradient-based Guidance
+BootPIG [41purushwalkam2024bootpig] Object PTA SD Data Augmentation
+CustomContrast [217chen2024customcontrast] Object PTA SD 1.5 Contrastive Learning
+MoMA [song2024moma] Object PTA SD 1.5 Multimodal LLM Adapter
+Diptych Prompting [209shin2024large] Object; Style PTA FLUX Attention-based Operation; Mask-assisted Generation
+StyleDrop [5sohn2024styledrop] Style; Multiple Subjects TTF Muse Data Augmentation
+StyleBoost [53park2023styleboost] Style TTF SD 1.5 Regularization Dataset
+StyleAligned [98hertz2023style] Style TTF SDXL Regularization
+GAL [151zhang2024generative] Style; Object TTF SD 1.5 Data Augmentation
+StyleForge [153park2024text] Style TTF SD 1.5 Parameter-efficient Fine-tuning
+FineStyle [213zhangfinestyle] Style TTF Muse Mask-assisted Generation
+Style-Friendly [211choi2024style] Style TTF FLUX-dev; SD 3.5 SNR Sampler
+StyleAdapter [26wang2023styleadapter] Style PTA SD Data Augmentation; Reference Feature Fxtraction
+ArtAdapter [42chen2023artadapter] Style PTA SD 1.5 Data Augmentation
+ProFusion [55zhou2023enhancing] Face TTF SD 2.0 Fusion Sampling
+Celeb Basis [60yuan2023inserting] Face TTF SD Face Feature Representation
+Banerjee et al. [39banerjee2023identity] Face TTF SD 1.4 Regularization
+Magicapture [90hyung2023magicapture] Face; Multiple Subjects TTF SD 1.5 Attention-based Operation; Mask-assisted Generation; Identity Loss
+Concept-centric [35cao2023concept] Face TTF SD 1.5 Modified Classifier-free Guidance
+Cross Initialization [66pang2023cross] Face TTF SD 2.1 Regularization
 Table 3: Paper summary on personalization.
 Paper Scope Framework Technique Backbone
-OMG [27] Face; Multiple Subjects TTF SDXL Attention-based Operation; Mask-assisted Generation; Identity Loss
-InstantFamily [185] Face; Multiple Subjects; Extra Conditions TTF SD 1.5 Mask-assisted Generation
-PersonaMagic [186] Face TTF SD 1.4 Attention-based Operation
-HyperDreamBooth [26] Face TTF; PTA SD 1.5 Pretraining and Fast Fine-tuning
-Su et al. [187] Face PTA LDM; StyleGAN Identity Loss; Multi-task Learning
-FastComposer [5] Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
-Face0 [107] Face PTA SD 1.4 Fusion Sampling
-DreamIdentity [48] Face PTA SD 2.1 Attention-based Operation
-Face-Diffuser [79] Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
-W-Plus-Adapter [53] Face PTA SD 1.5; StyleGAN Face Feature Representation
-Portrait Diffusion [188] Face PTA SD 1.5 Mask-assisted Generation
-RetriNet [74] Face PTA SD Mask-assisted Generation
-FaceStudio [109] Face PTA SD Attention-based Operation
-PVA [189] Face PTA LDM Mask-assisted Generation
-DemoCaricature [121] Face; Extra Conditions PTA SD 1.5 Regularization
-PhotoMaker [46] Face PTA SDXL Data Collection; Reference Feature Extraction
-Stellar [35] Face PTA SDXL Evaluation Prompts and Metrics
-PortraitBooth [50] Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
-InstantID [6] Face; Extra Conditions PTA SDXL Multi-feature Injection
-ID-Aligner [190] Face PTA SD 1.5; SDXL Feedback Learning
-MoA [191] Face; Multiple Subjects PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
-IDAdapter [192] Face PTA SD 2.1 Mixed Facial Features
-Infinite-ID [193] Face PTA SDXL Reference Feature Injection
-Face2Diffusion [162] Face PTA SD Multi-feature Injection
-FreeCure [194] Face PTA SD 1.5; SDXL Mask-assisted Generation
-Omni-ID [195] Face PTA FLUX Face representation enhancement
-Custom Diffusion [20] Multiple Subjects TTF SD 1.4 Parameter-efficient Fine-tuning; Constrained Optimization
-Cones [111] Multiple Subjects TTF SD 1.4 Concept Neurons Activation
-SVDiff [22] Multiple Subjects TTF SD Data Augmentation; Attention-based Operation
-Perfusion [21] Multiple Subjects TTF SD 1.5 Regularization
-Mix-of-Show [25] Multiple Subjects TTF SD 1.5 Attention-based Operation; Mask-assisted Generation
-Cones-2 [70] Multiple Subjects TTF SD 2.1 Attention-based Operation; Regularization; Mask-assisted Generation
-PACGen [75] Multiple Subjects TTF SD 1.4 Mask-assisted Generation
-Compositional Inversion [87] Multiple Subjects TTF SD Attention-based Operation; Regularization
-EM-style [72] Multiple Subjects TTF SD Attention-based Operation; Mask-assisted Generation
-MC2 [196] Multiple Subjects TTF SD 1.5 Attention-based Operation
-MultiBooth [64] Multiple Subjects TTF SD 1.5 Mask-assisted Generation
-Matsuda et al. [197] Multiple Subjects TTF LDM Mask-assisted Generation
-MagicTailor [198] Multiple Subjects TTF SD 2.1 Mask-assisted Generation; Attention-based Operation
-AnyDoor [78] Multiple Subjects PTA SD 2.1 Mask-assisted Generation
-Subject-Diffusion [63] Multiple Subjects PTA SD 2.0 Attention-based Operation; Data Augmentation; Mask-assisted Generation
-CustomNet [113] Multiple Subjects; Extra Conditions PTA SD Multi-condition Integration
-MIGC [114] Multiple Subjects PTA SD Attention-based Operation; Mask-assisted Generation
-Emu2 [199] Multiple Subjects PTA SDXL In-Context Learning
-SSR-Encoder [55] Multiple Subjects PTA SD 1.5 Attention-based Operation
-$\lambda$-eclipse [112] Multiple Subjects PTA Kandinsky v2.2 Contrasstive Learning
-MS-Diffusion [56] Multiple Subjects PTA SD Attention-based Operation
-ReVersion [115] High-level Semantic TTF SD Regularization
-Inv-ReVersion [200] High-level Semantic TTF SD 1.5 Regularization
-CusConcept [201] High-level Semantic TTF SD 2.1 Regularization
-ADI [117] High-level Semantic TTF SD 2.1 Mask-assisted Generation
-PhotoSwap [118] Extra Conditions TTF SD 2.1 Attention-based Operation
-PE-VITON [202] Extra Conditions TTF SD Shape and Texture Control
-Layout-Control [69] Extra Conditions TTF SD Attention-based Operation
-SwapAnything [203] Extra Conditions TTF SD 2.1 Mask-assisted Generation
-PE-VITON [202] Extra Conditions TTF SD Shape and Texture Control
-Viewpoint Control [122] Extra Conditions TTF SDXL 3D Feature Incorporation
-Prompt-Free Diffusion [204] Extra Conditions PTA SD 2.0 Reference Feature Injection
-Uni-ControlNet [205] Extra Conditions PTA SD Multi-feature Injection
-ViscoNet [206] Extra Conditions PTA SD Multi-feature Injection
-Context Diffusion [207] Extra Conditions PTA LDM Multi-feature Injection
-FreeControl [208] Extra Conditions PTA SD 1.5; SD 2.1; SDXL Fusion Guidance
-Li et al. [209] Extra Conditions PTA SD Attention-based Operation; Mask-assisted Generation
+OMG [115kong2024omg] Face; Multiple Subjects TTF SDXL Attention-based Operation; Mask-assisted Generation; Identity Loss
+InstantFamily [161kim2024instantfamily] Face; Multiple Subjects; Extra Conditions TTF SD 1.5 Mask-assisted Generation
+PersonaMagic [203li2024personamagic] Face TTF SD 1.4 Attention-based Operation
+HyperDreamBooth [58ruiz2023hyperdreambooth] Face TTF; PTA SD 1.5 Pretraining and Fast Fine-tuning
+Su et al. [49su2023identity] Face PTA LDM; StyleGAN Identity Loss; Multi-task Learning
+FastComposer [54xiao2023fastcomposer] Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
+Face0 [88valevski2023face0] Face PTA SD 1.4 Fusion Sampling
+DreamIdentity [89chen2023dreamidentity] Face PTA SD 2.1 Attention-based Operation
+Face-Diffuser [52wang2023high] Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
+W-Plus-Adapter [91li2023stylegan] Face PTA SD 1.5; StyleGAN Face Feature Representation
+Portrait Diffusion [32liu2023portrait] Face PTA SD 1.5 Mask-assisted Generation
+RetriNet [92tang2023retrieving] Face PTA SD Mask-assisted Generation
+FaceStudio [31yan2023facestudio] Face PTA SD Attention-based Operation
+PVA [27xu2024personalized] Face PTA LDM Mask-assisted Generation
+DemoCaricature [94chen2023democaricature] Face; Extra Conditions PTA SD 1.5 Regularization
+PhotoMaker [28li2023photomaker] Face PTA SDXL Data Collection; Reference Feature Extraction
+Stellar [95achlioptas2023stellar] Face PTA SDXL Evaluation Prompts and Metrics
+PortraitBooth [96peng2023portraitbooth] Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
+InstantID [37wang2024instantid] Face; Extra Conditions PTA SDXL Multi-feature Injection
+ID-Aligner [162chen2024id] Face PTA SD 1.5; SDXL Feedback Learning
+MoA [164ostashev2024moa] Face; Multiple Subjects PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
+IDAdapter [166cui2024idadapter] Face PTA SD 2.1 Mixed Facial Features
+Infinite-ID [169wu2024infinite] Face PTA SDXL Reference Feature Injection
+Face2Diffusion [170shiohara2024face2diffusion] Face PTA SD Multi-feature Injection
+FreeCure [210cai2024foundation] Face PTA SD 1.5; SDXL Mask-assisted Generation
+Omni-ID [205qian2024omni] Face PTA FLUX Face representation enhancement
+Custom Diffusion [6kumari2023multi] Multiple Subjects TTF SD 1.4 Parameter-efficient Fine-tuning; Constrained Optimization
+Cones [4liu2023cones] Multiple Subjects TTF SD 1.4 Concept Neurons Activation
+SVDiff [8han2023svdiff] Multiple Subjects TTF SD Data Augmentation; Attention-based Operation
+Perfusion [7tewel2023key] Multiple Subjects TTF SD 1.5 Regularization
+Mix-of-Show [46gu2024mix] Multiple Subjects TTF SD 1.5 Attention-based Operation; Mask-assisted Generation
+Cones-2 [62liu2023cones] Multiple Subjects TTF SD 2.1 Attention-based Operation; Regularization; Mask-assisted Generation
+PACGen [34li2023generate] Multiple Subjects TTF SD 1.4 Mask-assisted Generation
+Compositional Inversion [9zhang2023compositional] Multiple Subjects TTF SD Attention-based Operation; Regularization
+EM-style [44rahman2024visual] Multiple Subjects TTF SD Attention-based Operation; Mask-assisted Generation
+MC2 [154jiang2024mc] Multiple Subjects TTF SD 1.5 Attention-based Operation
+MultiBooth [156zhu2024multibooth] Multiple Subjects TTF SD 1.5 Mask-assisted Generation
+Matsuda et al. [165matsuda2024multi] Multiple Subjects TTF LDM Mask-assisted Generation
+MagicTailor [zhou2024magictailor] Multiple Subjects TTF SD 2.1 Mask-assisted Generation; Attention-based Operation
+AnyDoor [14chen2023anydoor] Multiple Subjects PTA SD 2.1 Mask-assisted Generation
+Subject-Diffusion [16ma2023subject] Multiple Subjects PTA SD 2.0 Attention-based Operation; Data Augmentation; Mask-assisted Generation
+CustomNet [29yuan2023customnet] Multiple Subjects; Extra Conditions PTA SD Multi-condition Integration
+MIGC [174zhou2024migc] Multiple Subjects PTA SD Attention-based Operation; Mask-assisted Generation
+Emu2 [223sun2024generative] Multiple Subjects PTA SDXL In-Context Learning
+SSR-Encoder [221zhang2024ssr] Multiple Subjects PTA SD 1.5 Attention-based Operation
+$\lambda$-eclipse [222patel2024lambda] Multiple Subjects PTA Kandinsky v2.2 Contrasstive Learning
+MS-Diffusion [220personalizationms] Multiple Subjects PTA SD Attention-based Operation
+ReVersion [63huang2023reversion] High-level Semantic TTF SD Regularization
+Inv-ReVersion [159zhang2024inv] High-level Semantic TTF SD 1.5 Regularization
+CusConcept [215xu2024cusconcept] High-level Semantic TTF SD 2.1 Regularization
+ADI [43huang2023learning] High-level Semantic TTF SD 2.1 Mask-assisted Generation
+PhotoSwap [15gu2024photoswap] Extra Conditions TTF SD 2.1 Attention-based Operation
+PE-VITON [38zhang2023two] Extra Conditions TTF SD Shape and Texture Control
+Layout-Control [152chen2024training] Extra Conditions TTF SD Attention-based Operation
+SwapAnything [155gu2024swapanything] Extra Conditions TTF SD 2.1 Mask-assisted Generation
+PE-VITON [38zhang2023two] Extra Conditions TTF SD Shape and Texture Control
+Viewpoint Control [163kumari2024customizing] Extra Conditions TTF SDXL 3D Feature Incorporation
+Prompt-Free Diffusion [110xu2023prompt] Extra Conditions PTA SD 2.0 Reference Feature Injection
+Uni-ControlNet [111zhao2024uni] Extra Conditions PTA SD Multi-feature Injection
+ViscoNet [93cheong2023visconet] Extra Conditions PTA SD Multi-feature Injection
+Context Diffusion [112najdenkoska2023context] Extra Conditions PTA LDM Multi-feature Injection
+FreeControl [113mo2023freecontrol] Extra Conditions PTA SD 1.5; SD 2.1; SDXL Fusion Guidance
+Li et al. [168li2024tuning] Extra Conditions PTA SD Attention-based Operation; Mask-assisted Generation
 Table 4: Paper summary on personalization.
 Paper Scope Framework Technique Backbone
-Tune-A-Video [127] Video TTF SD 1.4 Parameter-efficient Fine-tuning
-Gen-1 [130] Video TTF LDM Multi-feature Injection
-Make-A-Protagonist [210] Video TTF LDM Attention-based Operation
-Animate-A-Story [131] Video TTF SD Retrieval-augmented Paradigm
-MotionDirector [134] Video TTF ZeroScope; ModelScope Disentanglement Approach
-LAMP [132] Video TTF SDXL; SD 1.4 First-frame Conditioned Pipeline
-VMC [129] Video TTF Show-1 Parameter-efficient Fine-tuning
-SAVE [128] Video TTF VDM Attention-based Operation
-Customizing Motion [133] Video TTF ZeroScope Regularization
-DreamVideo [135] Video TTF ModelScope Disentanglement Approach
-MotionCrafter [211] Video TTF VDM Disentanglement Approach
-Customize-A-Video [212] Video TTF ModelScope Multi-stage Refinement
-Magic-Me [213] Video TTF SD 1.5; AnimateDiff Multi-stage Refinement
-CustomTTT [214] Video TTF VDM Fine-tuning and Alignment
-PersonalVideo [215] Video; Face TTF AnimateDiff Identity Injection
-CustomVideo [81] Video; Multiple Subjects TTF ZeroScope Data Augmentation; Attention-based Operation
-VideoDreamer [80] Video; Multiple Subjects TTF SD 2.1 Data Augmentation
-VideoAssembler [125] Video PTA VidRD Reference Feature Injection
-VideoBooth [124] Video PTA VDM Reference Feature Injection
-VideoMaker [216] Video PTA VDM Reference Feature Injection
-SUGAR [217] Video PTA VDM Attention-based Operation
-StyleCrafter [54] Video; Style PTA SD 1.5 Reference Feature Injection
-ID-Animator [218] Video; Face PTA AnimateDiff Reference Feature Injection
-ConsisID [219] Video; Face PTA CogVideoX frequency decomposition
-MotionCharacter [220] Video; Face PTA VDM ID-Preserving Optimization; Motion Control Enhancement
-DreaMoving [126] Video; Face; Extra Conditions PTA SD Multi-feature Injection
-Magic3D [138] 3D TTF Imagen; LDM Score Distillation Sampling
-DreamBooth3D [140] 3D TTF Imagen Multi-stage Refinement
-PAS [145] 3D PTA SD Text-to-3D-Pose
-StyleAvatar3D [146] 3D PTA LDM Multi-view Alignment
-AvatarBooth [147] 3D TTF SD Dual Model Fine-tuning
-MVDream [139] 3D TTF SD 2.1 Score Distillation Sampling
-Consist3D [141] 3D TTF SD Token Embedding Enhancement
-Animate124 [143] 3D TTF SD 1.5 Multi-stage Refinement
-Dream-in-4D [144] 3D TTF SD 2.1 Multi-stage Refinement
-TextureDreamer [142] 3D TTF LDM Texture Extraction
-TIP-Editor [221] 3D TTF SD Multi-stage Refinement; Mask-assisted Generation
-Anti-DreamBooth [148] Attack and Defense TTF SD 2.1 Perturbation Learning
-Concept Censorship [149] Attack and Defense TTF SD 1.4 Trigger Injection
-Huang et al. [222] Attack and Defense TTF SD backdoor attack
-Continual Diffusion [223] Others TTF SD Continual Learning
-SVGCustomization [150] Others TTF SD 1.5 Fine-tuning and Alignment
-StitchDiffusion [151] Others TTF LDM Parameter-efficient Fine-tuning
-MC-TI [224] Others TTF SD 1.5 Regularization
+Tune-A-Video [150wu2023tune] Video TTF SD 1.4 Parameter-efficient Fine-tuning
+Gen-1 [148esser2023structure] Video TTF LDM Multi-feature Injection
+Make-A-Protagonist [135zhao2023make] Video TTF LDM Attention-based Operation
+Animate-A-Story [136he2023animate] Video TTF SD Retrieval-augmented Paradigm
+MotionDirector [101zhao2023motiondirector] Video TTF ZeroScope; ModelScope Disentanglement Approach
+LAMP [102wu2023lamp] Video TTF SDXL; SD 1.4 First-frame Conditioned Pipeline
+VMC [140jeong2023vmc] Video TTF Show-1 Parameter-efficient Fine-tuning
+SAVE [103song2023save] Video TTF VDM Attention-based Operation
+Customizing Motion [104materzynska2023customizing] Video TTF ZeroScope Regularization
+DreamVideo [142wei2023dreamvideo] Video TTF ModelScope Disentanglement Approach
+MotionCrafter [106zhang2023motioncrafter] Video TTF VDM Disentanglement Approach
+Customize-A-Video [172ren2024customize] Video TTF ModelScope Multi-stage Refinement
+Magic-Me [173ma2024magic] Video TTF SD 1.5; AnimateDiff Multi-stage Refinement
+CustomTTT [202bi2024customttt] Video TTF VDM Fine-tuning and Alignment
+PersonalVideo [208li2024personalvideo] Video; Face TTF AnimateDiff Identity Injection
+CustomVideo [145wang2024customvideo] Video; Multiple Subjects TTF ZeroScope Data Augmentation; Attention-based Operation
+VideoDreamer [138chen2023videodreamer] Video; Multiple Subjects TTF SD 2.1 Data Augmentation
+VideoAssembler [82zhao2023videoassembler] Video PTA VidRD Reference Feature Injection
+VideoBooth [83jiang2023videobooth] Video PTA VDM Reference Feature Injection
+VideoMaker [201wu2024videomaker] Video PTA VDM Reference Feature Injection
+SUGAR [204zhou2024sugar] Video PTA VDM Attention-based Operation
+StyleCrafter [212liu2024stylecrafter] Video; Style PTA SD 1.5 Reference Feature Injection
+ID-Animator [158he2024id] Video; Face PTA AnimateDiff Reference Feature Injection
+ConsisID [207yuan2024identity] Video; Face PTA CogVideoX frequency decomposition
+MotionCharacter [206fang2024motioncharacter] Video; Face PTA VDM ID-Preserving Optimization; Motion Control Enhancement
+DreaMoving [105feng2023dreamoving] Video; Face; Extra Conditions PTA SD Multi-feature Injection
+Magic3D [120lin2023magic3d] 3D TTF Imagen; LDM Score Distillation Sampling
+DreamBooth3D [121raj2023dreambooth3d] 3D TTF Imagen Multi-stage Refinement
+PAS [122azadi2023text] 3D PTA SD Text-to-3D-Pose
+StyleAvatar3D [123zhang2023styleavatar3d] 3D PTA LDM Multi-view Alignment
+AvatarBooth [124zeng2023avatarbooth] 3D TTF SD Dual Model Fine-tuning
+MVDream [126shi2023mvdream] 3D TTF SD 2.1 Score Distillation Sampling
+Consist3D [127ouyang2023chasing] 3D TTF SD Token Embedding Enhancement
+Animate124 [128zhao2023animate124] 3D TTF SD 1.5 Multi-stage Refinement
+Dream-in-4D [130zheng2023unified] 3D TTF SD 2.1 Multi-stage Refinement
+TextureDreamer [133yeh2024texturedreamer] 3D TTF LDM Texture Extraction
+TIP-Editor [134zhuang2024tip] 3D TTF SD Multi-stage Refinement; Mask-assisted Generation
+Anti-DreamBooth [13van2023anti] Attack and Defense TTF SD 2.1 Perturbation Learning
+Concept Censorship [50zhang2023backdooring] Attack and Defense TTF SD 1.4 Trigger Injection
+Huang et al. [167huang2024personalization] Attack and Defense TTF SD backdoor attack
+Continual Diffusion [64smith2023continual] Others TTF SD Continual Learning
+SVGCustomization [36zhang2023text] Others TTF SD 1.5 Fine-tuning and Alignment
+StitchDiffusion [25wang2024customizing] Others TTF LDM Parameter-efficient Fine-tuning
+MC-TI [214wang2024multi] Others TTF SD 1.5 Regularization
 Acknowledgments
@@
 The authors declared that they have no conflicts of interest to this work.
+Xulu Zhang received the B.E., and M.S.
+degrees from Sichuan University in 2019 and 2022.
+He is currently working towards the Ph.D degree from the Department of Computing at the Hong Kong Polytechnic University.
+His research interests include image generation and active learning.
+E-mail: compxulu.zhang@connect.polyu.hk
+ORCID iD: 0000-0003-2473-460X
+Xiaoyong Wei received the Ph.D.
+degree in computer science from City University of Hong Kong, China in 2009, and has worked as a postdoctoral fellow in the University of California, Berkeley, USA from December, 2013 to December, 2015.
+He has been a professor and the head of Department of Computer Science, Sichuan University, China since 2010.
+He is an adjunct professor of Peng Cheng Laboratory, China, and a visiting professor of Department of Computing, Hong Kong Polytechnic University.
+He is a senior member of IEEE, and has served as an associate editor of Interdisciplinary Sciences: Computational Life Sciences since 2020, the program Chair of ICMR 2019, ICIMCS 2012, and the technical committee member of over 20 conferences such as ICCV, CVPR, SIGKDD, ACM MM, ICME, and ICIP.
+His research interests include multimedia computing, health computing, machine learning and large-scale data mining.
+E-mail: x1wei@polyu.edu.hk (Corresponding Author)
+ORCID iD: 0000-0002-5706-5177
+Wentao Hu received the B.E., and M.S.
+degrees from Shandong University and Sun Yat-sen University, in 2021 and 2024.
+He is currently working towards the Ph.D degree from the Department of Computing at the Hong Kong Polytechnic University.
+His research interests include image generation and 3D reconstruction.
+E-mail: wayne-wt.hu@connect.polyu.hk
+ORCID iD: 0000-0002-2071-9341
+Jinlin Wu is an Assistant Research Fellow at the Institute of Automation, Chinese Academy of Sciences (CAS).
+He received his Ph.D.
+from the University of Chinese Academy of Sciences in 2022 and his bachelor’s degree from the University of Electronic Science and Technology of China in 2017.
+His research interests include object detection, image recognition, and video understanding.
+He has served as the principal investigator of a National Natural Science Foundation of China (NSFC) Youth Science Fund project and has participated in several other NSFC-funded projects.
+Dr.
+Wu has accumulated a solid research foundation in areas such as video analysis for security and medical video understanding.
+He has published over 30 high-quality academic papers, with more than 500 citations.
+E-mail: jinlin.wu@cair-cas.org.hk
+ORCID iD: 0000-0001-7877-5728
+Jiaxin Wu is a Postdoctoral Fellow in the Department of Computing at The Hong Kong Polytechnic University.
+She earned her Ph.D.
+in Computer Science from the City University of Hong Kong.
+Her research interests include multimedia retrieval, AI for Science (AI4Science), and natural language processing (NLP).
+E-mail: jiaxwu@polyu.edu.hk
+ORCID iD: 0000-0003-4074-3442
+Wengyu Zhang is an Undergraduate Student in the Department of Computing at The Hong Kong Polytechnic University.
+His research interests include natural language processing (NLP), AI for Science (AI4Science), and Graph Learning.
+E-mail: wengyu.zhang@connect.polyu.hk
+ORCID iD: 0009-0001-2347-4183
+Zhaoxiang Zhang received his BSc degree in Department of Electronic Science and Technology from University of Science and Technology of China, and the PhD degree in Pattern Recognition and Intelligent Systems from the Institute of Automation, Chinese Academy of Sciences in 2004 and 2009, respectively.
+From 2009 to 2015, he worked as a Lecturer, Associate Professor, and later the deputy director of Department of Computer Application Technology at the Beihang University.
+Since July 2015, Dr.
+Zhang has joined the Institute of Automation, Chinese Academy of Sciences, where he is currently a professor.
+His major research interests include pattern recognition, computer vision and machine learning.
+Recently, he specifically focuses on deep learning models, biologically-inspired visual computing and human-like learning, and their applications on human analysis and scene understanding.
+He has published more than 200 papers in international journals and conferences, including reputable international journals such as IEEE TPAMI, IJCV, JMLR, IEEE TIP, IEEE TCSVT, and top-level international conferences like CVPR, ICCV, NIPS, ECCV, ICLR, AAAI, IJCAI and ACM MM.
+He has won the best paper awards in several conferences and championships in international competitions and his research has won the ‘Technical Innovation Award of the Chinese Association of Artificial Intelligence’.
+He has served as the PC Chair or area chair of many international conferences like CVPR, ICCV, AAAI, IJCAI, ACM MM, ICPR and BICS.
+He is serving or has served as associate editor of reputable international journals like IJCV, IEEE T-CSVT, IEEE T-BIOM, Pattern Recognition and NeuroComputing.
+E-mail: zhaoxiang.zhang@ia.ac.cn
+ORCID iD: 0000-0003-2648-3875
+Zhen Lei received the B.S.
+degree in automation from the University of Science and Technology of China, in 2005, and the Ph.D.
+degree from the Institute of Automation, Chinese Academy of Sciences, in 2010, where he is currently a professor.
+He is IEEE Fellow, IAPR Fellow and AAIA Fellow.
+He has published over 200 papers in international journals and conferences with 33000+ citations in Google Scholar and h-index 86.
+He was the program co-chair of IJCB2023, was competition co-chair of IJCB2022 and has served as area chairs for several conferences and is associate editor for IEEE Trans.
+on Information Forensics and Security, IEEE Transactions on Biometrics, Behavior, and Identity Science, Pattern Recognition, Neurocomputing and IET Computer Vision journals.
+His research interests are in computer vision, pattern recognition, image processing, and face recognition in particular.
+He is the winner of 2019 IAPR Young Biometrics Investigator Award.
+E-mail: zhen.lei@ia.ac.cn (Corresponding Author)
+ORCID iD: 0000-0002-0791-189X
+Qing Li is currently a Chair Professor (Data Science) and the Head of the Department of Computing, the Hong Kong Polytechnic University.
+Formerly, he was the founding Director of the Multimedia software Engineering Research Centre (MERC), and a Professor at City University of Hong Kong where he worked in the Department of Computer Science from 1998 to 2018.
+Prior to these, he has also taught at the Hong Kong University of Science and Technology and the Australian National University (Canberra, Australia).
+Prof.
+Li served as a consultant to Microsoft Research Asia (Beijing, China), Motorola Global Computing and Telecommunications Division (Tianjin Regional Operations Center), and the Division of Information Technology, Commonwealth Scientific and Industrial Research Organization (CSIRO) in Australia.
+He has been an Adjunct Professor of the University of Science and Technology of China (USTC) and the Wuhan University, and a Guest Professor of the Hunan University (Changsha, China) where he got his BEng.
+degree from the Department of Computer Science in 1982.
+He is also a Guest Professor (Software Technology) of the Zhejiang University (Hangzhou, China) – the leading university of the Zhejiang province where he was born.
+Prof.
+Li has been actively involved in the research community by serving as an associate editor and reviewer for technical journals, and as an organizer/co-organizer of numerous international conferences.
+Some recent conferences in which he is playing or has played major roles include APWeb-WAIM’18, ICDM 2018, WISE2017, ICDSC2016, DASFAA2015, U-Media2014, ER2013, RecSys2013, NDBC2012, ICMR2012, CoopIS2011, WAIM2010, DASFAA2010, APWeb-WAIM’09, ER’08, WISE’07, ICWL’06, HSI’05, WAIM’04, IDEAS’03,VLDB’02, PAKDD’01, IFIP 2.6 Working Conference on Database Semantics (DS-9), IDS’00, and WISE’00.
+In addition, he served as a programme committee member for over fifty international conferences (including VLDB, ICDE, WWW, DASFAA, ER, CIKM, CAiSE, CoopIS, and FODO).
+He is currently a Fellow of IEEE and IET/IEE, a member of ACM-SIGMOD and IEEE Technical Committee on Data Engineering.
+He is the chairperson of the Hong Kong Web Society, and also served/is serving as an executive committee (EXCO) member of IEEE-Hong Kong Computer Chapter and ACM Hong Kong Chapter.
+In addition, he serves as a councilor of the Database Society of Chinese Computer Federation (CCF), a member of the Big Data Expert Committee of CCF, and is a Steering Committee member of DASFAA, ER, ICWL, UMEDIA, and WISE Society.
+E-mail: qing-prof.li@polyu.edu.hk
+ORCID iD: 0000-0003-3370-471X